A Complete Bibliography of Publications in the Journal of Algorithms

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.05

Title word cross-reference

(1 ≤ k ≤ m ≤ n) [Sem84]. (2 + ϵ) [DIP02, QS02]. (h, k, 1) [Yao80]. (s, t) [Wei97]. 1 [GP87]. 2 [AS96b, ALS03, ADNP99, CHC99, Nag05, ŠŽ05b]. 2n^{2} [Hal97a]. 3 [ADNP99, BE05, DN99, DF86]. 4 [HZ01a]. 4, 5 [DN99]. 4/3 [Nag05, ŠŽ05b]. 5/3 [PS00]. 8/13 [Blä04]. [1, 2, , n] [Sem84]. ^{2} [FL06]. _{4} [BGLR99, NP04]. α [BDV02, II96]. B [GLW83]. d [vLW81]. Δ [KN88, MS03]. E [Mal94b]. ϵ [Car01]. f [Ans85]. g [Mal94a]. GF(2) [CRS^{+}00, SR03]. GF(qn) [LKPC05]. GF[2] [KL93]. GL(n, Z) [Fer87]. h [SSO01, CDF01, GL98]. k [AS96a, AIKS91, ATUW01, ALP04, Ano98f, BYR02, BR97, BM05, BS95b, Bod90, BDGW98, CT99, DLSS95, DGK01, FKN99, Fer98, FJ83, FS96, HK02, HNZ02, KP98b, NH01, NRP96, Pem05a, PSY93, Raj95, Sch98, Sem84, She95, SU97, Sub99, Tur88, WPYK06, ZiNN96, vB00, BGLR99]. K_{3,3} [Che98c]. K_{5} [Che98c]. LUP [IMH82]. m [Etz86, GLD94, JZC04]. N [CS92, Fer87]. NC [CGM^{+}95, ZSN97, NH01]. n log n [Kla99]. NR(1) [Asp80]. O(log ∗n) [PV98]. O(log n) [SV82a]. O(log n log log n) [CL95]. O(m + n log n) [BK97]. O(N) [IM01b, HS81]. O(n^{1.3289}) [BE05]. O(N^{2})

[TW84, EGMS94, HG88, MS94a, BK00]. O(n^{2.5}) [Chu87]. O(n^{2} log n) [SV82b]. o(n^{3}) [KS98]. O(n log log n) [Han04, Tho02]. O(n log m) [Llo83]. O(n log n) [DKR82, ML84, MZ86]. O(nm) [NI99]. O(\sqrt{E}) [Mal94b]. O(\sqrt{g}) [Mal94a]. p [AEH01, FJ83, PV98, NP04]. p(n) [JS93]. π(x) [LO87]. Q [Ron94]. r [DB97]. S [CS92, Gab98]. t [MW94]. V [Szp88]. w [TSB05b]. Z [Fer87, HK01].

- [FS96]. -Algorithms [CGM^{+}95]. -Approximate [Car01]. -Approximation [ADNP99, DIP02, DN99, Blä04, QS02, Nag05]. -Approximations [HK01]. -ary [Etz86, Pem05a, She95, vB00, Szp88]. -Center [PV98]. -centers [FJ83]. -Client [ATUW01]. -Colorable [BS95b, Sub99, HNZ02, Tur88]. -Coloring [CDF01, BE05]. -comparability [NP04]. -competitive [ŠŽ05b]. -connectivity [II96]. -d [DGK01]. -Dominating [KP98b]. -Edge-Connected [Fer98]. -Edge-Connectivity [NH01]. -Euclidean [BDV02]. -factor [Ans85]. -Free [GL98, Che98c]. -linearly [Fer87]. -Link [Sch98]. -local [ŠŽ05b]. -local-vertex-connectivity [Nag05]. -Node [CT99]. -Obstructions [GLD94]. -Overlap [NRP96]. -Partite [BYR02]. -Paths [Wei97]. -Point [DLSS95]. -Relations [SSO01]. -Round [MW94]. -SAT [DB97, HZ01a]. -Server [BR97, BM05]. -set [JZC04]. -Sets [Gab98]. -Shellsort [Yao80]. -Shredders [CT99]. -sorters [CS92]. -space [vLW81]. -Steiner [BDGW98, GP87]. -stepping [MS03]. -subsets [Sem84]. -Track [FKN99]. -Tree [PSY93, SU97, WPYK06]. -Trees [Ano98f, HK02, ZiNN96, Bod90, TSB05b, GLW83]. -Vertex-Connected [ADNP99, DN99].

0-density [Kla81].

1-20 [Ano96b]. 1-center [MZ86]. 1-polynomials [FM05]. 1-solvable [BMZ90]. 11-Competitive [CL94].

2- [PSK97]. 2-Approximation [Hen97, DLZZ03]. 2-Center [HSG02]. 2-Chain [MS94b]. 2-d [Cha98]. 2-Dimensional [Rup95, Ste87]. 2-Monotonic [Ano98e]. 2-Spanners [KP94]. 2-variable [MS90b]. 246 [Ano81c].

3-clique-ordering [Koc05]. 3-coloring [Koc05]. 3-connected [CM88]. 3-cut [He91b]. 3-d [Cha98]. 3-polytopes [BDRG04]. 3-Space [CKS^{+}98]. 3-Uniform [KNS01]. 3-Vertex [PSK97].

4-connected [CN89]. 416 [Ano02i]. 42 [BN03].

5-coloring [CNS81]. 54 [EIT^{+}93].

’95 [Cla98].

above [MR99]. Access [Rou01, HHL88, Vis83]. account [Bey83]. Achromatic [Ano02i, CV01]. acknowledgment [FLNU03]. Acyclic [BS97a, HT97, LS99, Squ98, BFCP^{+}05, Blu87, Vis92a]. ad [FFU04]. ad-hoc [FFU04]. Adaptive [BPK00, IS96, LP93, ABW03, BCd09, HS84]. addendum [SR03]. Addition [BN02, BBBD89, KKB88, Pip02, Tho02, BN03, Oli81, Zan91]. additions [Ran83]. additive [Pan83]. address [Vis83]. Adjacent [Rus88]. Adjusting [HM93, She95, Elm04]. Admission [KPP98, GHP05]. Advantages [DGK01]. Adversarially [You00]. Adversary [AH98]. again [Dol82]. Against [You00]. agent [GHCMS09]. AgentSpeak [GHCMS09]. aggregates [CS06]. Agreement [MW94]. Algebra [DK95]. Algebraic [FRW94, RT97, HY88, Hua91a, Mig82, SY82]. Algebras [Ron94, HH81, Rón88b]. Algorithm [AT95, AO00, AR00, Alt93, Ano98e, ADNP99, BKSD97, BK97, BS97b, BK93, BV99, BPK00, Bur96, CFFK98, CB01, Che98b, Che01a, CL94, CJM02, CZ95, Csi93, DHKP97, DN99, DC93, DF94, Dor98, EGMS94, Eng95, FKN99, FN01, FGP01, FR95, GKR00, GR98, Gol97, Goo11, Hal97a, HO94, He93, He01, Hen97, HM96, Hva96, IK94, IM01b, JMB96, JM95, KS96, KLST01, KPT96, KRT94, Kla99, Kle93, KR95b, KS97, LS97, MS94a, MLP99, NH01, NW00, PV98, PS00, RR96, Rup95, SU97, Sla97, WMM95, YM99, Yeh01, Yeo99, ZSN96, ZSN97, Zhu95, vB00, ADKP89, ABI86, ADK^{+}03, AS80, AK92, Atk87, BP04, BBK81, Bak85, BJMT92, BYE81, Ben80, BS91a, BBH^{+}06, BMX06, Blä04, Blu87, BF84, CG92, CNS81, Chu92, CG84, CS92, DF02, DLZZ03]. algorithm [DK85, DKR82, DA85, EA81, Etz86, Fer87, FGS92, Fra81, FL86, Fri88, GHP05, GMM03, Güt84, HS84, HM85, He91b, HG88, HNS86, Hof84, HK86, HS81, Hua90, IMH82, Ili82, Irv85, JMVW02, JZC04, Kam88, KN88, Kin88, Kla81, Kuc91, Lar87, LLZ05, LP82, LS87b, LL88, LP81a, Llo83, Loz83, MP89, Mai88, ML84, Mak02, MZ86, MS03, MS87, MS80, Nos85, OAMB86, Pem05b, PH87, QS02, Ral81, Sch88a, SHOS90, Sch88b, Sch05b, Sem84, SV82a, SV82b, ŠŽ05b, Spr89, TSB05a, Tiw86, Tol82, TW84, Vak03, Val91, Vis92a, Vis04, War86, WTM05, Wol86, ZLL09, ZG91]. Algorithmic [ADL^{+}94, BBDS00, JS93, KL94, Lie82, Ans85, Ave92, LLL86, SL06, SSZ86, RS86]. Algorithms [AR93, AMO99, AC98, AK98, AGK04, AT00, Ano02i, AW02, BN03, BR97, BGPW97, BSV93, BDMR01, BG99, BBDL01, BK96, BGLR99, BKRS96, Car01, CRS^{+}00, Cha98, CC00, CH01a, CGM^{+}95, CCyC^{+}99, CV01, CM96, Che01b, CJ92, CT99, CLRW97, CN00, CS99, CMN96, Dah00, DGM01, DLSS95, DK95, DW04, Epp94, FL01, FMSN00, Fro00, GT88, Gab98, GKT01, GGK^{+}97, GIK02, GG97, GPV93, GT01, GP00, GK99, GBH97, HZ01a, HZ01b, Heb98, HK93, Hen97, HKAA97, HN96, IWJ93, IZ94, KMSY98, KK06, KT93, KR96, KPRT97, KPR01, Lag96, LW95, LR99, MS98, MS94c, MT91, NRP96, NMR01, Old01, PV93, PSY93, PtL96, PSK97, Pis99, Poo93, Ram97, RT97, Rob86, Rou01, SCGD94, Sor94, TSB05b, Thu97, WG97, Wan00, Zha98, AP90, AP03, AM82, Ale85]. algorithms [ALP04, AW05, Aur88, AISN91, ABK92, AERW04, BNL03, BM05, BHS07, BP83, BDM03, BRG89, Bey83, BT83, Bod90, BDV02, Cab07, CGM92, CH03, CHT02, CY89, CN90, CL91, Chu87, CFGK91, CSS89, CR06, DM03, ES03, Fei81, Fei83a, Fei83b, FKL^{+}91, GT84, GKS04, GM08, Gol92, GS84, GNO89, HM89, HY88, He91a, HM87, JLR06, Kan85, KS87, KT80, KLM89, Lag80, Len87, Leu84, LC92, MMS90, Mey03, MS05, NT90, NR03, NP04, NPT92, Pac87, Pan80, Pan81a, Pan81b, Pan83, PC91, RS03, RH84, Ros82, SR03, Sch82, SZ05a, SL84, SU84, SZ90, SSW84, Sor88a, TAB86, Vit82, Wah04, WLK04, Wan05, WPYK06, Wil91]. Alignment [LZU01, WG97, SFBY04]. alignments [AS92]. All-norm [AERW04]. All-Pairs [ACZ98, CZ01, BHS07]. All-Sources [DK95]. Allocate [BKP93]. Allocation [FCL00, Hag95, KP98a, Knu96, Win01, BC82]. allowing [BKN04]. Almost [DG01, LR98, Tur88, COKN06]. Almost-Greedy [DG01]. Alphabet [AC00, Duv83]. Alphabet-Independent [AC00]. Alphabetic [LP98, BK06]. Alphabets [Bre95]. Amalgamation [BBDS00]. Amidst [AS96b, LS87b]. Amitabh [Ano02i]. among [CND03, OvdS96, Pri83]. amortized [Boy05]. Analogs [FHS03]. Analysis [CH03, FGP01, FS96, HW00a, Ili82, KRS98, Knu96, KPR00, Lar84, Lue98, MP89, Pip02, SS93, Sla97, Zhu04, dlTK97, BW85, BDV02, CHT02, CFGK91, FFV80, GM84, HM91, Hof84, HK86, KM03, Knu85a, Pan81b, PPR04, PM85, Sch82, SU84, Spr89, Yao80]. analytic [LO87]. Ancestors [AT00, BFCP^{+}05]. Ancient [KPT96]. AND/OR [CG92, DCD96, Hva96]. Anonymous [HKAA97, KK97, AS91]. any [Blu00, Tam85]. Appear [Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02j, Ano02k, Ano02l, Ano02q, Ano84c, Ano84d, Ano84e, Ano84f, Ano85c, Ano85d, Ano85e, Ano85f, Ano86c, Ano86d, Ano86e, Ano86f, Ano87c, Ano87d, Ano87e, Ano87f, Ano88c, Ano88d, Ano88e, Ano88f, Ano89c, Ano89d, Ano89e, Ano89f, Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano92c, Ano92d, Ano92e, Ano92f, Ano02m, Ano02n, Ano02o]. appear [Ano02p, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano07e]. Application [yHHsH^{+}00, Sie01, HIL03, dLFM07]. Applications [AST94, AEH01, Blu00, Dah00, DR99, DDSV93, ES03, HK02, He93, HR99, HW00b, JPS01, KP98b, Mir98, Shp00, Ver94, AP90, AS92, CY91, CJ91, CHT02, CM05, IMH82, LS91]. Applied [KVY94]. Approach [HS95a, NRP96, Pip02, dlTK97, CDBPL^{+}09, DLMM05, Dra05, Pan81b, PCBC^{+}09, dBCO92]. approximants [BGY80]. Approximate [AP98, BC99, Car01, Kri97, KS03, WMM95, ZSW94, LV89, SR06, Ukk85]. Approximately [Ind01, Lar87]. Approximating [BY01, BYR02, BHP01, BGHK95, BCD99, CL99, DPPS01, Fuj99, FR94, Hoc98, KL93, KKM99, KNS01, LK98, LR98, BYES05, MSV06]. Approximation [AC98, AP03, Ano02i, AW02, ADNP99, BDMR01, BDM03, BRG89, BF02, Cab07, CFFK98, CH01a, CH01b, CCyC^{+}99, CV01, CB01, Che98c, Che01b, CS99, Czu96, CGPR97, DIP02, DN99, DM03, FL01, Fer98, GKS04, GKR00, GIK02, Gup01, GBH97, HZ01a, Hen97, HN96, HMR^{+}98, Jan00, KT93, KVY94, KR96, KR95b, KPRT97, LR99, MS05, NRP96, Old01, PV98, PSK97, PS00, WG97, WCT00, Yan94, AERW04, BKL^{+}04, BYE81, BST99, Bla92, Blä04, Cha03, DLZZ03, FKL02, GMM03, HM87, JMVW02, JP02, KLM89, Mas06, Nag05, QS02, Vis04, Wu02, ZLL09]. Approximations [AHL06, BR94, HK01]. Arbitrary [AK97, HM96, adHV99, BBH^{+}06, FMSN03]. Arborescences [CMN96]. Arc [BK97, HS95b, KKW98, DGM^{+}06, FMSN03, GH88, KMNF92, Leu84, Ram88]. Area [BFG^{+}96, CL02, Lar84]. Arising [FL01]. Arithmetic [LZ99, BRG89, Pan81a, Pan83]. arrangement [SRB06, GS84]. Arrangements [EGT96, Lib04, Yao85]. Array [BS95a, Mel81, MS05]. Arrays [KL98, SS00b, Sad03]. Arrivals [KPP98]. article [Ano02i, EIT^{+}93]. ary [Etz86, Pem05a, She95, Szp88, vB00]. ASP [CPR08, CF09]. Aspect [DGK01]. Aspects [ADL^{+}94, BBDS00, RS86]. assembly [Bru08]. Assessing [Fro00]. Assignment [FKN99, JKNS00, ET03, RDL84]. Assignments [ANR95]. Asteroidal [KKM99]. Asymmetric [PV98, Blä04, Szp88]. Asymptotic [Fle99, Xu00, vV96, CHT02]. Asymptotically [BG99]. asymptotics [FJ02]. Asynchronous [CZ95]. ATSP [ZLL09]. attention [PRZP09]. auctions [VGC^{+}08]. Augmentation [CT99, KT93, PSK97, PB99, CW02, sH02a, Nag05, Sor88b]. Augmenting [BK00, Kan96, NI99]. Author [Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano93a, Ano93b, Ano94a, Ano94b, Ano95a, Ano95b, Ano96a, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a]. Automata [ELW97, NMR01]. Automatically [NT90]. Average [Dia80, EH97, Gal00, HM91, Lue98, Mey03, Rüb97, Sub99, AM82, DM02]. Average-Case [Lue98, Mey03, AM82].

Back [GK99, MRFL08]. backtracking [ABW03, BW85]. bad [Kuc91]. Baked [DKK99]. Baked-Potato [DKK99]. Balanced [And99, DGK01, Epp97, GT97, AF84, Geo04]. Balancing [AKP^{+}97, BCK00, Wes00, HLLT06]. balls [Aur88]. Bandwidth [Gup01, KKM99, GS84]. barriers [LS87b]. Based [AMO99, BC95, BSV93, Rou01, AJS08, FL86, JMVW02, LM90, SZ05a, WG08]. Bases [Har93, KGW99, Bla92]. basis [LKPC05, Sch88b]. Batched [EO85]. Bayesian [Wil08]. BDD [Kla99]. Be [KOPS00]. Behavior [Csi93, Xu00, vV96, Sch88a]. bend [RCS86]. benzenoids [BCH02]. Bernoulli [Fil92]. Best [BFF96, Hva96, KRS98, KR95b, CG92, HNS86]. Best-Possible [KR95b]. Better [AGM87, AS91, CFFK98, Fer98, KPT96, HNS86, Vak03]. between [BT83, Ede85, LN02, SZ90]. bias [BP04]. Biased [KRS98, You00]. Bichromatic [CDBPL^{+}09]. Biclique [Hoc98]. bicoloring [KT02]. Biconnected [Thu97]. biconnectivity [sH02a]. Bicriteria [MRS^{+}98]. bijective [FZ82]. bilinear [Fei81, Fei83a, Fei83b, Pan81b, Pan83, SL84]. Bin [Csi93, KRS98, AJKL84, BC82, Bak85, CFGK91, CW02, Hof84, HK86, RBLL89]. bin-packing [BC82, Bak85, HK86]. Binary [CF96, Dor98, DG01, HY88, HM96, HM02a, LZ99, LVR93, PR85, Tho96, AKKL02, Bri86, EK04b, HHHW86, HHL88, HW84, Hua90, Ili82, Kno82, Lar87, Luc87, PY92, PM85, RP90]. Biology [Blu00, JKL00]. Bipartite [BJY01, BNK98, DKST01, FN01, KLST01, KR00, KK95, PV93, WPMK86, AH87, HLLT06, KMNF92]. Birkhoff [LLL86]. Bit [Tho02, Pan81a]. bit-complexity [Pan81a]. Bit-wise [Tho02]. block [MS87]. block-interchange [MS87]. blocking [Vis92a]. Board [Ano80b, Ano81b, Ano82c, Ano83c, Ano84b, Ano85b, Ano86b, Ano87b, Ano88b, Ano89b, Ano90b, Ano91b, Ano92b, Ano02e, Ano02f, Ano02g, Ano02h, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano09a, Ano09b, Ano09c]. body [FP89, HG88]. book [Woo02]. Boolean [Ano98e, Mak02, MS08, Tho02]. Bootstrapping [BT95b]. bottleneck [GT88]. Bound [DB97, IM01a, MT97, DLO03, Fre81, Pac87, Sch88a, Sed86]. Boundary [Pal97, BL92]. Bounded [FKK^{+}97, FBS97, FS94, HM02b, HN96, Lag96, Pis99, Wan94, BM05, CFR03, CT06, CW02, DW84, EFF04, FKL02, GHT84, HHL88, KM03, TSB05b]. Bounded-Degree [FKK^{+}97, CT06]. Bounding [BHP01, MLP99]. Bounds [BS97a, BMR98, Bsh99, Gal00, GG98, Gus84, HRG00, NR00, PS97, Sie01, Wan97, AR05, ES03, FM02, Kla92, KS06, Lag80, Mey03, NNS85, SY82, WS90]. bowling [BR91]. Box [BHP01, He01, RiNN00]. Box-Rectangular [RiNN00]. boxes [CDBPL^{+}09]. brackets [SR03]. braids [MS90b, PR91]. Branch [Har97]. Branches [BT95a, BKL^{+}04]. branching [JN08]. branchings [GKR^{+}91]. Branchwidth [BT99]. Bridge [CL02]. Broadband [FST97]. Broadcast [BZ93, Liu02, CH03, CG89, EK04a, GM03, Lib04]. Broadcasting [BDP97, CR06, KKP01, CGR02, KKW06]. Brooks [GP00]. Brownian [CL02]. Bruijn [Etz86, Hua90, Ral81]. bucket [Dev85, HHL^{+}02, Kno82]. bucketing [AM82]. Buffer [BKVV00]. Building [Fle99, MR89, HM91]. Buses [IMK96, Raj98]. Byzantine [BDP97, Dol82, MW94].

C [AKM^{+}96]. cache [BDIW04]. cache-oblivious [BDIW04]. Caches [LL99]. Caching [CN00, KPR01, Knu85a]. Cactus [Fle99]. Calculating [MS90b, Bor85]. Call [GGK^{+}97]. calls [BNM04]. cancellation [Pem05b]. canonical [CG84]. canonization [Shi81]. Capacitated [GHKO03]. cardinalities [GS85]. Cares [ZSW94]. Carlo [KLM89]. carpool [Nao05]. Carry [Pip02]. Cascading [Sen95]. Case [BFF96, GG98, Lue98, Poo93, SA00, Wal98, Xu00, vV96, AM82, FM02, HM91, KLL88, Lag80, Mey03, Sch88a]. Cases [BN02, BN03]. Cauchy [CHT02]. Cellular [Fre95, JKNS00, NMR01]. Center [AEST93, HSG02, PV98, MZ86]. Centers [BKP93, FJ83]. Centralized [GM03]. Centroids [Gab95]. century [ABIR89]. Certain [Fei83a, HM89, Kla81]. Certificate [NH01]. Certificates [Thu97, Can81]. Certified [MS94c]. Chain [Czu96, Fle99, Gho96, MS94b, PW98]. Chaining [Rag93, Kno84, KdlT89, Lar84]. Chains [BN02, CDF01, DG00, BN03, BBBD89, Oli81, Zan91]. changes [BKN04]. Channel [FWW93]. characterization [BMZ90, SSZ86]. Characterizations [AF84]. Chaudhary [Ano02i]. Cheaply [CE01]. checking [HH81, MS08]. checkpointing [JZ90]. Choosing [EH97]. Chordal [BCD99, Gup01, KK95, Leu84]. Chordless [GLD94]. Chow [JPS01]. Chromatic [Jan00, Bod90, LG83]. CIOQ [KR06]. Circle [AS93, DIP02, KKW98, Spi94, DKR82, WPMK86]. Circuits [Ebe94]. Circular [BK97, HS95b, KKW98, DGM^{+}06, GH88, KMNF92, Leu84, Shi81, Lib04]. Circular-Arc [BK97, HS95b, KKW98, DGM^{+}06, KMNF92, Leu84]. circulations [AP86]. class [Bed07, DLZZ03]. classification [ABK92]. Clausal [ND07]. clause [JN08]. Client [ATUW01]. Clique [BK97, DKST01, Hoc98, IA83, Koc05, KT02]. cliques [KMNF92]. closed [MS90b]. Closest [DHKP97, Sto85]. Closest-Pair [DHKP97]. closure [BHS07, SSW84]. Cluster [Liu02]. Clustering [Dah00, AP03, DLSS95]. clusterings [CRW91, HN89]. co [PR91]. co-NP-complete [PR91]. coalescing [Kno84]. Coarsening [MTT99]. Code [KR93, vB00, Kli82, Sav89]. Codes [BGLR02, HNS97, PR85, RT84]. coding [Knu85b]. coefficients [Can81, Lot92]. Cograph [He93]. cographs [AP90]. Coins [AK97]. collecting [BYES05]. Collection [LMY98, Pri99]. collision [KdlT89]. Color [BNK98, HNS86, KN88, Tur88]. Colorable [BS95b, Sub99, HNZ02, Tur88]. Colored [BK93]. Coloring [BS95b, BK87, CD95, CDF01, Hal97b, HNZ02, KR00, KNS01, Min89, PW97, Sub99, ZiNN96, ZSN96, ZSN97, BE05, BW85, CNS81, CY89, CN90, DLZZ03, FM02, KS87, Koc05, KS03, Kuc91, SU84, TW84, Vis92b, Zve06]. Colorings [GP00]. colors [KN88]. column [Joh81, Joh82a, Joh82b, Joh82c, Joh82d, Joh83a, Joh83b, Joh83c, Joh83d, Joh84a, Joh84b, Joh84c, Joh84d, Joh85a, Joh85b, Joh85c, Joh86a, Joh86b, Joh87a, Joh87b, Joh88, Joh90, Joh92b]. combinations [Rus88]. Combinatorial [Mon80, Old01, BMZ90, GM08, Kli80, Kul90, Lie82, NNS85, VGC^{+}08]. Combining [DGK01]. come [PM89]. commitment [BYK88]. Common [AT00, BG02, LZU01, BFCP^{+}05, Chu92]. Communication [AAG^{+}97, CV95, LVV96, WCT00, CH03, DF02, FFU04]. Commuting [AW99]. Compact [Cow01, CW04, EGP03, LLY03, Jer86, Pri83]. Compaction [Rag93, GKR^{+}91]. comparability [NP04]. Comparing [KLST01]. Comparison [BCT98, Rou01, AKKL02, Epp90, FG03, MRFL08]. Comparison-Based [Rou01]. Comparisons [ES95]. Compatibility [War94]. Competitive [AMRR05, BNL03, BR97, BKVV00, CL94, CN00, FKL^{+}91, MMS90, GHP05, ŠŽ05b]. Competitiveness [ANR95, AW05, Geo04]. Complete [BYR02, FGP01, CM87, DF86, Hås90, PR91, Ron90, Tan91, Tho89]. completeness [Joh81, Joh82a, Joh82b, Joh82c, Joh82d, Joh83a, Joh83b, Joh83c, Joh83d, Joh84a, Joh84b, Joh84c, Joh84d, Joh85a, Joh85b, Joh85c, Joh86a, Joh86b, Joh87a, Joh87b, Joh88, Joh90, Joh92b, LG83, MO85]. completion [Kim05, vSL05]. completions [KP03]. complex [DB09]. complexities [Pan83, Tun87]. Complexity [AMO99, BS95a, BCT98, Bsh99, DHK93, FK96, GGP^{+}01, GN95, JJ00, LVV96, MS94c, NH82, PS96, SA00, Sit05, Aig82, AM82, AFN04, Ale85, Alt88, AP86, BDRG04, Bor85, CHS89, DA85, Er85, FS86, Fre81, FV82, HR83, KL91, KLS89, KT02, Lag80, Mey03, MP81, Nos85, Pan81a]. component [CSS89]. Components [CL95, Thu97, FG03, GS85, IA83]. composition [Ili82]. Compositional [AW05]. compositions [Kli82]. Compressed [ABF97, ALS03, Sad03]. Compression [BGPW97, MLP99]. Computable [FRW94]. Computation [HLW94, Kri96, MN98, AM86, BLW87, BGY80, Fil92, HY88, Sch88c, Sel80, SV81, Sni86]. Computational [JKL00, Mat96, Tun87, CHS89]. Computations [Coh97]. compute [Ran83, SSW84]. Computer [GGPS98, Sto85]. Computers [Raj98]. Computing [AEST93, AS96a, Ata86, Bes03, But83, Che96, CT06, Ede85, ELW97, FMO^{+}07, Gho91, Gho96, Har97, HRG00, JPS01, JM95, KK97, LO87, Pri99, Ron94, Sch98, Shp00, TL96, ZP93, AS91, BS91a, BT83, BMX06, EA81, HT90, Pea81]. Concave [Sch98, Epp90, Wil88]. Concurrency [BC95]. Concurrency-Based [BC95]. Concurrent [CF96]. condition [Tan91]. Conditional [TPMRC05, CF09]. conditioned [CL02]. conditions [Hei05]. conference [BNM04]. Configurations [HS96]. Confluently [BT95b, FK03]. congestion [KL91]. congestion-1 [KL91]. congruence [Atk87]. congruential [Kra92]. conjecture [BN03, OTZ88, BN02]. Conjugates [KGW99]. conjunctive [Sch05b]. Connected [ADNP99, CL95, DN99, Fer98, KR95a, PB99, Raj98, CM88, CN89, FG03, IA83, Koc05, Nag05]. connectedness [BJMT92]. Connecting [sH00, Pal97]. Connection [LMRR94]. connections [BL92]. Connectivity [BJJ00, BK00, CT99, CZ95, Hen97, HRG00, IK94, KR96, NI99, NH01, PSK97, Tho99, II96, JMVW02, KT80, MS89, Nag05, SV82a]. Conquer [EGMS94, Ver94, Wan97, Mon80]. Consecutive [Hsu02, CY91]. Consensus [Asp93, AH90, Asp02, CBS04]. Constant [GIK02, IM01b, KL97, Wal98, Boy05, Bru08, GMM03, Yel92]. constant-size [Bru08]. constants [Kam88, Pan80, Ven87]. Constrained [CS99, HK93, WPYK06, Woo02]. Constraint [FN01, FP09, WG08]. Constraints [BC95, DL93b, ADK^{+}03, BNM04]. Constructing [BS01, CV95, GS85, GGPS98, Gro02, GW97, Ben80, Etz86, Lar87]. Construction [DR99, HHL88, KP98b, LOÖ01, Sko03, HHHW86]. Constructions [KMSY98]. Constructive [BK96, BD97, BCH02]. constructs [Fer87]. Consumption [CE01, CF09]. containing [KL85]. containment [BFM86]. continued [BBBD89, Bey83]. continuous [CHS89]. Contour [Ano81c, Güt84, LP80]. contours [LP81b]. Contraction [yHHsH^{+}00, ADKP89, LL88]. contractions [SRB06]. contradictions [Bed07]. Control [AMO99, GGK^{+}97, KPP98, GHP05]. controllable [Mas06]. controlled [JMP09]. Controller [BKSD97]. Convergence [AMO99, NR02]. Conversion [KK01]. Convex [AS96b, BS97b, Cha98, CKS^{+}98, DT93, Gho96, JMB96, LW95, Mir98, SV87, TL96, Ata86, BDRG04, DK85, Ede85, Epp90, Gho91, GY83, HG88, HS81, KL85, LS87a, MS90c, Ski91, Swa85]. Convolution [Pan80]. Convolutions [TL96]. Conway [Sch88a, War86]. Cooperative [GS04, KPR01]. Coordinated [MW94, AZ04]. coprimes [Ber05]. Core [PSY93, PtL96, SU97, MP83, MS80, WPYK06]. Correction [Che01a]. Corrigendum [Ano82b, Ano83b]. coset [All81]. Cost [ABK^{+}95, BT93, CHC99, FST97, Gol97, Hoo96, Jan00, KMN^{+}99, SZ90, Wu02]. Costs [BGLR02, Gal00, Epp90]. count [CM05]. count-min [CM05]. counterparts [BCd09]. Counting [BM97, FJM^{+}96, GJS95, MT97, PS95, MW92]. Cover [BY01, CKJ01, FN01, MS94b, Sla97, BYE81, NR03]. Covering [CR94, Kri97, Yeo99, GKS04, GHKO03, HM87, MS90c]. Covers [KVY94]. Crane [COKN06]. Criteria [Hoo96]. critical [Lov87]. Crossing [SS00a]. crosslingual [PCBC^{+}09]. CTL [GHCMS09]. Cubesort [CS92]. cubic [BST99, HLZ04]. Cuckoo [PR04]. Cumulative [Ano96b]. Curved [AVO93, CDV06]. Curves [GW00, SV87]. Cut [ACZ98, GT01, HO94, Raj95, Yeh01, He91b, HLZ04, FKL02]. Cut-Through [Raj95]. Cuts [Epp97, Fle99, GVY04]. cutsets [LL88, Ros82]. Cutwidth [TSB05b, TSB05a]. Cycle [BEMS95, DHK93, Tho93, Yeo99, BST99, CN89, LL88, Tsa03]. Cycles [FJM^{+}96, FR95, GL96, GL98, Hva96, Pem05a, Zhu95, Bri87, CPR03, CM88, Fri88, Ric86, Sor88a]. cyclic [Er85, Lib04, Heb98]. Cyclotomic [PS95].

d [Cha98, DGK01, AS96b, ALS03]. DAGs [AR05]. Data [BGH99, BGPW97, BS95a, BT95b, Fre97, FJMO95, GG97, Goo11, HM93, Kim05, LMY98, Rou01, CM05, CS92, DM02, FK03, FFV80, FJ83, GHPT03, KK06, MR92, Ove88]. Data-Oblivious [Goo11]. Data-Structural [BT95b]. database [Wol86]. Datalog [dLFM07]. David [EIT^{+}93]. DAWG [Blu87]. Deadline [DL93b]. Deadlines [GPS00, SB95, Gus84]. deadlocks [LP81a]. decaying [CS06]. decision [SY82]. decisions [Mai88]. Decomposable [BS80, ALS91, BLW87, EO85, Ove81]. Decomposition [BV99, Coh96, Dah00, DGM01, EGMS94, MS94a, PS96, Hua91a, IMH82]. decompositions [MT91]. Decreasing [Csi93, Bak85]. Decremental [Epp00, Tho99, BHS07]. Deficiency [HK93]. defined [Len87]. Degree [FKK^{+}97, FR94, HK93, KL00, Woo02, ABK92, CFR03, CT06, FKL02, GP02, MW90, Pem05b, TSB05b]. Dehn [DA85]. Delaunay [Rup95]. delay [AZ04, BNM04]. Delays [BKRS96, LVV96, DF02]. deleted [GS85]. Deletion [BYR02, Fuj99, Vit82, Bri86]. delivery [Vak03]. demand [BNL03, Wan05]. Demands [AIMF02]. Dense [AITT00, DHK93, FWW93, FR95]. density [AH87, Kla81]. depend [Kam88]. dependent [Fer87]. Depth [Bod93, Ebe94, Sni86, Tiw86]. Depth-First [Bod93]. Depth-size [Sni86]. Deques [BT95b]. Derandomization [Mat96]. Design [ARV98, HN96, MRS^{+}98]. Designing [FST97]. Despite [AH98]. Detecting [YW04, LP81a]. Detection [HKAA97]. Determination [Eng95]. determine [Kla81]. Determining [DC93, KLW96, DK85]. Deterministic [Cha98, CM96, GZ94, Hag95, HMP01, Han04, KRT94, Ron94, SCK97, Asp02, CM87]. Deviations [MH96]. Diagnosability [CCMS02]. diagnosis [HS84]. Diagrams [CKS^{+}98, Aur88]. Diameter [ZP93, AIKS91, HS97]. Diameters [CH88]. Dictionaries [Bre95, HMP01]. Dictionary [AC00, AKL^{+}00, Bre95, YY97, BDIW04]. Dictionary-Matching [Bre95]. difference [Gol82]. Different [CS99]. differential [CHT02]. differentiated [AMRR05, KM03]. Digital [LMY98, Rou01, HHL^{+}02]. Digraph [Yeo99]. Digraphs [BJY01, Coh96, GL96, GL98, Tho93, BJMT92, CFR03, FMSN03, SSZ86]. Dijkstra [Nos85]. Dimension [CM96, YM99, NPT92]. Dimensional [ABF97, AE97, DVS95, GG97, IM01a, IM01b, Rup95, AJKL84, BBK81, BC82, BZ88, Ste87, Vel86]. Dimensions [BHP01, LC92]. Diophantine [Gre88, Tun87]. Direct [Kno84, GNO89, KdlT89]. Direct-chaining [Kno84]. Directed [BPT96, CCyC^{+}99, ELW97, Fu97, FR95, GL98, HT97, HO94, Kle93, KPRT97, Mot98, PW98, BFCP^{+}05, Blu87, CW04, Fri88, JLR06, MS89]. disc [ÓY85]. disconnected [CT06]. discrete [NH82]. discs [Aur88]. disguised [Asp80]. Disjoint [CC00, SS00a, Wei97, Vel86]. Disjunctive [FK96]. Disk [BKVV00, AF04]. disks [KK06]. Dispatching [BDPS03]. Dispersion [CH01a]. Distance [BCD99, CDV06, CKS^{+}98, GPPR04, yHHsH^{+}00, BT83, LN02, SZ90]. Distance-Hereditary [yHHsH^{+}00]. distances [Ede85, Sel80]. Distant [Ten99]. Distributed [AR93, ABF98, BR97, GP00, Hal97a, JKNS00, KK97, KP98b, KP99, PS96, SZ99, Thu97, AW05, BYK88, BMZ90, DKR82, FS86, HT90, JZ90, NT90, Pac87, SU84]. Distribution [CHC99, CH03, Dev85, HM89]. distributively [RV92]. Divide [EGMS94, Ver94, Wan97, Mon80]. Divide-and-Conquer [EGMS94, Ver94, Wan97, Mon80]. division [Fei81]. division-free [Fei81]. divisions [Fei83b]. divisors [Rón88b]. DLV [MRFL08]. DM [DF86]. documents [DB09]. Does [ES95, Kam88]. Domains [BMR98, RS03]. dominance [ADK^{+}03, dBCO92]. dominances [GNO89]. Dominating [KP98b, EFF04]. Domination [DIP02, FK85, AGK04]. dominator [Chu92]. Domino [BE97]. Don’t [ZSW94]. Double [HSH^{+}98, CH88, Yao85]. Doubling [Coh97]. Doubly [BSV93, VO93]. Doubly-Linked [VO93]. doughnut [SIN92]. Drawing [GW00]. Drawings [He01, RiNN00, RNG04]. Dual [KVY94, AJKL84, CFGK91, JMVW02]. Dualization [FK96]. Dually [BCD99]. duals [FHS03]. duration [ABW03]. Dyadic [CJM02]. Dynamic [AT00, BG02, BJM94, Blu00, BGLR02, DLSS95, EIT^{+}93, Epp94, Epp97, Fer97, FG99, FLNU03, FMSN00, GT97, HM02a, IA87, KL98, Knu85b, RT97, Tho99, Win01, Awe90, BDIW04, BS80, EO85, EIT^{+}92, FFV80, FV82, FMSN03, Geo04, GS84, LS91, Wac89, vO09]. Dynamical [BDV02]. Dynamically [Fre97]. Dynamization [GJS95, Ove81, VW82].

early [Wol86]. Easy [ALS91, Tur88]. EDF [LNT04]. Edge [BJJ00, BK00, CD95, DDSV93, Fer98, HT97, Hen97, KR00, MUI01, NI99, NH01, Wei97, ZiNN96, ZSN96, ZSN97, CY89, CN90, HNS86, KS87, MS89, MO85, SRB06]. Edge-Coloring [KR00, ZiNN96, ZSN96, ZSN97, CY89, CN90]. Edge-Connectivity [BJJ00, NI99]. Edge-Disjoint [Wei97]. Edges [BJJ00, Mal94b]. editing [SZ90]. Editor [Fre94, Joh92a, Ram95a, Tar98]. Editorial [Ano80b, Ano81b, Ano82c, Ano83c, Ano84b, Ano85b, Ano86b, Ano87b, Ano88b, Ano89b, Ano90b, Ano91b, Ano92b, Ano02e, Ano02f, Ano02g, Ano02h, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano09a, Ano09b, Ano09c]. effect [JN08]. Effective [BCd09]. Effectively [FRW94]. effects [Awe90]. Efficiency [AAP86, JN08]. Efficient [ABK^{+}95, ALV92, Asp93, Att00, BKSD97, BDMR01, BT83, BBDL01, BK96, BPK00, BFS97, CC00, CB01, Che98c, Che98b, CJ91, Coh96, DGM01, DM85, FN01, FFU04, FR95, GT84, GGK^{+}97, Gol92, Gol97, He91a, HM96, HM02a, IWJ93, IZ94, KS87, Lag96, LKPC05, Len87, LOÖ01, MN98, MRR01, NH01, Ove88, PV93, PtL96, Tho96, Vil00, Wan00, WLK04, WPYK06, ADK^{+}03, AK92, Irv85, JLR06, JZC04, LS87b, NN90, NR03, Pea81, Sch88b, Sem84, Tiw86, Wac89]. Efficiently [BHP01, dFKKR00, LS88]. Elastic [IL03]. Election [AS97, NMR01]. Element [BYR02, JMVW02]. Element-Weighted [BYR02]. Elementary [IM01a, Pip02, Bey83]. Elements [Ron94, Kan85, MP81]. Eliminating [KP01]. Elimination [BC99, BGHK95]. Ellipsoids [LR99]. Embedding [Fu97, GN95, HM96, HM02b, Tam96, KL91]. Embeddings [Fre95, HM02a, MO85, Woo02]. empty [CND03]. Enclosing [BFG^{+}96, OAMB86]. enclosure [LP82]. enclosures [GNO89]. encodings [Gol82]. End [AAG^{+}97, AZ04]. End-to-End [AAG^{+}97, AZ04]. Energy [Ten99]. Ensemble [FG93]. Entire [NI99]. entities [Wol86]. Enumeration [BD97, BCH02, KLM89]. Envelopes [EGT96]. environment [Asp02]. episodic [HMI09]. Equal [JJ00]. equality [Yel92]. equation [Gre88]. Equations [AEH01, IWJ93, Vil00, BGY80, CHT02, Tun87]. EREW [CL95, HZ01b]. Erratum [Ano81c, Ano02i, BN03, EIT^{+}93, Ano99d]. Error [AKL^{+}00, HLW94, YY97]. essential [MP81]. essentially [Ber05]. Establishing [BNM04]. estimates [Pea81]. Estimating [Dra05, GLLF05]. Euclidean [BDV02, FRW94, MZ86]. Euler [CHT02, HSH^{+}98]. Evaluation [FL86, BGPGS06, GM08, Loz83, MRFL08, Pan80, VGC^{+}08]. evaluation-interpolation [Pan80]. Even [GL98, KLST01, Rüb97, Tho93]. Evidential [KL94]. Evolutionary [KLW96, LOÖ01, War94, BKL^{+}04, Sel80]. Exact [FN01, Knu96, NU02, SA00, Wah04, AF04, CGM92, COKN06]. exchange [CSS89]. Exhaustive [HH81, McK98]. existence [Tan91]. Expander [FM99]. expanders [AGM87]. Expected [ABW03, KS98, MS98, CO07, Dev85, DF89, KT80, Nos85, TAB86]. expected-time [KT80, TAB86]. Experimental [GM08, GT01, WG08, JN08]. Experimenting [CPR08]. Experiments [KLW96]. Explanation [Sub96]. Exploiting [SH02b]. Exploration [Gal00, DFKP04]. Exploring [PP99]. explosion [GM08]. exponential [AFN04]. exponentially [FS86]. Exponentiation [Gor98, LKPC05]. Expression [WMM95, dLFM07]. Expressions [LZ99, BRG89]. extend [dLFM07]. Extended [BJY01, HW00b]. External [Sib06, ABT04]. external-memory [ABT04]. extrema [DKR82]. extremal [Lie82]. extreme [Ede85].

F [EIT^{+}93, Wac89]. F. [Wac89]. facet [Swa85]. facilities [TPMRC05]. Facility [GK99, KPR00, GS04, GMM03]. Factor [BDS93, Ans85, EGP03, GMM03]. Factorable [PS95]. factorial [CG84]. factorials [Bor85]. Factoring [Ber05, Rón88a, HM89, Hua91b, Sch82]. Factorization [FGP01, Hua91a]. Factorizing [Duv83]. Factors [HK93, GP02, Wei84]. Failures [BDP97, BKRS96]. Fairness [AAN^{+}98, Nao05]. Families [JJ00, Dra05]. Family [Ind01, GT84, Kli80]. Fast [AM86, Ano98e, AH90, Asp02, BGLR99, BGY80, CRS^{+}00, CT99, CY89, CGR02, Czu96, DK95, Gor98, GP00, GNO89, Hag95, HM87, KP98b, LV89, Leu84, Lot92, Pri83, SR03, SZ90, Shi81, Sor94, Sor88a, Sor88b, ABI86, CL91, Dia80, IMH82, KN88, LP81a, MR89, Tsa03]. Faster [AO00, ALP04, Fil92, GZ94, HO94, HHHW86, yHHsH^{+}00, KLST01, KRT94, ST99, sH02a]. Fastest [Col94]. Fat [OvdS96, Cha03]. Fault [BYK88, KKP01, KP99, GMM03, NT90, SHOS90]. Fault-Local [KP99]. fault-tolerance [NT90]. Fault-Tolerant [KKP01, GMM03]. feasible [FP89]. feedback [Awe90, Ram88]. FFD [Xu00]. FFT [SZ05a]. FFT-based [SZ05a]. FGK [MLP99]. field [Pan80]. Fields [FGP01, HR99, ABK92, GP02, Hua91a, Hua91b, Rón88a, SL84]. FIFO [And04, BC82]. Fill [KKW98]. Fill-in [KKW98]. Filtering [BPK00]. find [HS81]. Finding [AIKS91, AT00, Ano81c, AITT00, ADNP99, BRS97, BSV93, BT93, CFFK98, CC00, CDHX03, CM88, CL95, DN99, FP89, FKK^{+}97, FJ83, FR95, GY83, HZ01b, HO94, Hav00, HS91a, IA83, Kan90, KR95a, KL85, KS98, LW81, LP80, Lov87, LS97, MS89, NW00, PtL96, Ram88, Ric86, SV81, SU97, Swa85, Ukk85, Wei84, Yeo99, dFKKR00, BST99, BDRG04, BL92, CH03, DKR82, Fri88, HM85, HR83, Kan85, Len87, LG83, LL88, ML84, MT91, MP83, OAMB86, RV92, Sor88a, SR06, Vis04, Wah04]. findpath [BZ88]. Finite [FGP01, HR99, ABW03, ABK92, BT83, GP02, Hua91a, Hua91b, Rón88a, SL84]. First [Bod93, Csi93, Hva96, Bak85, CG92, PM89, Tiw86]. First-Fit [Csi93, Bak85]. Fit [Csi93, KRS98, vV96, Bak85, Hof84]. Fitting [HIL03]. Fixed [CM96, Kno82, LMRR94, Raj98, Boy05, EFF04, MSV06, NR03, NPT92, TSB05a]. Fixed-bucket [Kno82]. Fixed-Connection [LMRR94]. fixed-parameter [NR03]. Floats [Tho00]. floor [LLY03]. floor-planning [LLY03]. Flow [AMO99, AIMF02, BT93, DL93a, DL93b, FKK^{+}97, Gol97, KRT94, LMY98, Old01, Ram97, KS06, Ram88, SV82b, Vis92a]. Flow-Control [AMO99]. Flows [Gab95, HS83]. Fly [CE01]. forbid [Vis83]. Forest [EIT^{+}93, KR93, EIT^{+}92]. Forests [HZ01b, Tho96, Len87]. Foreword [Ano99e, Fre94, Kar00, Koz94, Mor01, Pru03, Ram95a, Shm04, Tar98, Joh92a]. form [Bed07, Ran83, Sch05b]. formal [Bey83, JMP09]. Formed [Wal98]. Forms [FK96, JPS01, Fei81, Fei83a, Fei83b, Ili82, Lag80, SL84, Tol82, Tre09]. formula [FZ82]. Formulae [DB97, Val84]. formulas [Sch05b, SH02b]. Forthcoming [Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02j, Ano02k, Ano02l, Ano02q, Ano84c, Ano84d, Ano84e, Ano84f, Ano85c, Ano85d, Ano85e, Ano85f, Ano86c, Ano86d, Ano86e, Ano86f, Ano87c, Ano87d, Ano87e, Ano87f, Ano88c, Ano88d, Ano88e, Ano88f, Ano89c, Ano89d, Ano89e, Ano89f, Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano92c, Ano92d, Ano92e, Ano92f, Ano02m, Ano02n, Ano02o]. forthcoming [Ano02p, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano07e]. Forward [LW95, Tsa03]. Forward-Convex [LW95]. Fountains [MS02a]. Four [sH00, Hei05]. Four-Connecting [sH00]. Fourier [Dia80]. fraction [Bey83]. Fractional [Sen95]. fractions [BBBD89]. Fragments [BG02]. Framework [GW00, ET03, JMP09, MOPV07]. France [Ano81c]. Free [BGLR02, GL98, KKM99, McK98, MT97, Che98c, Fei81, GP02, Wil81]. Frequency [JKNS00]. fringe [PM85]. Frontsize [BGHK95]. Fuel [CE01]. Full [BKSD97]. Fullerenes [BD97]. Fully [FMSN00, FMSN03]. Function [Bur96, Hoo96, HM85, Val84]. functionalities [Sad03]. Functions [Ano98e, Bsh99, CKS^{+}98, CLRW97, Ind01, KRS98, TL96, BP83, Loz83, Mak02, RW84]. Further [CKJ01, GJS95]. fusenes [BCH02]. future [Tre09]. fuzzy [Bed07].

Game [Alt93, Zha98, Alt88, LM90]. games [GS04]. Gamma [Har97]. Garsia [Kin88]. Gauss’ [Val91, KS96, RT84]. Gaussian [BC99]. GCD [Sor94, WTM05]. GCDs [Sch88c]. Gene [GGP^{+}01]. General [And99, ABF98, BJM94, DB97, KPP98, SM02, Ver94, BP83, CG92, CDBPL^{+}09, ET03, JP02, RV89, dBCO92]. generalised [Er85]. Generalization [RR96, IMH82, LLL86]. Generalized [BJL^{+}90, GJS95, KS96, Old01, Win86, HR83, KM86, KKW06, QS02, Hua91b]. generals [Dol82]. Generate [CRS^{+}00, PW98, SR03]. Generating [All81, DB09, FJM^{+}96, Gav87, HJW05, KL97, KP94, Kul90, RP90, RSW92, SR03, Squ98, Wor84, FS86, FJ83, Leu84, Sem84]. Generation [KMNF92, KL00, LVR93, McK98, Rup95, Wal98, Boy05, Hua90, KV83, MW90, Pem05a, Rus88]. generator [EG81]. Generators [ARV98, BP04, Kra92]. Generic [PW98]. Generosity [CL94]. genomes [LN02]. Genus [Mal94a, EFF04, GHT84, Tho89]. Geodesic [GT97, MN98]. Geometric [AST94, AF04, CRW91, DPPS01, HMR^{+}98, BYES05, LB92, PV84]. geometrical [Atk87]. Geometry [Mat96]. Get [PW98]. Giuseppe [EIT^{+}93]. Given [Yeo99, HW92, KL85]. Good [ARV98, MW92, FJ83, Tam85]. Gosper [Pem05b]. Gossip [LN02]. Gossiping [MS02b, CGR02, KKW06]. gradually [II96]. Graph [BJJ00, CZ95, CDF01, Dah00, ES98, EIT^{+}93, FL01, Fre95, FM99, Fu97, Gal00, GR98, GKS95, Hal97b, HO94, HK93, sH00, IK94, KT93, NRT02, Pri99, PW98, RS86, SRB06, Squ98, Tam96, AFN04, ADK^{+}03, ABT04, AJS08, BW85, BS91b, Blu87, Bod90, CO07, CDN89, CG84, Dra05, EIT^{+}92, Gav87, IA83, II96, KN88, KMNF92, Luc87, Nag05, NN90, SU84, Tho89, Vis92b, WPMK86, Zve06]. graph-based [AJS08]. Graphic [Gab98]. Graphs [BPT96, BNK98, BNHK^{+}00, BK97, BS95b, BK93, BK96, BT99, BCD99, Bur96, CD95, Che98c, Che01b, DHK93, DIP02, DCD96, DPPS01, DDSV93, DG01, EGMS94, FBS97, FS94, FJM^{+}96, FR95, GLD94, GW00, Gup01, HK02, HT97, He01, HM02b, HW00a, yHHsH^{+}00, HS95b, HMR^{+}98, Hva96, IZ94, Kan96, KR00, Kle93, KK95, KKW98, KKM99, Lag96, LW95, Mal94b, Mal94a, Mot98, NW00, PP99, PW97, PB99, RiNN00, Ram97, Ram95b, Sch98, Spi94, Sub99, SM02, Wei97, AP90, AF04, ALS91, BM05, BST99, BFCP^{+}05, BLW87, BF02, BK87, CFR03, CPR03, CG92, CT06, CDV06, CM88, CNS81, CN89, CY89, CN90, DW83, DW84, DGM^{+}06, EFF04, EC90, FK85, FKL02, Fri88, GVY04, GPPR04, GHT84, Gol92, GH88, HNZ02, HLZ04, HLLT06, HY88, He91a, JLR06]. graphs [Koc05, KT02, LLZ05, Len87, Leu84, LG83, MS89, MT91, MW90, NNS85, NP04, Pla84, PC91, RV89, RNG04, Ram88, Ric86, RT84, Sor88b, ŠŽ05b, Tsa03, TW84, Tur88, Vis04, Wor84, ZG91]. gray [PR85, HNS97, Kli82, KR93, Sav89, vB00]. Gray-Code [vB00]. Greed [Win01]. Greedily [AITT00]. Greedy [BGPW97, CST96, CH01b, Dor98, DG01, FKN99, GK99, KL98, LK98, MP93, SCGD94, Sla97, Bla92, Kuc91]. Grid [Mir98, Pal97, BL92, BR91, LS87a, MO85, Ove88]. Grids [CC00]. Group [GKR00, BP04]. Groups [KGW99, All81, Bla92, But83, DA85, Jer86]. Growing [BT95a, APP98, CDHX03]. grows [Pit87]. growth [BM05]. growth-rate [BM05]. guarantee [LNT04]. Guaranteed [MR99]. Guessing [DLMM05]. guide [Joh81, Joh82a, Joh82b, Joh82c, Joh83a, Joh83b, Joh83c, Joh83d, Joh84a, Joh84b, Joh84c, Joh84d, Joh85a, Joh85b, Joh85c, Joh86a, Joh86b, Joh87a, Joh87b, Joh88, Joh90, Joh92b]. gulde [Joh82d].

half [Ski91]. half-planes [Ski91]. Hamilton [Fri88, FJM^{+}96]. Hamiltonian [BEMS95, BJMT92, BST99, CN89, DHK93, FR95, Hav00, LW95, Luc87, RT84, Sor88a, Vis04]. Hamiltonian-connectedness [BJMT92]. Hanoi [Er85, Min89]. Hao [Fle99]. haplotype [DR08]. Hard [CR94, GPS00, HMR^{+}98, CO07, DF89, EG81]. harder [CBS04]. Harmonic [vV96]. Hash [Ind01, KdlT89, Szy85]. hashing [Lar84, MV91, PR04, PM89, Vit82, Yao85]. Hasten [SSO01]. Heap [DR99, HM91]. heaps [Elm04, MR89]. Heapsort [BFF96, LP93, SS93]. height [DW84, KS02]. Helpers [SSO01]. Helps [CL94]. Hereditary [yHHsH^{+}00]. Hermite [Ebe94]. Heterogeneous [Liu02]. Heuristic [DCD96, KPR00, LC92, LMY98, Pla84, Xu00, BS85a, COKN06, PH87, PM85]. Heuristics [FS96, KLS89, MRFL08]. hexagonal [ŠŽ05b]. Hidden [Pel96]. Hierarchical [AC98, Dah00, KPR01]. hierarchically [Len87]. High [IS96, AZ04]. High-Speed [IS96, AZ04]. Hilbert [HW00b]. histograms [MS02a]. History [War94]. hoc [FFU04]. Holes [GLD94]. homeomorphism [Chu87]. Homeomorphisms [GW97]. Homogeneous [BY01]. Homomorphisms [Zhu95]. homotopic [Bes03]. Hong [OTZ88]. hook [FZ82, SL06]. hook-length [FZ82]. hop [HL03]. Hot [ANS97]. Hot-Potato [ANS97]. Hough [BS91a]. Huffman [HJW05, Knu85b]. Hull [BS97b, Ata86, GY83, Swa85]. Hulls [DT93, SV87]. Hybrid [KMSY98, Kle93]. Hypercube [CSS89, HM96, KK97, KL91]. Hypercubes [HM02a, HM02b]. Hypergraph [LS97, LS99]. Hypergraphs [BV99, Kri97, KNS01, KT02, KS03, Wah04]. Hypertext [ALL00]. Hypothesis [Hua91b].

IC [Len84]. IC-layout [Len84]. Ideals [HNS97, KR93]. identical [Vak03]. Identification [Mig82, SHOS90]. identify [KT86]. Identifying [Ano98e, MP81]. Identity [KOPS00]. II [RS86, TSB05b]. image [CSS89]. images [ALS03]. Immersion [BGLR99]. Implementation [Gol97, yHHsH^{+}00, MRFL08, Vis83]. Imprecise [HLW94]. Improved [AT95, AW02, Aur88, ABNLP90, BHS07, BR94, CN90, FKL02, FG99, GK99, Gup01, GS84, KR96, KS06, PSK97, Ven87, WG97, CM05, He91b, LP82, MS87, ZLL09]. Improvement [CH01b]. Improvements [CKJ01, Sch82]. Improving [KMN^{+}99, SCGD94]. in-place [Mel81, Spr89]. inapproximability [EK04a]. Inclusion [Che98b]. Incomplete [LN02]. Increasing [FSO99, NT90]. Incremental [AISN91, Epp00, Hen97, RR96]. Independent [AC00, Che01b, DG00, HS95b, Ind01, KNS01, LS97, AF04, ABI86, CM88, HS91b, KMNF92, Leu84, Rob86]. Index [Ano93a, Ano93b, Ano94a, Ano94b, Ano95a, Ano95b, Ano96a, Ano96b, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02d, Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano02c, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Bod90, DLO03, LG83]. Indexed [AP98]. Indexing [AKL^{+}00, Fer97, FG99, Sad03, WLK04]. induced [CM88]. Inequality [ABK^{+}95, Len84, Pla84, ZLL09]. inference [DR08]. Inferring [War94]. Infimal [TL96]. Influence [LL99]. Inplace [ALS03]. Inputs [You00]. insertion [BS85b, HM91, MP89]. Instability [And04]. Instance [FS96]. instances [Asp80, DB09, DR08]. Instancy [AW99]. instantiation [CPR08]. Insufficiency [Hei05]. integer [HM89, LS88, MV91, Sch88c, WTM05]. Integers [PS95, Shp00, Tho00, Gre88, Lot92]. integral [Can81, Fer87, Lag80]. Intensional [vO09]. interchange [MS87, Rus88]. Interlaced [Heb98]. Internal [DC93]. Interpolation [Ebe94, HR99, Pan80]. intersecting [BL92, DE87]. Intersection [AVO93, GJS95, JMB96, CJ92, Fra81, GT84, Gro02, IA83, IA87, LW81, VW82]. intersections [Ata86]. Interval [Ano98f, ES03, GG98, HNS97, SH98, FGS92, Leu84]. intervals [CJ91]. invariant [MNU05]. Invariants [KGW99]. Inverse [AO00, AAL97, SFBY04]. Involution [JS93, SL06]. Irreducibility [FM05, HW00b]. Irreducible [Can81, CRS^{+}00, GP02, SR03]. Iso [Ano81c, Güt84, LP80, SSW84, Vel86]. Iso-Oriented [Ano81c, Güt84, LP80, Vel86]. iso-rectangles [SSW84]. Isomorph [McK98]. Isomorph-Free [McK98]. Isomorphism [BPT96, GR98, ST99, Bod90]. issue [Bra07]. Issues [Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02j, Ano02k, Ano02l, Ano02q, Ano84c, Ano84d, Ano84e, Ano84f, Ano85c, Ano85d, Ano85e, Ano85f, Ano86c, Ano86d, Ano86e, Ano86f, Ano87c, Ano87d, Ano87e, Ano87f, Ano88c, Ano88d, Ano88e, Ano88f, Ano89c, Ano89d, Ano89e, Ano89f, Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano92c, Ano92d, Ano92e, Ano92f, Ano02m, Ano02n, Ano02o]. issues [Ano02p, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano07e]. Items [Xu00, CS92].

J [BN03]. Jacobi [Shp00]. Jeffery [EIT^{+}93]. Job [BG99, JP02, KP03]. Jobs [Sga96]. joining [SWMV03]. Jordan [SV87]. Josephus [Llo83]. Jr [Ano81c]. jumps [HM85].

Kanellakis [AKM^{+}96]. Kayles [BK02]. kernel [ZC09]. Kernels [GLD94]. Key [AAG^{+}97]. Keys [dlTK97, Yao85]. Khachiyan [AS80]. kind [Zve06]. Knapsack [Lue98, Pis99, Joh80]. knots [MS90b]. known [Hei05].

L [GHCMS09]. Labeled [NRT02, Kli80]. labeling [CDV06, CSS89, GPPR04]. labels [IL03]. lacunary [FM05]. ladder [LS87b]. language [GHCMS09, PCBC^{+}09]. languages [dLFM07]. Large [HR99, MH96, AGK04, DR08, GP02]. large-scale [DR08]. Larger [GZ94]. Largest [CND03, Pit87]. Las-Vegas [KOPS00]. Last [PM89]. Last-come-first-served [PM89]. Lattice [NRP96, Sch88b]. laws [KS02]. Layered [Bur96, Ram95b, VW82]. Layout [DPPS01, Len84]. layouts [Sko03]. LCFS [SB95]. LCS [WCP93]. Leader [AS97, NMR01]. Leaf [LR98, NRT02]. Leaf-Labeled [NRT02]. Learning [PRZP09, BCd09, HMI09, JN08]. Least [BPK00, FST97, Wil88]. Least-Cost [FST97]. least-weight [Wil88]. Lemma [ADL^{+}94, Geo04]. Length [Bre95, FWW93, NU02, PtL96, Wan00, ZSW94, AISN91, Boy05, Dev85, EG81, FZ82, MO85, MP83]. Letter [BGLR02]. level [Bur81, HS84]. levels [Tam85]. Lexicographic [DK95]. lexicographical [Sem84]. lifting [Joh80]. Limit [KS02]. Line [AK98, ANR95, AKP^{+}97, BGPW97, BCK00, DC93, FKN99, GGK^{+}97, GG97, Hal97b, HS96, LW95, Lue98, Pal97, PW97, Ram95b, You00, Ata86, APP98, BNL03, BZ88, CJ91, HIL03, LS83, LS91, LC92, NN90, Pea81, RBLL89, Sga96, Tam96, VW82, WPMK86, YW04, vSL05]. line-triangle [BZ88]. Linear [Alt93, AEH01, BS95a, BLW87, BZ93, Bod93, BK93, Cha98, CM96, CG89, GW97, GS84, He01, KT80, LR98, LS97, Mat93, Moh93, Pis99, Pit87, SU97, Tho02, TL96, Vil00, Wan05, Wei97, ZSN96, AS80, BYE81, Ber05, BF84, CNS81, CN89, DF02, DLO03, DK85, EA81, Fer87, GM84, Gre88, Han04, LS88, Lot92, Mak02, Mas06, MS80, Pan83, RBLL89, Ros82, SRB06, Sko03, TSB05a, Thu97]. Linear-Time [CM96, BLW87, BYE81, CN89]. Linearity [LN97]. linearly [Fer87]. Lines [CKS^{+}98, DT93, EGT96, Meg85]. Link [BDP97, CGM^{+}95, Sch98]. Linked [VO93]. Lipski [Ano81c]. List [CN00, HRJ^{+}89]. Listing [vB00, Gol92]. Lists [AP98, SZ99, VO93, Col86, HS91b, Kno84]. little [DFKP04]. Load [AKP^{+}97, BCK00, Wes00, AMZ06, HLLT06]. Loading [BKN00, Wan05]. Local [CH01b, FCL00, KPR00, KP99, CL02, DR08, Nag05, RT87, ŠŽ05b]. locality [BDIW04]. locality-preserving [BDIW04]. locally [Koc05]. Locating [AIMF02, Wan00]. Location [BJM94, DVS95, GK99, KPR00, OvdS96, SA00, GS04, GMM03, TPMRC05]. locational [Mai88]. locked [LP81a]. locking [Wol87]. log [Har97]. log-Gamma [Har97]. Logarithmic [BSV93, BMR98, BDP97, Ebe94, EK04a, CM87]. logarithmically [Pit87]. logarithms [Sch05a]. LOGCFL [Wan94]. logging [JZ90]. logic [Tre09, Wil08]. logical [Pan83]. logics [Bed07, ND07]. LogP [FLNU03]. long [Vis04]. Longest [BG02, Dev85, AR05]. Look [MRFL08, YY97]. Look-back [MRFL08]. Look-Up [YY97]. Lookahead [SCGD94]. Lookahead-Search [SCGD94]. Loop [ZP93, CH88]. Loopless [KL00, vB00]. Lorenz [Mai88]. Loss [KM03, MLP99]. Loss-bounded [KM03]. Low [FKK^{+}97, EGP03, Ten99]. Low-Weight [FKK^{+}97]. Lower [BMR98, Bsh99, Gal00, IM01a, MT97, NNS85, PS97, SY82, DLO03, ES03, Mey03, Pac87, WS90]. Lowest [BFCP^{+}05]. LSFA’06 [Bra07]. Lyapunov [KRS98]. Lyndon [SR03].

Machine [DL93a, Hoo96, Gus84, Mas06, vSL05]. Machines [BCK00, CB01, CS99, BRG89, EF05, Sit05, Vak03]. Maintaining [CS06, Fre97, FMSN00, BHS07]. Maintenance [EIT^{+}92, EIT^{+}93, Epp00, HS96, CJ91]. majority [BYK88, Val84]. Makespan [CB01]. Making [FK03]. Management [BKVV00]. Many [ES95]. Map [Che01b, IL03]. mapping [HIL03]. mappings [FG03]. maps [AERW03, Pem05a]. Markov [CDF01, DG00, PW98]. MAST [FHS03]. Master [Ver94]. match [FV82]. Matching [ABK^{+}95, AR00, AERW03, AAL97, ABF97, AC00, ALL00, AAL^{+}00, AKL^{+}00, BBDL01, Bre95, BCT98, Col94, DHK93, Fu97, GKT01, GG97, GPV93, GZ94, KP93, Mut00, WMM95, ZSW94, ALV92, ALS03, ALP04, BJL^{+}90, LV89, MNU05, Pla84, Ron90, SZ05a, Tan91, WPMK86]. Matchings [KLST01, AH87, HLLT06, RV89]. mathematical [GM84]. Matrices [AS96a, Mot98, AF84]. Matrix [BFS97, CL99, Czu96, Ale85, AK92, IMH82, KKB88, Kla92, Wu91]. matroid [CGM92, Fra81, GT84]. Matroidal [Fuj99]. Max [GBH97, AHL06, SV82b, AC98, AW02, BDM03, HZ01a, HLZ04, FKL02]. Max-Cut [FKL02]. max-flow [SV82b]. MAX-MIN [BDM03]. MAX-SAT [AC98]. MaxCut [MR99]. Maximal [LS97, PV93, ABI86, Leu84]. Maximin [Wan97]. Maximization [Che98c, FL01, MSV06]. Maximizing [KP03]. Maximum [BS97a, BK97, BT93, CC00, EH97, GKT01, HLW94, Hoo96, KRT94, KNS01, LR98, NR00, Pal97, RV89, Yan94, BT83, Blä04, CH03, Gav87, GLLF05, IA83, KMNF92, MO85, Rob86, SV81]. Maximum-Clique [BK97]. MaxSat [MR99]. may [Fei83b]. Mean [DL93a, DL93b]. measure [Fre81, Mai88, vLW81]. media [BNL03]. media-on-demand [BNL03]. Median [Sie01]. Meet [AIMF02]. Memoriam [AKM^{+}96]. memories [LS83]. Memory [AS97, Att00, ABT04, AH90, DFKP04, Vis83]. memoryless [Ral81]. Mending [KP99]. Merge [Rüb97]. Mergesort [CHC99]. Mergesorts [CHC99]. Merging [CJM02, LOÖ01, BNL03, SS87, SV81, Spr89, TAB86]. Mesh [ABE99, IMK96, Raj95, Raj98, Rup95]. Meshes [IM01a, IM01b, MTT99, SCK97]. Message [Att00, JZ90]. Message-Passing [Att00]. Messages [Hal97a]. Method [Ron94, Ver94, LO87, ÓY85, Pea81, Wil91]. Methods [Gor98, PS95, SWMV03, Joh80]. Migration [CLRW97, KP01, Kim05]. Miller [BMX06]. Min [ACZ98, GBH97, Ind01, NRP96, AHL06, CM05, BDM03]. Min- [NRP96]. Min-Cut [ACZ98]. Min-Max [GBH97, AHL06]. Min-Wise [Ind01]. MinCut [GS84]. Minima [BMR98, AK92]. Minimal [Fei83b, FRW94, Gol82, KR95a, AISN91, BBH^{+}06, Fei81, Fei83a, FGS92, OAMB86, PR91, Tol82]. minimality [AAP86]. Minimization [Mir98, GS84]. Minimize [Hoo96, Kim05]. Minimizing [AZ04, CB01, DL93a, DL93b, HLW94, Zan91, vSL05, MO85]. Minimum [ABK^{+}95, BJY01, BNK98, BHP01, BFG^{+}96, BMR98, Bla92, BT93, BM97, CE01, CGM^{+}95, Cow01, DIP02, EIT^{+}93, Epp94, Epp97, Fer98, Fle99, FWW93, FSO99, FR94, Gol97, HO94, Har93, HL03, HK93, JM95, KKW98, KMN^{+}99, LK98, MUI01, Sch98, Sub99, AIKS91, AHL06, AH87, Ben80, EIT^{+}92, EG81, FG03, Len87, MR92, Nag05, Ram88, Ran83, SRB06, Wah04, Wu02]. Minimum-Cost [BT93, Gol97]. Minimum-Degree [FR94]. minimum-length [EG81]. Minimum-Volume [BHP01]. Minor [Bod93]. minors [RS86]. minsum [QS02, Sit05]. mismatches [ALP04, SZ05a]. mixed [Epp90, LS88, Sor88b, VGC^{+}08]. mixed-integer [LS88]. Mobile [BGH99]. mod [AEH01]. modal [ND07]. modality [AM86]. Model [BC99, Bsh99, MS08, BC82, Bru08, CH03, FLNU03, Pem05b, SV81]. Modeling [CF09]. models [Vis83, ZC09]. moderate [MW90]. Modifications [KL94]. Modular [BV99, DGM01, EGMS94, WTM05]. Modularity [GL96]. modulo [Shp00]. Molecular [Blu00, JKL00]. Monge [BGLR02, Sch98]. Monotone [AS96a, BFS97, FK96, AK92, HM85, Val84]. Monotonic [Ano98e]. Monotonicity [BS91b]. Monte [KLM89]. Most [BT99]. Moti [EIT^{+}93]. Motion [FG93, LS87b, ÓY85]. movement [MR92]. moving [Blu87, HRZ03]. moving-target [HRZ03]. MST [ABT04]. much [Blu87]. Multi [GG97, KP01, ABT04, DW04, VGC^{+}08]. Multi-Dimensional [GG97]. Multi-processor [KP01]. multi-resolution [DW04]. multi-unit [VGC^{+}08]. multi-way [ABT04]. multicast [GHP05, MSV06]. Multichannel [BPK00]. Multicoloring [BNHK^{+}00, HK02, ŠŽ05b]. Multicuts [CFR03, KPRT97]. Multidimensional [LR99, SS00b, DM02, Mon80]. MULTIFIT [FL86].

MULTIFIT-based [FL86]. Multigraphs [ZSN96, ZSN97, HNS86]. multihop [GM03]. Multilevel [CN00, SRB06]. Multiobjective [DCD96]. Multipartite [DKST01, Yeo99]. Multiple [Mut00, RT87, ZP93, Gus84]. Multiple-Loop [ZP93]. Multiplication [CL99, Ale85, Kam88]. Multiplying [RW84, ABK92]. multiprocessor [JP02, LS83]. Multiresolution [Heb98]. Multiset [KL97, KL00]. Multiterminal [HS83]. Multivariate [HR99, HHL^{+}02]. Multiway [GVY04, Yeh01, Hua87]. Mutually [Ten99].

naive [Gus84]. names [Wol87]. natural [PCBC^{+}09]. Navigation [BBFY94]. NC [HMR^{+}98]. NC-Approximation [HMR^{+}98]. near [HS81, LLZ05]. near-optimum [HS81]. near-unanimity [LLZ05]. Nearest [AS96a, AT00, BSV93, Car01, Chu92]. Nearly [KR95b, LS83]. necessary [Hei05, KT86, Tan91]. necklace [FP09]. Necklaces [CRS^{+}00, RSW92, SR03]. Neighbor [Car01, MN98, SWMV03]. neighbor-joining [SWMV03]. neighborhoods [DM03, dBGK^{+}05]. Neighbors [AS96a]. Nested [TL96]. Network [BKP93, BT93, FKK^{+}97, HN96, LMY98, MRS^{+}98, PS96, Wan00, Awe90, FFU04, GHPT03, HW92, IR85, WPYK06, HS97]. Network-Flow [FKK^{+}97, LMY98]. Networks [AIMF02, ACZ98, ABF98, CV95, ELW97, FST97, HSH^{+}98, IS96, JKNS00, KPP98, KK01, KK97, KKP01, LMRR94, ZP93, adHV99, ABNLP92, And04, AZ04, BS85a, CIM^{+}02, CH88, CGR02, CW04, CR06, GM03, HS83, Vis92a, Win86]. Neumann [LLL86]. Neuroevolution [HMI09]. Next [Hof84]. Next-Fit [Hof84]. NFD [HK86]. Nimbers [BK02]. Node [CT99, Fuj99, KR95b, GVY04]. Node-Deletion [Fuj99]. Node-Weighted [KR95b]. Nodes [KMN^{+}99, Pal97]. noisy [Asp02, BGPGS06]. non [BL92, CDV06, CG84, EF05, HG88, KK06]. non-factorial [CG84]. non-intersecting [BL92]. non-positively [CDV06]. non-preemptive [EF05]. non-rotating [HG88]. non-uniform [KK06]. Nonblocking [FST97]. nonconvex [HM87]. nonhomogeneous [BF84]. noninductive [Fer87]. Nonintersecting [Pal97]. nonnegative [Gre88]. nonparallel [II96]. Nonpreemptive [FG93, SB95]. nonquadratic [Tol82]. nonseparating [CM88]. Nonstar [BN02, BN03]. Norm [FRW94, AERW04]. Normal [FK96, Bed07, LKPC05, ND07, Sch05b, Tre09]. normalizers [But83, Kan90]. Note [Gho96, Ano03u, PSK97]. nother [BST99]. NP [CO07, DF86, DF89, EG81, Hås90, HMR^{+}98, Joh81, Joh82a, Joh82b, Joh82c, Joh82d, Joh83a, Joh83b, Joh83c, Joh83d, Joh84a, Joh84b, Joh84c, Joh84d, Joh85a, Joh85b, Joh85c, Joh86a, Joh86b, Joh87a, Joh87b, Joh88, Joh90, Joh92b, LG83, MO85, PR91, Ron90, Tho89]. NP- [HMR^{+}98]. NP-complete [DF86, Hås90, Ron90, Tho89]. NP-completeness [Joh81, Joh82a, Joh82b, Joh82c, Joh82d, Joh83a, Joh83b, Joh83c, Joh83d, Joh84a, Joh84b, Joh84c, Joh84d, Joh85a, Joh85b, Joh85c, Joh86a, Joh86b, Joh87a, Joh87b, Joh88, Joh90, Joh92b, MO85]. NP-hard [CO07, DF89, EG81]. Number [ARV98, Ano81c, Ano02i, CC00, CV01, EIT^{+}93, FRW94, KL93, Pal97, FG03, Hua91a, KT86, Pem05a, Pit87, Pri83, Ran83, Wei84]. numbering [CG84]. numbers [BF84, Ehr81, Fer87, Fil92, Mig82]. numeric [Boy05].

Objective [Wil08, QS02]. Objects [AW99, AVO93, BFG^{+}96, OvdS96, Cha03, DE87, Kul90, PY92]. Oblivious [Goo11, IM01a, IM01b, MS02b, BDIW04]. Observation [BKSD97]. Observations [CKJ01]. obstacle [BP04]. Obstacles [HSG02, MN98]. Obstructions [GLD94]. Octrees [DGK01]. Odd [GLD94, Rüb97]. Odd-Even [Rüb97]. Off [BJJ00, HS96, Lue98]. Off-Line [HS96, Lue98]. Offline [Epp94]. offs [Kri96, Sni86]. Old [HRG00]. Omitting [Blu00]. On-Line [AK98, ANR95, AKP^{+}97, BGPW97, BCK00, FKN99, GGK^{+}97, GG97, Hal97b, Lue98, PW97, Ram95b, You00, LS91, RBLL89, Sga96, Tam96, BNL03, LS83, LC92, Pea81, vSL05]. One [AKL^{+}00, FR94, YY97, AJKL84]. one-dimensional [AJKL84]. Ones [Hsu02, CY91]. ongoing [Joh81, Joh82a, Joh82b, Joh82c, Joh82d, Joh83a, Joh83b, Joh83c, Joh83d, Joh84a, Joh84b, Joh84c, Joh84d, Joh85a, Joh85b, Joh85c, Joh86a, Joh86b, Joh87a, Joh87b, Joh88, Joh90, Joh92b]. Online [BBFY94, Cha98, CV95, GPS00, JKNS00, KP93, Kla99, CW02, EF05, GHP05, KP03, Vis92b]. ontology [DB09]. Open [DL93a, JKL00, KAO99, QS02]. Operations [Tho02, FFV80]. Optical [BHNS00, KK01, CIM^{+}02]. optima [RT87]. Optimal [AS97, AR00, AT00, ABE99, ABF97, BSV93, BG99, BS97b, BZ93, Blu00, BGLR02, BS01, Car01, CHC99, Chu92, EF05, FBS97, FG03, FR94, GG97, HZ01b, HM96, HM02b, HKAA97, HW84, Hua87, JMB96, KMSY98, LP98, MS95, MTT99, Mut00, PY92, SZ99, SSW84, TAB86, Wu91, WCT00, AKKL02, Atk87, BLW87, CDHX03, DW04, ES81, Güt84, HM85, HHHW86, HHL88, Lar87, OAMB86, Sko03, SR06, Vel86, Yao85]. Optimality [Win01]. Optimally [Che96, Wan00]. Optimistic [AMO99, JZ90]. Optimization [AST94, CM96, Liu02, Mat93, GT88, Lie82]. Optimum [ADNP99, DN99, Jan00, Per84, HS81, Knu85a]. Order [SZ99, Kan85, Ove81, Sem84]. Order-Preserving [SZ99]. Ordered [Che98b, Che01a, Eng95, KL00, SH98, Duv83, EK04b, LS85]. ordering [BBH^{+}06, Koc05]. orderly [LLY03]. Orders [HNS97, Ste87]. Organizing [VO93, HS91b]. orientation [Sor88b]. Orientations [Squ98]. Oriented [Ano81c, Hav00, Zhu95, And04, Güt84, LP80, Vel86]. origin [CL02]. Orlin [Fle99]. orthogonal [DM02, IA87, PY92]. other [Alt88]. Outerplanar [DDSV93, Kan96, NW00, BF02, HS83, ZG91]. outforests [DF02]. Output [BS97b]. Output-Sensitive [BS97b]. Overlap [NRP96]. overload [LNT04].

P [Ano81c, ZC09]. P-stable [ZC09]. Pack [GI99]. Packet [BG99, MP93]. Packing [AE97, BDMR01, CPR03, CH01b, Csi93, KRS98, MS90c, AJKL84, BBK81, BC82, Bak85, Cha03, CFGK91, CW02, Hof84, HK86, JZC04, LC92, RBLL89]. Padé [BGY80]. Page [AK98, CLRW97, LW95]. Pagenumber [Mal94b, Mal94a, SS00a]. pages [Ano02i, EIT^{+}93]. Paging [ABF98, You00, FKL^{+}91]. Pair [DHKP97, Sto85]. pairing [BMX06]. Pairs [ACZ98, CZ01, Fre95, Hal97a, BHS07, Dra05]. palettes [HNZ02]. Pancake [HS97]. Papers [Ano84c, Ano84d, Ano84e, Ano84f, Ano85c, Ano85d, Ano85e, Ano85f, Ano86c, Ano86d, Ano86e, Ano86f, Ano87c, Ano87d, Ano87e, Ano87f, Ano88c, Ano88d, Ano88e, Ano88f, Ano89c, Ano89d, Ano89e, Ano89f, Ano90c, Ano90d, Ano90e, Ano90f, Ano91c, Ano91d, Ano91e, Ano91f, Ano92c, Ano92d, Ano92e, Ano92f, Ano96c, Ano96d, Ano96e, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano97l, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano98m, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p]. Papers [Ano02q, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano07e, Cla98]. Paradigm [AR93]. Parallel [AP90, Aig82, Alt93, BEMS95, BKVV00, BSV93, BMR98, BT93, BFS97, BKRS96, CD95, CY91, CS99, Coh96, Coh97, CZ95, CGPR97, DHK93, Dah00, DK95, DR99, FR95, GR98, GPV93, GZ94, GN95, Hal97b, He93, HSH^{+}98, yHHsH^{+}00, IWJ93, IK94, JM95, KLM99, KVY94, KS97, Kri96, Lag96, LB92, LS97, MS98, Mut00, NP04, PV93, PC91, Rag93, Ram97, Sga96, SH98, Tho97, Vil00, Wan00, ZSN96, ZSN97, ADKP89, ABI86, AK92, Ben80, BK87, BR91, CS92, GKR^{+}91, Goo89, HY88, He91a, JLR06, KN88, KS87, KK06, KLS89, LV89, LKPC05, Loz83, MV91, NN90, Sch88a, SHOS90, SV81, SV82a, SV82b, Sit05, Sni86, Sor88a, Sor88b, Tiw86, Vis92a, War86, Win86]. Parallelism [Kle93, CPR08]. parallelizable [MS03]. parameter [EFF04, NR03, TSB05a, ZG91]. Parameterized [AFN04, Elm04, AF04, JZC04, ZLL09]. Parameterizing [MR99]. parameters [Tun87]. Parametric [AST94, Csi93, FBS97, FS94, Gab98, FBS89, SFBY04]. Parenthesis [Wal98]. Paris [AKM^{+}96]. Parity [Dah00, AP90, PC91]. parsimony [DR08]. Partial [BY01, BK06, HK02, ZiNN96, Bod90, FV82, GKS04, Ste87, TSB05b]. Partially [Eng95]. Partite [BYR02]. Partition [GBH97, Jan00, HS81]. Partitioning [FL01, GIK02, MS95, Meg85, SM02, BP83, Joh80, MS05]. Partitions [NRP96, dFKKR00, Boy05, HS91a, PY92, Sav89]. Passing [Att00]. Passive [BHNS00]. Past [Tre09, DM85]. Past-future [Tre09]. Path [BPT96, BEMS95, CGM^{+}95, Coh97, DK95, GG98, Har93, Hav00, NU02, RR96, Sch98, Wan00, adHV99, CDHX03, DLZZ03, HG88, MS03, Nos85, TPMRC05]. Path-Algebra [DK95]. Path-Doubling [Coh97]. Paths [CC00, Coh96, CZ01, Fre95, FMSN00, GLD94, GT97, KS97, KS98, LW95, MP93, NW00, SS99, Tho00, Wei97, AISN91, BHS07, Bes03, Bri87, Dra05, FMO^{+}07, FP89, FJ83, FMSN03, HR83, Mey03, RV92, Sor88a, SR06, Vis04, Hal97a].

Pathwidth [BGHK95, BK96, BF02]. PATRICIA [KS02]. Pattern [AAL97, ALL00, AAL^{+}00, CL99, Col94, Fu97, GG97, Mut00, Sel80, Tho96, ALV92]. patterns [Ukk85]. Pedestrian [HS95a]. OR [CG92, DCD96, Hva96]. Tate [BMX06]. perfect [TW84]. Perfectly [PW98, BDPS03]. perfectly-periodic [BDPS03]. Performance [Fro00, SWMV03, LL99, LNT04, PS00, AR05, DM02, Pem05a, SU84]. perimeter [IL03]. Periodic [BJJ97, HW00a, BDPS03]. Permutation [IZ94, KGW99, SCK97, Tsa03, All81, Bla92, But83, FK85, Jer86, KT86, Mel81]. Permutations [KL97, KL00, RT84, Tsa03]. Permuting [PPR04]. Persistent [BT95b, FK03]. personnel [RDL84]. perspective [AJS08]. Petersen [BBDL01]. phylogenetic [CT06, SWMV03]. Phylogenies [GGPS98]. Piecewise [GW97, TL96, SV87]. Piecewise-Linear [TL96]. piecewise-smooth [SV87]. piercing [Cha03]. pipelined [BRG89]. place [Mel81, PPR04, Spr89]. Placement [ABE99, GGP^{+}01, KPR01, KK06]. Placing [APP98]. Plaisted [OTZ88]. Planar [AEST93, BS97b, CFFK98, DDSV93, DF86, Epp00, GL96, GR98, GT97, GW00, HK02, HS96, PB99, SA00, Tam96, Tho93, Wei97, Yeh01, AFN04, AERW03, ABT04, BJL^{+}90, BT83, BK87, CNS81, CN89, CY89, CN90, He91b, RNG04, Ric86, RT84, TW84]. Planarity [Moh93]. Plane [BDGW98, EIT^{+}93, He01, RiNN00, Bes03, CDV06, DM03, EIT^{+}92, IA83, Meg85, MS90c, dBCO92]. planes [Ski91]. Planning [FG93, LS87b, LLY03, ÓY85]. play [BR91]. Point [AEST93, ABE99, BHP01, BJM94, DLSS95, DVS95, OvdS96, PY86, SA00, CND03, EA81, FP89]. Pointer [AT00]. points [AIKS91, BL92, Cab07, HIL03, MW92, MSV06, WPMK86]. Poisson [KPP98]. policies [AMRR05, KR06, Zhu04]. Polygon [AS93, AAP86, BFM86, Che96, DC93, ES98, Gho96, EA81, Gho91, Goo89, GY83, HS81, KL85, MS90c]. polygonal [LS87b]. Polygons [AS96b, CR94, GW97, JMB96, AM86, Ede85, EC90, PY86, Ski91]. Polyhedra [Pel96, DK85]. Polyhedral [CKS^{+}98, LR99]. Polylines [GW00]. Polylogarithmic [GKR00]. Polymatroids [Gab98]. Polynomial [AAG^{+}97, Bod90, Cha03, DF94, DGM^{+}06, FGP01, HW00a, JP02, Kan85, KT88, KL93, KGW99, MW94, NW00, Sub99, WCT00, Yeo99, Zhu95, BKL^{+}04, BJMT92, BGPGS06, CGM92, CO07, DF89, FGS92, Kam88, Kan90, LLZ05, Lot92, MS90b, Pem05b, Sch88c, Wei84, Wu02]. Polynomial-Time [HW00a, NW00, Cha03, Kan85, KT88, LLZ05]. Polynomials [CRS^{+}00, GP02, HR99, Shp00, ABK92, Can81, FM05, Hua91a, Hua91b, Rón88a, SR03]. polytopes [BDRG04]. poor [Pem05a]. Poset [KR93, YM99]. Positive [Ano98e]. positively [CDV06]. Possible [KR95b, HNS86]. Posted [MSV06]. Potato [ANS97, DKK99]. Power [CHC99, Aur88]. Power-of- [CHC99]. Powers [KC98, NRT02]. pp [Ano81c]. Practical [ARV98, BS97b, DGM01]. PRAM [CL95, HZ01b]. Precedence [CS99, DW84]. Precedence-Constrained [CS99]. predicate [Tre09]. predict [Kra92]. prediction [LS91]. preemptive [BC82, EF05, QS02, Sit05]. Preface [BI03, OO09, PV02, SdGG09]. preferences [CF09]. Prefetching [BKVV00]. Prefix [BGLR02, BCT98, Sni86]. Prefix-Free [BGLR02]. Prefix-Matching [BCT98]. Prefixing [dlTK97]. Prefixing-Keys [dlTK97]. Preparata [Ano81c]. Preprocessing [Kle93, Tho96]. Prescribed [GL96, GS85, Gro02]. Presence [MN98, ZSW94, BCd09, HHL88]. Presented [KAO99, MS90b]. preserve [Vit82]. preserves [Bed07]. Preserving [BJJ00, SZ99, BDIW04]. price [MSV06]. Primal [KVY94, JMVW02]. primal-dual [JMVW02]. Prime [EGMS94, Kan85, Lar84, Pri83]. primes [Hua91a]. Principal [Har97, NRP96]. priority [BS85b, SIN92]. privacy [JMP09]. prize [BYES05]. prize-collecting [BYES05]. Probabilistic [BC99, CFGK91, HKAA97, Sch88c, BS91a, DLMM05, Hof84, Kuc91, Pac87, Pem05b, Wil08]. probabilities [HHL88]. Probability [LM90, HRJ^{+}89]. Probability-based [LM90]. probable [Pit87]. probe [Dev85]. Probing [Ski91, CY87, GM84, LB92, Pit87]. Problem [AO00, AK98, ATUW01, BPT96, BJY01, BY01, BR97, BS97a, BR94, BKN00, BK97, BCT98, CE01, Che01a, Czu96, DHK93, DHKP97, FKN99, Fer98, GKR00, HSG02, Jan00, KPT96, KOPS00, LZU01, LP98, LN97, NBHB00, NU02, PV98, PS00, RR96, SS99, Tho93, Yeh01, AMZ06, AF04, ABI86, ABIR89, Asp80, AJKL84, BYE81, BYES05, BW85, BZ88, CN89, Chu92, Chu87, DLMM05, DR08, EK04a, EFF04, Er85, EG81, FHS03, GP87, GHPT03, GMM03, GS84, He91b, HRZ03, HM87, Irv85, II96, JMVW02, KS06, LP82, LS88, Mas06, MZ86, Min89, Nao05, PV84, QS02, RDL84, SZ05a, Sch05b, Sto85, Tho89, WLK04, Wan05, WPYK06, Wil88, Win86, vLW81, AT95, CIM^{+}02, COKN06, GIK02, HL92, IM02, Llo83]. Problems [AS99, BYR02, BDMR01, vEB91a, vEB91b, CH01a, CGM^{+}95, CCyC^{+}99, CM96, Che98c, CS99, CM87, DLSS95, DKST01, DK95, DPPS01, Epp94, FL01, FS94, Fre95, Fuj99, GPV93, GKS95, Gui80a, Gui80b, Gui80c, Gui81a, Gui81b, Gui81c, Gui82a, Gui82b, Gui83a, Gui83b, Gui83c, Gui84a, Gui84b, Gui85, GJS95, GN95, HT97, HN96, Hoc98, HMR^{+}98, IMK96, JKL00, KR96, Khu97, Khu98, KAO99, KPR00, Lue98, MS94b, MRS^{+}98, NRP96, Old01, PSK97, Pis99, PB99, RT97, WCP93, AIKS91, Aig82, AFN04, AHL06, ALS91, BS80, CGM92, CO07, CFGK91, DA85, DF89, EO85, ET03, FM02, FBS89, FG03, FP09, GT84, GT88, GKS04, GM08, Gho91, Joh80, KS87, KT80, KLM89, Kla92, Lie82, MMS90, MS05, NNS85, Ove81, PV84, Ron90, Sor88b]. Procedures [HZ01a]. processing [Mas06]. Processor [BKRS96, Hag95, KOPS00, Vil00, KP01, KLL88]. processors [DF02, MS90a]. Product [Czu96, WCT00, BP04]. Product-Requirement [WCT00]. production [CF09]. products [KKB88]. profit [MSV06]. Programming [AEH01, BG02, Blu00, BGLR02, Cha98, AS80, GS84, LS91, LS88, Wac89, vO09]. Programs [HZ01a, KP98a, CPR08, FL06, GHCMS09, ZC09]. Projections [LS99]. Projective [Moh93, AP03]. proof [Ans85, Ave92, Bak85, FZ82, Geo04, Kin88, LLL86]. Propagation [Pip02]. Proper [He01]. Properties [BN02, Fuj99, BN03]. Property [BYR02, BGLR02, Hsu02, Sch98, CY91]. protocols [Awe90]. Provably [ARV98]. pruning [LM90]. pseudo [CGM92]. pseudo-polynomial [CGM92]. PSPACE [HMR^{+}98]. PSPACE-Hard [HMR^{+}98]. Publisher [Ano03u]. pure [DR08]. Pyramid [Sto85].

QBF [MRFL08]. QoS [Zhu04]. Quadrangle [ABK^{+}95]. quadratic [Fei83a, Ili82, Lag80, Ran83, Tol82]. Quadripartite [Che98a]. Quadtree [Heb98]. Quality [Rup95]. quantified [SH02b]. quantifying [JMP09]. quantile [Pea81]. quantization [DW04, Wu91]. quartet [SWMV03]. Quartets [BS01]. Quasi [LK98]. Quasi-Greedy [LK98]. quaternion [Rón88b]. Queries [AVO93, Mat93, KT86]. Query [SA00, Fre81]. Queue [CHC99, IM01b, AMRR05, BS85b, SIN92]. Queue-Mergesort [CHC99]. Queue-Size [IM01b]. queueing [Zhu04]. Quicksort [ES95, FJ02, MH96, NR02]. Quincunx [Heb98].

Radio [KKP01, ABNLP92, CGR02, CR06, EK04a, GM03]. Radius [JMB96]. rainbow [Min89]. RAM [Bsh99]. Random [ARV98, BS95b, CGM92, DT93, DPPS01, DG01, DB97, FS96, FJM^{+}96, Gué83, KRS98, PW97, PW98, SM02, You00, BS85b, DW83, DF89, Fri88, HHL^{+}02, Kul90, MP89, MW90, Pem05a, Sag80, SU84, Wor84]. randomization [RV89]. Randomized [ANS97, Asp93, BM05, BS97b, DHKP97, Fro00, GR98, Goo11, HZ01b, KS97, LMRR94, LS97, Sga96, Tho02, Vis92b, ABI86, AH90, CH03]. randomizing [MZ86]. Randomly [CDF01]. Randomness [Kri96, Vit82]. Range [BMR98, Epp97, NI99, OvdS96, BBH^{+}06, DM02, Fre81, LW81, Ove88]. ranges [vLW81]. Rank [DR99, Hås90, Wah04]. rank-3 [Wah04]. Ranking [GLW83, MUI01, CDN89]. Rapidly [AH98]. Rate [AMO99, BM05]. Rate-Based [AMO99]. Rates [NR02]. Ratio [BDGW98, DGK01, Fer98, GIK02, PS00, AGK04]. Ray [AS96b, GT97, HS95a, CJ92]. ray-shooting [CJ92]. RCRA [GM08]. Reachability [Kle93]. Real [KS98, Ehr81, Fer87, HIL03, Pan80]. Real-Valued [KS98]. Realistic [Bsh99]. Realization [Mot98]. Reasoning [KL94]. rebalancing [AMZ06]. Recognition [CL99, Dah00, ES98, He93, HSH^{+}98, Spi94, CY91, DGM^{+}06, Sel80]. Recognizing [Asp80, CST96, EC90, JJ00, Mak02]. Reconfigurable [BS95a, Raj98]. Reconstructing [BGPGS06]. reconstruction [NN90]. records [Pit87]. Recovery [JZ90]. rectangies [LP80]. rectangle [CND03, LP82, GIK02]. Rectangles [Ano81c, BFG^{+}96, BDMR01, Güt84, IA83, LW81, LP81b, SSW84, Vel86]. Rectangular [DVS95, He01, RiNN00, RNG04, SS00b, OW88, vLW81]. Rectilinear [BDGW98, MN98, VW82, GIK02]. Recurrence [IWJ93, Wan97]. Recurrences [Ver94, Mon80]. recursion [Alt88]. Recursive [BPK00]. Reducible [Ram97, Ram88]. Reduction [BEMS95, BS95a, HT97, KS96, Sch88b]. redundant [MP81]. references [HS91b]. Refined [Sch82]. Refinement [BDS93, Kla99, Rup95].

Refinements [BMX06]. Reflected [CL02]. regions [SR06]. Register [FCL00, KP98a]. Regular [FJM^{+}96, NW00, WMM95, dLFM07, CCMS02, LG83, Mak02, MW90, Wor84]. Regularity [ADL^{+}94]. reinforcement [HMI09]. rejection [EKK^{+}03]. Related [BCK00, CGM^{+}95, CB01, GPV93, MS94b, Poo93, AIKS91, EF05, Gho91, PV84, Sor88b]. Relations [SSO01, Fer87]. relatively [HNZ02]. Relaxed [CN00]. Release [DL93b, Gus84, Vak03]. relevant [Len84]. Reliable [BDP97, CV95, DHKP97]. Removal [Pel96]. reorganization [Szy85]. Repeated [BS85b, HM91]. repetitions [ML84]. Repetitive [Pel96]. replacement [BP04]. Replication [AK98]. Reporting [GJS95]. representable [Ave92]. representation [AJS08, BZ88, FGS92, Jer86, Vel86]. Representations [DF94, Fle99, Gab95, LKPC05]. representatives [All81]. Representing [Yel92]. required [Ran83]. Requirement [WCT00]. requirements [KKB88]. rerouting [CIM^{+}02]. reservation [BKN04]. residue [BGPGS06]. resolution [DW04, KdlT89, ND07]. Resource [CW02, CF09]. resources [APP98]. respect [Sko03]. Response [Wes00]. restarts [vSL05]. restricted [DLZZ03, Dra05, RS03]. restrictions [HL03]. Restructuring [EK04b]. Results [GJS95, Jan00, NH82, Szp88]. retraction [ÓY85]. Retrieval [LW95, DLO03, FV82, PY86]. Reversing [ABIR89, SL06]. Revisited [LP98, Sen95, Val91, Wil88]. reweighing [Geo04]. RH [HIL03]. Riemann [Hua91b]. rim [SL06]. rim-hook [SL06]. ring [AS91, Kam88, Lot92, MS90a, Wan05]. Rings [HKAA97, BGPGS06, DLZZ03, Min89]. RNA [LS91]. Robert [EIT^{+}93]. Roberto [EIT^{+}93]. robot [BZ88]. robotics [PRZP09]. Robust [Att00, BPK00, RS03, Ros82]. Room [BBFY94]. Roommates [IM02, Irv85]. root [Tiw86]. Rooted [BPT96, Fre97, Tho97]. roots [CT06]. Rosetta [vO09]. rotating [HG88]. Rotation [LZ99, Luc87]. Rotations [BJJ97, LVR93]. Round [MW94, ABNLP92]. Rounding [HZ01a, FL06]. roundtrip [CW04]. routability [LS87a]. Routing [Ano98f, ANS97, BG99, CIM^{+}02, Cow01, DKK99, FWW93, GG98, IS96, IMK96, IM01a, IM01b, KL98, KPP98, KM86, LMRR94, Raj95, SCK97, adHV99, AHL06, ABNLP90, CHS89, CDV06, CG89, CW04, EGP03, GHP05, SIN92, Wu02]. row [AK92]. RPR [FL06]. Rules [CHC99, PPR04]. Rumors [AH98]. Run [CS99, Fle99]. Running [AR93, Rüb97, Dia80, HM89].

Safety [NBHB00, LP81a]. salesman [BYES05, HRZ03, PV84]. Salesmen [FJMO95]. Same [Fle99]. Sample [PW98]. Samples [GZ94]. Sampling [Ten99, Tsa03]. Sandwich [GKS95]. sat [FS96, AC98, AW02, AJS08, DB97, HZ01a, JN08]. Satisfiability [DB97, NR00, Yan94, AS03, Asp80, Bru08, FM02, Sch05b, MS08]. satisfying [Pla84]. scale [DR08]. Scaled [AC00, WLK04]. Scaling [Gol97, ALV92]. SCG’98 [KAO99]. schedule [AZ04]. schedules [BDPS03]. Scheduling [AAN^{+}98, BC95, BG99, BKN00, CS99, DW84, GHPT03, GPS00, Hoo96, IR85, KP01, KPT96, KR06, LVV96, Liu02, MP93, SB95, Sga96, SH98, BRG89, DF02, EKK^{+}03, EF05, FL86, Gus84, JP02, KLL88, LS83, Lib04, Mas06, NH82, Sit05]. schema [JMVW02, dLFM07]. Scheme [DIP02, yHHsH^{+}00, WCT00, BKL^{+}04, Mas06, Wu02]. Schemes [Che98c, HMR^{+}98, Cha03, CDV06, EGP03, JP02]. Scholz [BN03, BN02]. Schur [RW84]. Scores [BRS97]. Search [Alt93, Bod93, DCD96, DK95, Dor98, DG01, FBS97, GG97, HM93, Hva96, KPR00, SCGD94, She95, ABW03, AKKL02, BS85a, CG92, DR08, DM02, HHL^{+}02, IA87, Lar87, LW81, MP89, Pit87, PM85, RT87]. Searching [AST94, BFS97, CF96, Col86, Ehr81, GJS95, LS85, OvdS96, Ste87, Alt88, BS80, BS91b, CJ92, Dev85, DE87, EO85, FJ83, Kla92, Ove88, Wu91]. Seat [BKN04]. secondary [LS91]. Secrets [DLMM05]. Security [HT90]. Segment [AEST93, DC93, IA87, VW82]. Segmentation [AS99]. Segments [LP81b, Pal97, YW04]. Selecting [BKL^{+}04, Sag80]. Selection [AS96a, DR99, Raj98, DW83, ES03, RH84, Sib06, WG08, Wil81]. Selective [Coh97]. Self [AS92, BHNS00, HS91b, HM93, She95, VO93, Elm04]. Self-Adjusting [HM93, She95, Elm04]. Self-alignments [AS92]. Self-Organizing [VO93, HS91b]. Self-Simulation [BHNS00]. semantic [MOPV07]. Semi [BS95b, HLLT06, EF05]. Semi-matchings [HLLT06]. semi-online [EF05]. Semi-Random [BS95b]. Semicomplete [BJY01, Yeo99, BJMT92]. Semidefinite [AEH01, HZ01a, FL06]. semiorders [Ave92]. semirandom [CO07]. Sensitive [BS97b]. separability [CDBPL^{+}09]. separable [HN89]. separation [AF04, ABT04, DK85, NPT92, Sko03, Tre09]. Separator [Coh96, GHT84, Ven87]. Separators [DDSV93, Ric86]. Sequence [Epp90, FFV80, KL00, Dev85, EG81, SFBY04]. Sequences [MS95, Bri87, Etz86, FHS03, HJW05, Hua90, Ral81, Sav89]. sequencing [Vak03]. Sequential [CGPR97, DGM01, PV93, SU84]. serial [LV89]. Series [BT93, CD95, HSH^{+}98, ZSN96, ZSN97, He91a, JLR06, Win86]. Series-Parallel [BT93, HSH^{+}98, ZSN96, ZSN97, JLR06, Win86]. served [PM89]. Server [BR97, BM05, MMS90]. Servers [CL94, CL91]. services [AMRR05, KM03]. session [And04]. session-oriented [And04]. Set [AEST93, BHP01, BFG^{+}96, CH01b, Eng95, IWJ93, JMB96, KVY94, Kri97, LN97, LS97, Sla97, WCP93, Yeo99, AF04, ABI86, CND03, EFF04, Gho91, GHPT03, Gro02, JZC04, Ove81, PR91, Ram88, Sem84, SSW84, YW04]. Set-Set [WCP93]. Sets [Che01b, DG00, Gab98, HS95b, JJ00, KNS01, KP98b, PV93, Pri99, SS00a, BT83, KMNF92, Kla81, Leu84, Lov87, Rob86, Yel92]. setting [FV82]. Shape [CY87]. shaped [TPMRC05]. Shared [AS97, AH90, PRZP09]. Sharing [Att00]. Shellsort [Goo11, PS97, Poo93, Sed86, WS90, Yao80]. Shift [Tho02]. shifted [Sag80]. shifting [Tiw86, BP83]. Shoot [HS95a]. Shooting [AS93, AS96b, GT97, HS95a, CJ92]. Shop [BG99, JP02, QS02]. Shops [DL93a]. Short [BJJ97, Val84, Ric86]. Shortcutting [Tho97]. Shortest [BGHK95, Che96, Coh96, Coh97, CGPR97, Fre95, FMSN00, GG98, GT97, Hal97a, KS97, KS98, MP93, RR96, SS99, Tho00, adHV99, BHS07, Bes03, Dra05, FMO^{+}07, FMSN03, HG88, HW92, Mey03, MS03, Nos85, RV92]. Shortest-Path [Coh97, RR96, adHV99]. Shortest-Paths [Coh96, Mey03]. Shredders [CT99]. shuffle [CSS89]. shuffle-exchange [CSS89]. sieve [Wil91]. sieves [Pri83]. sign [SL06]. sign-reversing [SL06]. similar [Col86]. Simple [AS93, AR00, Ano98e, BS97b, BK93, Boy05, Che96, DC93, FS96, Goo11, GW97, He01, Hsu02, Mut00, NW00, Ron94, Tho96, Yeh01, ADKP89, ABI86, AZ04, GY83, HY88, LS87b, Spr89]. Simpler [sH02a, IK94]. Simplicity [Rag93]. Simplified [SS87]. Simulation [BHNS00, ABNLP92]. simulations [MOPV07]. simultaneous [Vis83]. Single [ABNLP92, ANS97, Hoo96, KS97, KS98, RCS86, SS99, Tho00, Mas06, Mey03, vSL05]. Single-Machine [Hoo96]. Single-Source [KS97, KS98, SS99, Mey03]. Single-Target [ANS97]. Size [Fer98, IM01b, Bru08, DLO03, KK06, Sni86, dBGK^{+}05]. sizes [Gro02]. sketch [CM05]. Skew [dFKKR00]. Sky [LMY98]. Slide [AAG^{+}97]. Sloan [LMY98]. Small [BMR98, CV95, CZ01, DR99, Ind01, KP98b, Xu00, ABK92, BDRG04, GP02, HNZ02, LL88]. Small-Rank [DR99]. Small-Stretch [CZ01]. Smaller [BSV93]. smallest [KL85]. smooth [SV87]. Smoothing [ABE99]. Snakes [ELW97]. Snapshots [AW99]. sociable [PRZP09]. SODA [Cla98]. Solitude [HKAA97]. Solution [Blu00, Gre88, Vil00, Wan97, BGY80, DF89, Len84, Lot92]. solutions [AF04, EO85, Mon80, Sto85]. solvable [BMZ90, CN89, LS88]. solvers [MRFL08]. Solving [BZ88, Bru08, CO07, FBS89, FP09, Fre95, IWJ93, GM08, JN08]. Some [BN02, BN03, Blu00, IZ94, JPS01, JKL00, Shp00, Szp88, Ale85, Bey83, DF89, Sch82, Ven87]. Sort [Che98a, Rüb97, GM84]. sortable [RT84]. sorters [CS92]. Sorting [Goo11, LL99, LMRR94, MS90a, MR92, Raj95, Tho02, Aig82, ABIR89, CS92, Ehr81, Han04, MV91, Sch88a, SV81, War86]. sortings [KV83]. Source [IS96, KS97, KS98, SS99, Tho00, Mey03, Wu02]. Sources [AIMF02, DK95]. Space [AS97, Asp93, CKS^{+}98, DE87, JLR06, MRR01, SZ99, Tho02, CM87, CW02, Han04, MP89, Mel81, PY92, Pea81, vLW81]. Space-Efficient [Asp93, Pea81]. spaces [SU84]. Spanners [KP94]. Spanning [AO00, ADNP99, BJY01, BM97, DN99, EIT^{+}93, Epp94, FKK^{+}97, Fer98, FSO99, HZ01b, JM95, KMN^{+}99, LR98, MUI01, PW98, WCT00, Ben80, CM88, CDN89, EIT^{+}92, Fre81, Gav87, Gué83, HL03, Len87, LLY03, Tiw86, Wu02]. Sparse [ACZ98, BG02, HR99, KP94, NH01, PW97, Shp00, Thu97, HH81]. Sparser [AR93]. Special [Bra07, KLL88, SL06]. specification [GHCMS09]. specifications [Tre09]. Specified [BJJ00, KL00, PtL96, Wan00, MP83]. spectra [BF84]. Speed [IS96, AFN04, AZ04]. speed-up [AFN04]. Speeding [Blu00]. Speeds [CS99]. Speedup [Alt93]. spiral [EC90]. Splay [Geo04]. Splaying [Sub96]. Split [Dah00, MS94a, HHHW86, HHL88, HW84, Hua87, Per84]. splitters [HRJ^{+}89, MW92]. Splitting [BJJ00, FM99, Ron94, Wan05]. Spreading [AH98, Cab07]. Squares [BPK00]. SSSP [ABT04]. Stabbing [GIK02].

Stability [GH88]. Stable [Knu96, Irv85, Ron90, SS87, Tan91, ZC09, IM02]. stack [RT84]. stack-sortable [RT84]. Stacker [COKN06]. Stacks [Bur81, FG94]. Star [BHNS00]. starting [NH82]. Static [DLSS95, Hen97, BS80]. Static-to-dynamic [BS80]. statistical [HS91b, PCBC^{+}09]. Steiner [BDGW98, BR94, CCyC^{+}99, FR94, GKR00, GP87, KR95b, KPRT97, PS00, Win86]. step [HM85]. step-function [HM85]. stepping [MS03]. Steps [BN02, BN03]. Stochastic [DR08, KRS98, HK86]. stone [vO09]. stopping [PPR04]. storage [BC82, Kno82]. Store [dlTK97, Bur81]. storing [Col86]. straightline [BL92]. Strategies [LN97, MOPV07, ABNLP90, HMI09, LB92, WG08]. Stream [CJM02, BNL03, CS06, CM05]. Stretch [CZ01, Cow01, EGP03]. strike [Dol82]. Strikes [GK99]. String [AR00, BCT98, Fer97, GZ94, ALP04, Gol82, Hei05, LV89, ML84, MNU05, SZ05a]. Strings [BJJ97, Col94, Wal98, Shi81, Ukk85]. Strong [BP04, BJY01, Sor88b, ZC09]. Structural [BT95b, JN08]. Structure [Fre97, LS91, SH02b, Tsa03]. Structured [BC99, KP98a, vO09]. Structures [BGH99, CST96, EGMS94, FJMO95, GG97, HM93, Rou01, dlTK97, DM02, FK03, FFV80, FJ83, LS85, Ove88]. studies [WG08]. Study [GT01, JN08, SWMV03]. Sub [Thu97]. Sub-linear [Thu97]. Subadditive [Joh80]. Subdigraph [BJY01]. Subdivisions [BJM94, DVS95, GT97]. Subedge [Che96]. Subgraph [AITT00, ADNP99, BS97a, Fer98, MS94b, Chu87]. Subgraphs [CFFK98, DN99, KR95a, BLW87, GS85]. Subgroups [KLM99, Kan85]. Sublinear [GPV93]. Sublinear-Time [GPV93]. Submodular [Gab95]. Subquadratic [WMM95, Lar87]. Subsequence [BG02, Wil88]. Subset [BJJ00, Epp97, Pri99]. Subsets [EH97, HN96, Sem84]. Substring [LZU01]. Subtree [ST99]. succinct [ABNLP90, Can81]. sufficient [Tan91]. Suffix [MRR01, Sad03]. suite [VGC^{+}08]. Sum [BNK98, BNHK^{+}00]. summary [CM05]. Sums [Epp97, TL96, Zan91]. Sundar [Ano02i]. superconcentrators [AGM87]. Superlinear [Kla92]. Superstrings [BJJ97, CGPR97]. surely [COKN06]. Surface [Pel96]. Survey [Gor98, LMY98]. Swaps [AAL^{+}00]. switchbox [KM86]. switches [KR06, Zhu04]. Sylow [Kan85, KT88, Kan90, KLM99]. symbolic [Wol87]. Symbols [Shp00]. symmetric [SL84]. synergy [HS91b]. syntenic [LN02]. synteny [LN02]. Synthesis [BKSD97]. system [HS84, Tol82, AR05]. Systems [Att00, Liu02, Vil00, BGY80, CCMS02, Fei81, Fei83a, Gro02, JZ90, LS83, Len84, LP81a, Lot92].

table [KdlT89, Szy85]. tableau [Sag80]. tableaux [SL06]. tables [ABNLP90]. tailored [HS91a]. Take [HS95a]. Target [ANS97, HRZ03]. Tasks [AKP^{+}97, CL99, HLW94, SH98, BMZ90, NH82, PCBC^{+}09, SS87]. tautologies [Bed07]. TCP [FLNU03]. Technique [FKK^{+}97, KVY94, FL06, Wac89]. Techniques [EKK^{+}03, HRG00, MRFL08]. temporal [Tre09]. Temporary [AKP^{+}97]. Tensor [Hås90]. Test [Hsu02, VGC^{+}08]. Testing [AS03, BGLR99, FM05, LP81a, Yel92]. Tests [Bod93]. tetrahedralizations [YW04]. Text [AKL^{+}00, Fer97, FG99, DLO03, Sad03]. th [AS96a, FJ83]. Their [HM02b, Sie01, AS92, GS85]. Theorem [BBDL01, Ver94, Ans85, GHT84, KT88, LLL86, Nos85]. theorems [Ven87]. theoretical [BW85]. Theory [KL94, CHT02, Lag80, Sel80]. threaded [Bri86]. Three [BHP01, BK93, BT99, CL94, DVS95, FG94, Hoo96, IM01a, DF02, LC92]. Three-Colored [BK93]. Three-Dimensional [DVS95, IM01a]. Threshold [DB97]. Thresholds [AKKL02]. throughput [GHP05]. throughput-competitive [GHP05]. Ties [IM02]. Tight [AR05, BS97a, Sla97, WS90]. Tighter [Wan97, Sch88a]. tile [Bru08]. tileset [Bru08]. Tiling [BDMR01, SS00b, BDM03]. Time [AS97, Asp93, BK00, BDP97, BKN00, Bod93, BK93, CM96, CL95, DF94, DL93a, DL93b, Fle99, GPV93, He01, HW00a, KS98, KL97, Kri96, LR98, MS98, Moh93, MW94, NI99, NW00, Pis99, Rüb97, SS99, SU97, Sub99, Tho02, Wal98, Wei97, Wes00, WCT00, Yeo99, BKL^{+}04, BYE81, BYES05, BE05, BLW87, Ber05, Boy05, Cha03, CL02, CN89, Chu87, CS06, CO07, DF02, Dia80, DGM^{+}06, DF89, FM02, Han04, JP02, Kan85, KT88, Kan90, KT80, Kim05, LLZ05, Mak02, Mas06, Mel81, Pit87, RBLL89, Sko03, TAB86, TSB05a, Wan05, Wu02, Yel92, vSL05]. Time- [Asp93]. time-decaying [CS06]. Time-Randomness [Kri96]. Time-Space [AS97]. Time-Work [SS99]. time/space [Mel81]. times [Gus84, HM89, Mas06, NH82, Vak03]. Toeplitz [BGY80]. tolerance [NT90]. Tolerant [KKP01, BYK88, GMM03]. Tools [HK02]. Topological [Fu97, KV83]. Topology [KPP98, CR06, HW92]. Tori [MS02b]. Toroidal [HW00a]. Total [FWW93, Kim05, SSZ86, vSL05]. Totally [BFS97, AF84, AK92]. Tournament [Hav00, Sor88a]. Tournaments [BRS97, BEMS95, HR83]. Towers [Er85, Min89]. Track [FKN99]. tractable [EFF04]. Trade [Kri96, Sni86]. Trade-offs [Kri96, Sni86]. Tradeoff [AS97, Mel81]. Tradeoffs [SS99]. Trails [HSH^{+}98]. trains [ABIR89]. transaction [LP81a]. transactions [Wol86]. transfers [GHPT03]. transform [Dia80, GM84]. transformation [BS80]. Transformations [FH98, LZ99, dLFM07]. transforms [BS91a]. Transitive [GKR^{+}91, BHS07]. translates [MS90c]. translation [BFM86]. Translator [BR97]. transmissions [IR85]. Transportation [ABK^{+}95]. transposed [KKB88]. Transposition [MNU05]. transpositions [RP90]. transversals [Wah04]. Traveler [AT95, HL92]. Traveling [FJMO95, BYES05, HRZ03]. travelling [PV84]. Traversal [MW94, Bri87]. Traversing [Bur96, Ram95b]. Tree [AO00, Alt93, BR94, BBDS00, BGHK95, BT95a, Che98b, Che01a, DFKP04, EGMS94, Epp94, FBS97, FS94, FG93, FR94, GKR00, GT01, GN95, GBH97, HM96, yHHsH^{+}00, KLW96, KC98, Lag96, LP98, PSY93, PtL96, PS00, PW98, Rou01, SU97, WG97, Wan00, Wan94, War94, ZSW94, Zha98, ADKP89, ALS91, BKL^{+}04, BP83, CFR03, CDHX03, FG03, GP87, Gué83, HL03, HY88, LM90, MT91, MP83, MS80, PR85, RS86, Sko03, TPMRC05, Tiw86, WPYK06]. tree-decomposable [ALS91]. tree-decompositions [MT91]. tree-layouts [Sko03]. Tree-to-Tree [Che01a, FG03]. Tree-Width [FBS97, FS94, Lag96, Wan94, CFR03, RS86]. Trees [AT00, And99, Ano98f, BCD99, BM97, BS01, CF96, DGK01, FKK^{+}97, Fre97, FSO99, FMSN00, GI99, Gup01, HK02, HM02a, JM95, KLST01, KR95b, KL00, KMN^{+}99, LVV96, LZ99, LOÖ01, LR98, LVR93, MUI01, MRR01, NRT02, PW97, She95, Tho97, WCT00, ZiNN96, vB00, Alt88, AKKL02, BK06, BFCP^{+}05, Ben80, Bod90, Bri86, CM88, Chu87, COKN06, CDN89, DLZZ03, ES81, EK04b, FHS03, FBS89, GS85, Gav87, Geo04, GLW83, HHHW86, HHL88, HW84, Hua87, HHL^{+}02, Kli80, Kno82, Lar87, LLY03, Luc87, MP89, Mai88, Per84, PM85, RP90, SZ90, SY82, TPMRC05, TSB05b, VW82, Vel86, Wil81, Wu02]. Treewidth [BGHK95, BK96, BE97, HM02b, KK95]. triangle [BZ88, Pla84, ZLL09]. triangles [KL85, MW92, OAMB86]. Triangulating [BK93, Goo89]. Triangulation [LK98, AAP86, BBH^{+}06, PH87]. Triangulations [GT97, LK98, BDRG04]. Triconnected [sH00]. Trie [dlTK97]. tries [KS02, Szp88]. Triple [KKM99]. Triple-Free [KKM99]. Triply [BMR98]. Triply-Logarithmic [BMR98]. True [BN02, BN03]. TSP [Blä04, DM03, KLS89, dBGK^{+}05]. turn [ABIR89]. Tutte [Ans85]. Twist [LZ99]. Twist-Rotation [LZ99]. Two [AS99, ABF97, AE97, BFG^{+}96, CMN96, DVS95, DL93a, EGMS94, FS96, Hoo96, IM01b, KR95a, PB99, Sor94, Tam85, Wil91, Zha98, BBK81, BC82, BT83, BZ88, Bur81, CH03, CL91, CDBPL^{+}09, Ede85, EF05, FP89, GT88, Meg85, PPR04, PV84, Vel86, Wu02]. Two-Connected [KR95a, PB99]. Two-Dimensional [ABF97, DVS95, IM01b, BBK81, BC82, BZ88, Vel86]. two-level [Bur81]. Two-Machine [DL93a]. two-point [FP89]. two-source [Wu02]. Two-Structures [EGMS94].

Ultrafast [MS98]. unanimity [LLZ05]. unavoidability [Hei05]. Unbalanced [KLST01]. Unbounded [Bre95]. uncertainty [FMO^{+}07]. Undirected [AIMF02, BK00, CZ95, MS94a, PP99, Wei97, CPR03]. Unequal [BGLR02]. Unfolding [DF94]. Unidimensional [DF94]. unidirectional [DKR82]. unified [Dra05, Pan81b]. Uniform [Bre95, CBS04, KR96, Kri97, KNS01, MW90, dlTK97, Dev85, KK06, KS03, KLL88, Wil81]. Uniformly [CB01]. Unifying [KLST01]. Union [Ano81c, LN97, CJ91, LP80]. Unions [JJ00]. Unique [GKT01, KOPS00]. unit [DGM^{+}06, SZ90, VGC^{+}08]. Universal [Bri87]. universality [HT90]. Unknown [PP99, CR06]. Unlabeled [CRS^{+}00, DW83, Gol92, SR03]. unlocking [Wol86]. Unprovability [LN97]. Unranking [CDN89, CMN96, GLW83]. unrelated [Sit05]. unsplittable [KS06]. Unstructured [MTT99]. Unsuccessful [HM93]. unweighted [CFR03, SWMV03]. Update [CN00]. Updates [Fer97]. updating [RV92]. Upgrading [KMN^{+}99]. Upper [BMR98, DB97, Gal00, NR00, Mey03, Sch88a, Sed86]. Use [ES95, GM84]. uses [DM85]. Using [ABK^{+}95, AEH01, BY01, Bur96, CLRW97, Coh97, Fre95, Hal97a, KLW96, NRP96, NPT92, Tho02, AH90, Aur88, BBBD89, BZ88, HNZ02, JZ90, LKPC05, Ric86, vSL05].

validator [dLFM07]. Valued [KS98]. Values [BSV93, MR99]. Variable [SIN92, ZSW94, MS90b, WG08]. Variable-priority [SIN92]. Variance [Eng95, Zha98]. variate [FS86]. varying [dBGK^{+}05]. Vector [ET03]. vectorial [Oli81]. Vectors [Dor98, DG01, Pan80]. Vegas [KOPS00]. vehicle [AHL06]. Venetian [CIM^{+}02]. verifiability [HT90]. version [AJKL84]. versions [KT88]. Vertex [ADNP99, BYR02, CKJ01, DN99, FN01, Hen97, HRG00, KVY94, PSK97, BYE81, GS85, GHKO03, Nag05, NR03, Sko03]. Vertices [Yeo99]. Very [BPK00, Czu96]. via [BT95b, Epp97, GKR^{+}91, GZ94, GT97, KLST01, LLY03, PS95]. view [HHL^{+}02]. Virtual [Sch05a]. virtues [Wol87]. Virus [GGPS98]. Vishwanathan [Ano02i]. Visibility [DC93, Gho96, EA81, EC90, Gho91, OW88]. Visible [Che96]. Vizing [GP00]. Volume [Ano93a, Ano93b, Ano94a, Ano94b, Ano95a, Ano95b, Ano96a, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02d, BHP01, EIT^{+}93, Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano02c, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano81c, Ano02i]. Volumes [Ano96b]. Voronoi [CKS^{+}98, MN98]. voter [JMP09]. voter-controlled [JMP09]. vs [Geo04].

W [Ano81c]. Wachs [Kin88]. Wait [MT97]. Wait-Free [MT97]. Walk [HS95a]. Walks [KRS98]. Wavelength [CIM^{+}02, KK01]. Way [AEH01, ABT04]. Weakly [Che96]. Weight [BM97, CV95, FKK^{+}97, FSO99, HT97, LK98, Sch98, Wil88]. Weighted [BS85a, BYR02, CH01b, CF96, EH97, HLW94, KP93, KVY94, KR95b, PSK97, Wan00, BYE81, CH88, Fra81, GVY04, Gav87, MZ86, NR03, SRB06, SR06]. weighting [BP83]. Weights [AK97, BY01, Pis99, FMSN03, WG08]. Weil [BMX06]. Weil/Tate [BMX06]. Well [Wal98]. Well-Formed [Wal98]. Where [BN02, BN03]. which [Bed07, COKN06, LS88, Sko03]. whose [Fei81, Fei83a]. wide [BBH^{+}06]. wide-range [BBH^{+}06]. Width [Epp00, FBS97, FS94, Lag96, Wan94, CFR03, RS86]. Wildcards [GG97]. window [Blu87]. windows [BYES05]. Wire [FWW93]. wireless [BNM04]. wiring [RCS86]. Wise [Ind01, Tho02]. within [BJJ00, FR94]. Without [GLD94, Blu00, CDHX03]. word [Blu87, DA85]. Words [FH98, AS92, Duv83]. Work [Bur96, CLRW97, SS99]. Worst [FM02, GG98, Lag80, Poo93, SA00, Wal98, Xu00, vV96, Sch88a]. Worst-Case [Wal98, Xu00, FM02, Lag80]. Write [MS94c]. Write-All [MS94c].

XML [dLFM07].

Yao [Wac89]. young [Sag80].

Zero [Rón88b]. Zeroes [KL93]. Zone [NBHB00].

References

Afek:1997:SKP

[AAG^{+}97] Yehuda Afek, Baruch Awerbuch, Eli Gafni, Yishay Mansour, Adi Rosén, and Nir Shavit. Slide — the key to polynomial end-to-end communication. Journal of Algorithms, 22(1):158–186, January 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908196.

Amir:1997:IPM

[AAL97] Amihood Amir, Alberto Apostolico, and Moshe Lewenstein. Inverse pattern matching. Journal of Algorithms, 24(2):325–339, August 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908500.

Amir:2000:PMS

[AAL^{+}00] Amihood Amir, Yonatan Aumann, Gad M. Landau, Moshe Lewenstein, and Noa Lewenstein. Pattern matching with swaps. Journal of Algorithms, 37(2):247–266, November 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911209.

Ajtai:1998:FS

[AAN^{+}98] Miklos Ajtai, James Aspnes, Moni Naor, Yuval Rabani, Leonard J. Schulman, and Orli Waarts. Fairness in scheduling. Journal of Algorithms, 29(2):306–357, November 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909531.

Asano:1986:PTE

[AAP86] Takao Asano, Tetsuo Asano, and Ron Y. Pinter. Polygon triangulation: Efficiency and minimality. Journal of Algorithms, 7(2): 221–231, June 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900052.

Amenta:1999:OPP

[ABE99] Nina Amenta, Marshall Bern, and David Eppstein. Optimal point placement for mesh smoothing. Journal of Algorithms, 30(2):302– 322, February 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909841.

Amir:1997:OTD

[ABF97] Amihood Amir, Gary Benson, and Martin Farach. Optimal two-dimensional compressed matching. Journal of Algorithms, 24(2):354–379, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908609.

Awerbuch:1998:DPG

[ABF98] Baruch Awerbuch, Yair Bartal, and Amos Fiat. Distributed paging for general networks. Journal of Algorithms, 28(1):67–104, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909245.

Alon:1986:FSR

[ABI86] Noga Alon, László Babai, and Alon Itai. A fast and simple randomized parallel algorithm for the maximal independent set problem. Journal of Algorithms, 7(4):567–583, December 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900192.

Amato:1989:RTT

[ABIR89] Nancy Amato, Manuel Blum, Sandra Irani, and Ronitt Rubinfeld. Reversing trains: a turn of the century sorting problem. Journal of Algorithms, 10(3):413–428, September 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900370.

Averbuch:1992:CAM

[ABK92] Amir Averbuch, Nader H. Bshouty, and Michael Kaminski. A classification of algorithms for multiplying polynomials of small degree over finite fields. Journal of Algorithms, 13(4):577–588, December 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290057J.

Aggarwal:1995:EMC

[ABK^{+}95] A. Aggarwal, A. Barnoy, S. Khuller, D. Kravets, and B. Schieber. Efficient minimum cost matching and transportation using the quadrangle inequality. Journal of Algorithms, 19(1):116–143, July 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710309.

Awerbuch:1990:IRS

[ABNLP90] Baruch Awerbuch, Amotz Bar-Noy, Nathan Linial, and David Peleg. Improved routing strategies with succinct tables. Journal of Algorithms, 11(3):307–341, September 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900179.

Alon:1992:SRS

[ABNLP92] Noga Alon, Amotz Bar-Noy, Nathan Linial, and David Peleg. Single round simulation on radio networks. Journal of Algorithms, 13(2):188–210, June 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900155.

Arge:2004:EMM

[ABT04] Lars Arge, Gerth Stølting Brodal, and Laura Toma. On externalmemory MST, SSSP and multi-way planar graph separation. Journal of Algorithms, 53(2):186–206, November 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000641.

Alexander:2003:ESD

[ABW03] D. L. J. Alexander, D. W. Bulger, and G. R. Wood. Expected search duration for finite backtracking adaptive search.

Journal of Algorithms, 47(2):78–86, July 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740300018X.

Agarwal:1998:AAH

[AC98] Sameet Agarwal and Anne Condon. On approximation algorithms for hierarchical MAX-SAT. Journal of Algorithms, 26(1):141– 165, January 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497909020.

Amir:2000:AIS

[AC00] Amihood Amir and Gruia Calinescu. Alphabet-independent and scaled dictionary matching. Journal of Algorithms, 36(1):34–62, July 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400910812.

Arikati:1998:APM

[ACZ98] Srinivasa R. Arikati, Shiva Chaudhuri, and Christos D. Zaroliagis. All-pairs min-cut in sparse networks. Journal of Algorithms, 29(1):82–110, October 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909610.

aufderHeide:1999:SPR

[adHV99] Friedhelm Meyer auf der Heide and Berthold Vöcking. Shortestpath routing in arbitrary networks. Journal of Algorithms, 31(1): 105–131, April 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909804.

Althaus:2003:EGA

[ADK^{+}03] Ernst Althaus, Denys Duchier, Alexander Koller, Kurt Mehlhorn, Joachim Niehren, and Sven Thiel. An efficient graph algorithm for dominance constraints. Journal of Algorithms, 48(1):194– 219, August 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000506.

Abrahamson:1989:SPT

[ADKP89] K. Abrahamson, N. Dadoun, D. G. Kirkpatrick, and T. Przytycka. A simple parallel tree contraction algorithm. Journal of Algorithms, 10(2):287–302, June 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900175.

Alon:1994:AAR

[ADL^{+}94] N. Alon, R. A. Duke, H. Lefmann, V. Rodl, and R. Yuster. The algorithmic aspects of the regularity lemma. Journal of Algorithms, 16(1):80–109, January 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710054.

Auletta:1999:AAF

[ADNP99] Vincenzo Auletta, Yefim Dinitz, Zeev Nutov, and Domenico Parente. A 2-approximation algorithm for finding an optimum 3vertex-connected spanning subgraph. Journal of Algorithms, 32 (1):21–30, July 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910064.

Azar:1997:TDP

[AE97] Yossi Azar and Leah Epstein. On two dimensional packing. Journal of Algorithms, 25(2):290–310, November 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908762.

Andersson:2001:NWU

[AEH01] Gunnar Andersson, Lars Engebretsen, and Johan Håstad. A new way of using semidefinite programming with applications to linear equations mod p. Journal of Algorithms, 39(2):162–204, May 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911544.

Alt:2003:MPM

[AERW03] Helmut Alt, Alon Efrat, Günter Rote, and Carola Wenk. Matching planar maps. Journal of Algorithms, 49(2):262–283, November 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000853.

Azar:2004:ANA

[AERW04] Yossi Azar, Leah Epstein, Yossi Richter, and Gerhard J. Woeginger. All-norm approximation algorithms. Journal of Algorithms, 52(2):120–133, August 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000392.

Agarwal:1993:CSC

[AEST93] P. K. Agarwal, A. Efrat, M. Sharir, and S. Toledo. Computing a segment center for a planar point set. Journal of Algorithms, 15(2):314–323, September 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710436.

Anstee:1984:CTB

[AF84] R. P. Anstee and Martin Farber. Characterizations of totally balanced matrices. Journal of Algorithms, 5(2):215–230, June 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677484900282.

Alber:2004:GSE

[AF04] Jochen Alber and Jirí Fiala. Geometric separation and exact solutions for the parameterized independent set problem on disk graphs. Journal of Algorithms, 52(2):134–151, August 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001639.

Alber:2004:PCE

[AFN04] Jochen Alber, Henning Fernau, and Rolf Niedermeier. Parameterized complexity: exponential speed-up for planar graph problems. Journal of Algorithms, 52(1):26–56, July 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000471.

Alon:2004:ALD

[AGK04] Noga Alon, Gregory Gutin, and Michael Krivelevich. Algorithms with large domination ratio. Journal of Algorithms, 50(1):118– 131, January 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001603.

Alon:1987:BES

[AGM87] N. Alon, Z. Galil, and V. D. Milman. Better expanders and superconcentrators. Journal of Algorithms, 8(3):337–347, September 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900149.

Atallah:1987:BMM

[AH87] Mikhail J. Atallah and Susanne E. Hambrusch. On bipartite matchings of minimum density. Journal of Algorithms, 8(4):480– 502, December 1987. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677487900472.

Aspnes:1990:FRC

[AH90] James Aspnes and Maurice Herlihy. Fast randomized consensus using shared memory. Journal of Algorithms, 11(3):441–461, September 1990. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677490900216.

Aspnes:1998:SRR

[AH98] James Aspnes and William Hurwood. Spreading rumors rapidly despite an adversary. Journal of Algorithms, 26(2):386–411, February 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667749790907X.

Arkin:2006:AMM

[AHL06] Esther M. Arkin, Refael Hassin, and Asaf Levin. Approximations for minimum and min-max vehicle routing problems. Journal of Algorithms, 59(1):1–18, April 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000258.

Aigner:1982:PCS

[Aig82] Martin Aigner. Parallel complexity of sorting problems. Journal of Algorithms, 3(1):79–88, March 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900104.

Aggarwal:1991:FPM

[AIKS91] Alok Aggarwal, Hiroshi Imai, Naoki Katoh, and Subhash Suri. Finding k points with minimum diameter and related problems. Journal of Algorithms, 12(1):38–56, March 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190022Q.

Arata:2002:LSM

[AIMF02] Kouji Arata, Satoru Iwata, Kazuhisa Makino, and Satoru Fujishige. Locating sources to meet flow demands in undirected networks. Journal of Algorithms, 42(1):54–68, January 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401912039.

Ausiello:1991:IAM

[AISN91] Giorgio Ausiello, Giuseppe F. Italiano, Alberto Marchetti Spaccamela, and Umberto Nanni. Incremental algorithms for minimal length paths. Journal of Algorithms, 12(4):615–638, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190036X.

Asahiro:2000:GFD

[AITT00] Yuichi Asahiro, Kazuo Iwama, Hisao Tamaki, and Takeshi Tokuyama. Greedily finding a dense subgraph. Journal of Algorithms, 34(2):203–221, February 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910623.

Assmann:1984:DVO

[AJKL84] S. F. Assmann, D. S. Johnson, D. J. Kleitman, and J. Y.-T Leung. On a dual version of the one-dimensional bin packing problem. Journal of Algorithms, 5(4):502–525, December 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748490004X.

Audemard:2008:SGB

[AJS08] Gilles Audemard, Said Jabbour, and Lakhdar Sais. SAT graph-based representation: a new perspective. Journal of Algorithms, 63(1–3):17–33, January/July 2008. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677408000047.

Atallah:1992:EPA

[AK92] Mikhail J. Atallah and S. Rao Kosaraju. An efficient parallel algorithm for the row minima of a totally monotone matrix. Journal of Algorithms, 13(3):394–413, September 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290046F.

Alon:1997:CAW

[AK97] Noga Alon and Dmitry N. Kozlov. Coins with arbitrary weights. Journal of Algorithms, 25(1):162–176, October 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908713.

Albers:1998:NLA

[AK98] Susanne Albers and Hisashi Koga. New on-line algorithms for the page replication problem. Journal of Algorithms, 27(1):75– 96, April 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497909068.

Anderson:2002:TOB

[AKKL02] Richard Anderson, Sampath Kannan, Howard Karloff, and Richard E. Ladner. Thresholds and optimal binary comparison search trees. Journal of Algorithms, 44(2):338–358, August 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402002031.

Amir:2000:TID

[AKL^{+}00] Amihood Amir, Dmitry Keselman, Gad M. Landau, Moshe Lewenstein, Noa Lewenstein, and Michael Rodeh. Text indexing and dictionary matching with one error. Journal of Algorithms, 37(2):309–325, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911040.

Abiteboul:1996:MPC

[AKM^{+}96] Serge Abiteboul, Gabriel M. Kuper, Harry G. Mairson, Alex A. Shvartsman, and Moshe Y. Vardi. In memoriam: Paris C. Kanellakis (1953–1995). Journal of Algorithms, 20(2):203–204, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900103.

Azar:1997:LLB

[AKP^{+}97] Yossi Azar, Bala Kalyanasundaram, Serge Plotkin, Kirk R. Pruhs, and Orli Waarts. On-line load balancing of temporary tasks. Journal of Algorithms, 22(1):93–110, January 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485707997.

Alekseyev:1985:CSA

[Ale85] Valery B. Alekseyev. On the complexity of some algorithms of matrix multiplication. Journal of Algorithms, 6(1):71–85, March 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677485900197.

Allison:1981:GCR

[All81] Lloyd Allison. Generating coset representatives for permutation groups. Journal of Algorithms, 2(3):227–244, September 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677481900249.

Amir:2000:PMH

[ALL00] Amihood Amir, Moshe Lewenstein, and Noa Lewenstein. Pattern matching in hypertext. Journal of Algorithms, 35(1):82–99, April 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910635.

Amir:2004:FAS

[ALP04] Amihood Amir, Moshe Lewenstein, and Ely Porat. Faster algorithms for string matching with k mismatches. Journal of Algorithms, 50(2):257–275, February 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740300097X.

Arnborg:1991:EPT

[ALS91] Stefan Arnborg, Jens Lagergren, and Detlef Seese. Easy problems for tree-decomposable graphs. Journal of Algorithms, 12(2):308– 340, June 1991. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749190006K.

Amir:2003:IMC

[ALS03] Amihood Amir, Gad M. Landau, and Dina Sokol. Inplace 2D matching in compressed images. Journal of Algorithms, 49(2):240–261, November 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000889.

Althofer:1988:CSG

[Alt88] Ingo Althöfer. On the complexity of searching game trees and other recursion trees. Journal of Algorithms, 9(4):538–567, December 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900168.

Althofer:1993:PGT

[Alt93] I. Althofer. A parallel game tree search algorithm with a linear speedup. Journal of Algorithms, 15(2):175–198, September 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710370.

Amir:1992:EPM

[ALV92] Amihood Amir, Gad M. Landau, and Uzi Vishkin. Efficient pattern matching with scaling. Journal of Algorithms, 13(1):2– 32, March 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290003U.

Akl:1982:ACC

[AM82] Selim G. Akl and Henk Meijer. On the average-case complexity of “bucketing” algorithms. Journal of Algorithms, 3(1):9– 13, March 1982. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677482900037.

Aggarwal:1986:FCM

[AM86] Alok Aggarwal and Robert C. Melville. Fast computation of the modality of polygons. Journal of Algorithms, 7(3):369–381, September 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900283.

Afek:1999:CCO

[AMO99] Yehuda Afek, Yishay Mansour, and Zvi Ostfeld. Convergence complexity of optimistic rate-based flow-control algorithms. Journal of Algorithms, 30(1):106–143, January 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909701.

Aiello:2005:CQP

[AMRR05] William A. Aiello, Yishay Mansour, S. Rajagopolan, and Adi Rosén. Competitive queue policies for differentiated services. Journal of Algorithms, 55(2):113–141, May 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000665.

Aggarwal:2006:LRP

[AMZ06] Gagan Aggarwal, Rajeev Motwani, and An Zhu. The load rebalancing problem. Journal of Algorithms, 60(1):42–59, July 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404001518.

Andersson:1999:GBT

[And99] Arne Andersson. General balanced trees. Journal of Algorithms, 30(1):1–18, January 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909671.

Andrews:2004:IFS

[And04] Matthew Andrews. Instability of FIFO in session-oriented networks. Journal of Algorithms, 50(2):232–245, February 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000968.

Anonymous:1980:AIV

[Ano80a] Anonymous. Author index for volume 1. Journal of Algorithms, 1(4):396, December 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667748090019X.

Anonymous:1980:EB

[Ano80b] Anonymous. Editorial Board. Journal of Algorithms, 1(1): ??, March 1980. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677480900012.

Anonymous:1981:AIV

[Ano81a] Anonymous. Author index for volume 2. Journal of Algorithms, 2(4):406, December 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900389.

Anonymous:1981:EB

[Ano81b] Anonymous. Editorial Board. Journal of Algorithms, 2(1): ??, March 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900018.

Anonymous:1981:EVN

[Ano81c] Anonymous. Erratum: Volume 1, Number 3 (1980), “Finding the Contour of a Union of Iso-Oriented Rectangles,” by W. Lipski, Jr., and France P. Preparata, pp. 235–246. Journal of Algorithms, 2 (1):105, March 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900122. See [LP80]; the er-

ratum just supplies three additional references that were inadvertently omitted.

Anonymous:1982:AIV

[Ano82a] Anonymous. Author index for volume 3. Journal of Algorithms, 3(4):396, December 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677482900335.

Anonymous:1982:C

[Ano82b] Anonymous. Corrigendum. Journal of Algorithms, 3(3):301–302, September 1982. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748290027X.

Anonymous:1982:EB

[Ano82c] Anonymous. Editorial Board. Journal of Algorithms, 3(1): ??, March 1982. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677482900013.

Anonymous:1983:AIV

[Ano83a] Anonymous. Author index for volume 4. Journal of Algorithms, 4(4):412, December 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677483900202.

Anonymous:1983:C

[Ano83b] Anonymous. Corrigendum. Journal of Algorithms, 4(3):301, September 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900299.

Anonymous:1983:EB

[Ano83c] Anonymous. Editorial Board. Journal of Algorithms, 4(1): ??, March 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900305.

Anonymous:1984:AIV

[Ano84a] Anonymous. Author index for volume 5. Journal of Algorithms, 5(4):612–613, December 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677484900129.

Anonymous:1984:EB

[Ano84b] Anonymous. Editorial Board. Journal of Algorithms, 5(1): ??, March 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900348.

Anonymous:1984:PAFa

[Ano84c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 5(1):161–162, March 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900464.

Anonymous:1984:PAFb

[Ano84d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 5(2):300–301, June 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677484900336.

Anonymous:1984:PAFc

[Ano84e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 5(3):448–449, September 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900233.

Anonymous:1984:PAFd

[Ano84f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 5(4):610–611, December 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900117.

Anonymous:1985:AIV

[Ano85a] Anonymous. Author index for volume 6. Journal of Algorithms, 6(4):597–598, December 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900355.

Anonymous:1985:EB

[Ano85b] Anonymous. Editorial Board. Journal of Algorithms, 6(1): ??, March 1985. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677485900148.

Anonymous:1985:PAFa

[Ano85c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 6(1):160–161, March 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900264.

Anonymous:1985:PAFb

[Ano85d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 6(2):306–307, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900471.

Anonymous:1985:PAFc

[Ano85e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 6(3):452–453, September 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900136.

Anonymous:1985:PAFd

[Ano85f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 6(4):596, December 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900343.

Anonymous:1986:AIV

[Ano86a] Anonymous. Author index for volume 7. Journal of Algorithms, 7(4):603–604, December 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900222.

Anonymous:1986:EB

[Ano86b] Anonymous. Editorial Board. Journal of Algorithms, 7(1): ??, March 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900350.

Anonymous:1986:PAFa

[Ano86c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 7(1):157, March 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900441.

Anonymous:1986:PAFb

[Ano86d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 7(2):306–307, June 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667748690012X.

Anonymous:1986:PAFc

[Ano86e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 7(3):448, September 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900349.

Anonymous:1986:PAFd

[Ano86f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 7(4):602, December 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900210.

Anonymous:1987:AIV

[Ano87a] Anonymous. Author index for volume 8. Journal of Algorithms, 8(4):600–601, December 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677487900551.

Anonymous:1987:EB

[Ano87b] Anonymous. Editorial Board. Journal of Algorithms, 8(1): ??, March 1987. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748790023X.

Anonymous:1987:PAFa

[Ano87c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 8(1):157, March 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677487900356.

Anonymous:1987:PAFb

[Ano87d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 8(2):304, June 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677487900447.

Anonymous:1987:PAFc

[Ano87e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 8(3):449, September 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900228.

Anonymous:1987:PAFd

[Ano87f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 8(4):598–599, December 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748790054X.

Anonymous:1988:AIV

[Ano88a] Anonymous. Author index for volume 9. Journal of Algorithms, 9(4):601–602, December 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677488900211.

Anonymous:1988:EB

[Ano88b] Anonymous. Editorial Board. Journal of Algorithms, 9(1): ??, March 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900016.

Anonymous:1988:PAFa

[Ano88c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 9(1):148–149, March 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900119.

Anonymous:1988:PAFb

[Ano88d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 9(2):297–298, June 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677488900442.

Anonymous:1988:PAFc

[Ano88e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 9(3):445–447, September 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748890034X.

Anonymous:1988:PAFd

[Ano88f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 9(4):599–600, December 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748890020X.

Anonymous:1989:AIV

[Ano89a] Anonymous. Author index for volume 10. Journal of Algorithms, 10(4):598–599, December 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677489900096.

Anonymous:1989:EB

[Ano89b] Anonymous. Editorial Board. Journal of Algorithms, 10(1): ??, March 1989. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677489900199.

Anonymous:1989:PAFa

[Ano89c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 10(1):155–156, March 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900308.

Anonymous:1989:PAFb

[Ano89d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 10(2):303–304, June 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900187.

Anonymous:1989:PAFc

[Ano89e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 10(3):449–450, September 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900394.

Anonymous:1989:PAFd

[Ano89f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 10(4):596–597, December 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900084.

Anonymous:1990:AIV

[Ano90a] Anonymous. Author index for volume 11. Journal of Algorithms, 11(4):657–658, December 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677490900168.

Anonymous:1990:EB

[Ano90b] Anonymous. Editorial Board. Journal of Algorithms, 11(1): ??, March 1990. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749090025A.

Anonymous:1990:PAFa

[Ano90c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 11(1):152, March 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749090036E.

Anonymous:1990:PAFb

[Ano90d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 11(2):305, June 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677490900083.

Anonymous:1990:PAFc

[Ano90e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 11(3):522, September 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900249.

Anonymous:1990:PAFd

[Ano90f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 11(4):655–656, December 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900157.

Anonymous:1991:AIV

[Ano91a] Anonymous. Author index for volume 12. Journal of Algorithms, 12(4):706–707, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190044Y.

Anonymous:1991:EB

[Ano91b] Anonymous. Editorial Board. Journal of Algorithms, 12(1): ??, March 1991. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749190019U.

Anonymous:1991:PAFa

[Ano91c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 12(1):186–187, March 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190031S.

Anonymous:1991:PAFb

[Ano91d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 12(2):357–358, June 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190008M.

Anonymous:1991:PAFc

[Ano91e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 12(3):530–531, September 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190018T.

Anonymous:1991:PAFd

[Ano91f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 12(4):704–705, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190043X.

Anonymous:1992:AIV

[Ano92a] Anonymous. Author index for volume 13. Journal of Algorithms, 13(4):700–701, December 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749290065K.

Anonymous:1992:EB

[Ano92b] Anonymous. Editorial Board. Journal of Algorithms, 13(1): ??, March 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290001S.

Anonymous:1992:PAFa

[Ano92c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 13(1):175–176, March 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900133.

Anonymous:1992:PAFb

[Ano92d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 13(2):350–351, June 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900236.

Anonymous:1992:PAFc

[Ano92e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 13(3):525–526, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290053F.

Anonymous:1992:PAFd

[Ano92f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 13(4):698–699, December 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290064J.

Anonymous:1993:AIVa

[Ano93a] Anonymous. Author index for volume 14. Journal of Algorithms, 14(3):490, May 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710278.

Anonymous:1993:AIVb

[Ano93b] Anonymous. Author index for volume 15. Journal of Algorithms, 15(3):504, November 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677483710515.

Anonymous:1994:AIVa

[Ano94a] Anonymous. Author index for volume 16. Journal of Algorithms, 16(3):493, May 1994. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677484710236.

Anonymous:1994:AIVb

[Ano94b] Anonymous. Author index for volume 17. Journal of Algorithms, 17(3):504, November 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710467.

Anonymous:1995:AIVa

[Ano95a] Anonymous. Author index for volume 18. Journal of Algorithms, 18(3):651, May 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710243.

Anonymous:1995:AIVb

[Ano95b] Anonymous. Author index for volume 19. Journal of Algorithms, 19(3):505, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677485710498.

Anonymous:1996:AIV

[Ano96a] Anonymous. Author index for volume 20. Journal of Algorithms, 20(3):633, May 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900322.

Anonymous:1996:CIV

[Ano96b] Anonymous. Cumulative index for volumes 1-20. Journal of Algorithms, 20(3):634–660, May 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900334.

Anonymous:1996:PAFa

[Ano96c] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 20(1):202, January 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900097.

Anonymous:1996:PAFb

[Ano96d] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 20(3):633, May 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900310.

Anonymous:1996:PAFc

[Ano96e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 21(3):663, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900656.

Anonymous:1997:AIVa

[Ano97a] Anonymous. Author index for volume 22. Journal of Algorithms, 22(2):393, February 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496908469.

Anonymous:1997:AIVb

[Ano97b] Anonymous. Author index for volume 23. Journal of Algorithms, 23(2):405, May 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908580.

Anonymous:1997:AIVc

[Ano97c] Anonymous. Author index for volume 24. Journal of Algorithms, 24(2):422, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908798.

Anonymous:1997:AIVd

[Ano97d] Anonymous. Author index for volume 25. Journal of Algorithms, 25(2):360, November 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497909019.

Anonymous:1997:PAFa

[Ano97e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 22(1):197, January 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496908342.

Anonymous:1997:PAFb

[Ano97f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 22(2):392, February 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496908457.

Anonymous:1997:PAFc

[Ano97g] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 23(1):206, April 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908579.

Anonymous:1997:PAFd

[Ano97h] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 23(2):404, May 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908592.

Anonymous:1997:PAFe

[Ano97i] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 24(1):221, July 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908701.

Anonymous:1997:PAFf

[Ano97j] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 24(2):421, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908786.

Anonymous:1997:PAFg

[Ano97k] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 25(1):203, October 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908907.

Anonymous:1997:PAFh

[Ano97l] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 25(2):359, November 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909007.

Anonymous:1998:AIVa

[Ano98a] Anonymous. Author index for volume 26. Journal of Algorithms, 26(2):414, February 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909270.

Anonymous:1998:AIVb

[Ano98b] Anonymous. Author index for volume 27. Journal of Algorithms, 27(2):358, May 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909415.

Anonymous:1998:AIVc

[Ano98c] Anonymous. Author index for volume 28. Journal of Algorithms, 28(2):367, August 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909506.

Anonymous:1998:AIVd

[Ano98d] Anonymous. Author index for volume 29. Journal of Algorithms, 29(2):412, November 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909786.

Anonymous:1998:FSA

[Ano98e] Anonymous. A fast and simple algorithm for identifying 2monotonic positive Boolean functions. Journal of Algorithms, 26(2):291–305, February 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908968.

Anonymous:1998:IRT

[Ano98f] Anonymous. Interval routing on k-trees. Journal of Algorithms, 26(2):325–369, February 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908804.

Anonymous:1998:PAFa

[Ano98g] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 26(1):??, January 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497909123.

Anonymous:1998:PAFb

[Ano98h] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 26(2):413, February 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909269.

Anonymous:1998:PAFc

[Ano98i] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 27(1):167, April 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909336.

Anonymous:1998:PAFd

[Ano98j] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 27(2):357, May 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909403.

Anonymous:1998:PAFe

[Ano98k] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 28(1):196, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909452.

Anonymous:1998:PAFf

[Ano98l] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 28(2):366, August 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749890949X.

Anonymous:1998:PAFg

[Ano98m] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 29(1):201, October 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909750.

Anonymous:1999:AIVa

[Ano99a] Anonymous. Author index for volume 30. Journal of Algorithms, 30(2):455, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498910033.

Anonymous:1999:AIVb

[Ano99b] Anonymous. Author index for volume 31. Journal of Algorithms, 31(2):356, May 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910192.

Anonymous:1999:AIVc

[Ano99c] Anonymous. Author index for volume 33. Journal of Algorithms, 33(2):320, November 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749991057X.

Anonymous:1999:E

[Ano99d] Anonymous. ERRATUM. Journal of Algorithms, 32(1):62, July 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910271.

Anonymous:1999:F

[Ano99e] Anonymous. Foreword. Journal of Algorithms, 30(2):??, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910040.

Anonymous:1999:PAFa

[Ano99f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 30(1):209, January 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909920.

Anonymous:1999:PAFb

[Ano99g] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 30(2):454, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498910021.

Anonymous:1999:PAFc

[Ano99h] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 31(1):269, April 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749991009X.

Anonymous:1999:PAFd

[Ano99i] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 31(2):355, May 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910180.

Anonymous:1999:PAFe

[Ano99j] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 32(1):63, July 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910283.

Anonymous:1999:PAFf

[Ano99k] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 33(1):185, October 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910519.

Anonymous:1999:PAFg

[Ano99l] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 33(2):319, November 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910568.

Anonymous:2000:AIVa

[Ano00a] Anonymous. Author index for volume 34. Journal of Algorithms, 34(2):398, February 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910660.

Anonymous:2000:AIVb

[Ano00b] Anonymous. Author index for volume 35. Journal of Algorithms, 35(2):261, May 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400910848.

Anonymous:2000:AIVc

[Ano00c] Anonymous. Author index for volume 36. Journal of Algorithms, 36(2):277, August 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740091112X.

Anonymous:2000:AIVd

[Ano00d] Anonymous. Author index for volume 37. Journal of Algorithms, 37(2):579–580, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740091126X.

Anonymous:2000:PAFa

[Ano00e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 34(1):202, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910659.

Anonymous:2000:PAFb

[Ano00f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 34(2):397, February 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910672.

Anonymous:2000:PAFc

[Ano00g] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 35(1):127, April 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400910824.

Anonymous:2000:PAFd

[Ano00h] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 35(2):260, May 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400910836.

Anonymous:2000:PAFe

[Ano00i] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 36(1):117, July 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911027.

Anonymous:2000:PAFf

[Ano00j] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 36(2):276, August 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911118.

Anonymous:2000:PAFg

[Ano00k] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 37(1):236, October 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911234.

Anonymous:2000:PAFh

[Ano00l] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 37(2):578, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911258.

Anonymous:2001:AIVa

[Ano01a] Anonymous. Author index for volume 38. Journal of Algorithms, 38(2):467, February 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911599.

Anonymous:2001:AIVb

[Ano01b] Anonymous. Author index for volume 39. Journal of Algorithms, 39(2):242, May 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911666.

Anonymous:2001:AIVc

[Ano01c] Anonymous. Author index for volume 40. Journal of Algorithms, 40(2):235, August 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911812.

Anonymous:2001:AIVd

[Ano01d] Anonymous. Author index for volume 41. Journal of Algorithms, 41(2):472–473, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401912088.

Anonymous:2001:PAFa

[Ano01e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 38(1):334, January 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911568.

Anonymous:2001:PAFb

[Ano01f] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 38(2):466, February 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911575.

Anonymous:2001:PAFc

[Ano01g] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 39(1):135, April 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911642.

Anonymous:2001:PAFd

[Ano01h] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 39(2):241, May 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911654.

Anonymous:2001:PAFe

[Ano01i] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 40(1):134, July 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911794.

Anonymous:2001:PAFf

[Ano01j] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 40(2):234, August 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911800.

Anonymous:2001:PAFg

[Ano01k] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 41(1):114, October 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911903.

Anonymous:2001:PAFh

[Ano01l] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 41(2):471, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401912076.

Anonymous:2002:AIVa

[Ano02a] Anonymous. Author index for volume 42. Journal of Algorithms, 42(2):368, February 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402912277.

Anonymous:2002:AIVb

[Ano02b] Anonymous. Author index for volume 43. Journal of Algorithms, 43(2):285, May 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402000093.

Anonymous:2002:AIVc

[Ano02c] Anonymous. Author index for volume 44. Journal of Algorithms, 44(2):380, August 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002407.

Anonymous:2002:AIVd

[Ano02d] Anonymous. Author index for volume 45. Journal of Algorithms, 45(2):214, November 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740200278X.

Anonymous:2002:EBa

[Ano02e] Anonymous. Editorial Board. Journal of Algorithms, 44(1):??, July 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402002274.

Anonymous:2002:EBb

[Ano02f] Anonymous. Editorial Board. Journal of Algorithms, 44(2): ??, August 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002365.

Anonymous:2002:EBc

[Ano02g] Anonymous. Editorial Board. Journal of Algorithms, 45(1): ??, October 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002626.

Anonymous:2002:EBd

[Ano02h] Anonymous. Editorial Board. Journal of Algorithms, 45(2):??, November 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002754.

Anonymous:2002:EVN

[Ano02i] Anonymous. Erratum: Volume 41, Number 2 (2001), in the article “Approximation Algorithms for the Achromatic Number,” by Amitabh Chaudhary and Sundar Vishwanathan, pages 404–416. Journal of Algorithms, 43(1):153, April 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740291223X. See [CV01].

Anonymous:2002:PAFa

[Ano02j] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 42(1):203, January 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740191209X.

Anonymous:2002:PAFb

[Ano02k] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 42(2):367, February 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402912265.

Anonymous:2002:PAFc

[Ano02l] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 43(1):154, April 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402912319.

Anonymous:2002:PAFd

[Ano02m] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 43(2):284, May 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402000081.

Anonymous:2002:PAFe

[Ano02n] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 44(1):286, July 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002304.

Anonymous:2002:PAFf

[Ano02o] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 44(2):379, August 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002390.

Anonymous:2002:PAFg

[Ano02p] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 45(1):91, October 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002651.

Anonymous:2002:PAFh

[Ano02q] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 45(2):213, November 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002985.

Anonymous:2003:AIVa

[Ano03a] Anonymous. Author index for volume 46. Journal of Algorithms, 46(2):191, February 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000257.

Anonymous:2003:AIVb

[Ano03b] Anonymous. Author index for volume 47. Journal of Algorithms, 47(2):136, July 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000646.

Anonymous:2003:AIVc

[Ano03c] Anonymous. Author index for volume 48. Journal of Algorithms, 48(2):453, September 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001184.

Anonymous:2003:AIVd

[Ano03d] Anonymous. Author index for volume 49. Journal of Algorithms, 49(2):305, November 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001457.

Anonymous:2003:EBa

[Ano03e] Anonymous. Editorial Board. Journal of Algorithms, 46(1): ??, January 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000075.

Anonymous:2003:EBb

[Ano03f] Anonymous. Editorial Board. Journal of Algorithms, 46(2):??, February 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667740300021X.

Anonymous:2003:EBc

[Ano03g] Anonymous. Editorial Board. Journal of Algorithms, 47(1): ??, April 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000348.

Anonymous:2003:EBd

[Ano03h] Anonymous. Editorial Board. Journal of Algorithms, 47(2):??, July 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000609.

Anonymous:2003:EBe

[Ano03i] Anonymous. Editorial Board. Journal of Algorithms, 48(1): ??, August 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001019.

Anonymous:2003:EBf

[Ano03j] Anonymous. Editorial Board. Journal of Algorithms, 48(2):??, September 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001147.

Anonymous:2003:EBg

[Ano03k] Anonymous. Editorial Board. Journal of Algorithms, 49(1): ??, October 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001330.

Anonymous:2003:EBh

[Ano03l] Anonymous. Editorial Board. Journal of Algorithms, 49(2):??, November 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001421.

Anonymous:2003:PAFa

[Ano03m] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 46(1):96, January 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000105.

Anonymous:2003:PAFb

[Ano03n] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 46(2):190, February 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000245.

Anonymous:2003:PAFc

[Ano03o] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 47(1):62, April 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000373.

Anonymous:2003:PAFd 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000634.

Anonymous:2003:PAFe

[Ano03q] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 48(1):271, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001044.

Anonymous:2003:PAFf

[Ano03r] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 48(2):452, September 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001172.

Anonymous:2003:PAFg

[Ano03s] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 49(1):217, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001354.

Anonymous:2003:PAFh

[Ano03t] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 49(2):304, November 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001445.

Anonymous:2003:PN

[Ano03u] Anonymous. Publisher’s note. Journal of Algorithms, 46(1):iii– iv, January 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000142.

Anonymous:2004:AIVa

[Ano04a] Anonymous. Author index for volume 50. Journal of Algorithms, 50(2):277, February 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000240.

Anonymous:2004:AIVb

[Ano04b] Anonymous. Author index for volume 51. Journal of Algorithms, 51(2):217, May 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000537.

Anonymous:2004:AIVc

[Ano04c] Anonymous. Author index for volume 52. Journal of Algorithms, 52(2):208, August 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000963.

Anonymous:2004:AIVd

[Ano04d] Anonymous. Author index for volume 53. Journal of Algorithms, 53(2):208, November 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001324.

Anonymous:2004:EBa

[Ano04e] Anonymous. Editorial Board. Journal of Algorithms, 50(1): ??, January 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001706.

Anonymous:2004:EBb

[Ano04f] Anonymous. Editorial Board. Journal of Algorithms, 50(2):??, February 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000033.

Anonymous:2004:EBc

[Ano04g] Anonymous. Editorial Board. Journal of Algorithms, 51(1): ??, April 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000318.

Anonymous:2004:EBd

[Ano04h] Anonymous. Editorial Board. Journal of Algorithms, 51(2):??, May 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404000495.

Anonymous:2004:EBe

[Ano04i] Anonymous. Editorial Board. Journal of Algorithms, 52(1):??, July 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740400080X.

Anonymous:2004:EBf

[Ano04j] Anonymous. Editorial Board. Journal of Algorithms, 52(2): ??, August 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000938.

Anonymous:2004:EBg

[Ano04k] Anonymous. Editorial Board. Journal of Algorithms, 53(1): ??, October 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001221.

Anonymous:2004:PAFa

[Ano04l] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 50(1):132, January 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001731.

Anonymous:2004:PAFb

[Ano04m] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 50(2):276, February 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000069.

Anonymous:2004:PAFc

[Ano04n] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 51(1):105, April 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000343.

Anonymous:2004:PAFd

[Ano04o] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 51(2):216, May 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000525.

Anonymous:2004:PAFe 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000835.

Anonymous:2004:PAFf

[Ano04q] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 52(2):207, August 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000951.

Anonymous:2004:PAFg

[Ano04r] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 53(1):113, October 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001245.

Anonymous:2004:PAFh

[Ano04s] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 53(2):207, November 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001312.

Anonymous:2005:AIVa

[Ano05a] Anonymous. Author index for volume 54. Journal of Algorithms, 54(2):223, February 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000088.

Anonymous:2005:AIVb

[Ano05b] Anonymous. Author index for volume 55. Journal of Algorithms, 55(2):204, May 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405000453.

Anonymous:2005:AIVc

[Ano05c] Anonymous. Author index for volume 56. Journal of Algorithms, 56(2):155, August 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000738.

Anonymous:2005:AIVd

[Ano05d] Anonymous. Author index for volume 57. Journal of Algorithms, 57(2):149, November 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405001021.

Anonymous:2005:EBa

[Ano05e] Anonymous. Editorial Board. Journal of Algorithms, 54(1): ??, January 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001609.

Anonymous:2005:EBb

[Ano05f] Anonymous. Editorial Board. Journal of Algorithms, 54(2):??, February 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405000040.

Anonymous:2005:EBc

[Ano05g] Anonymous. Editorial Board. Journal of Algorithms, 55(1): ??, April 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405000295.

Anonymous:2005:EBd

[Ano05h] Anonymous. Editorial Board. Journal of Algorithms, 55(2):??, May 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677405000416.

Anonymous:2005:EBe

[Ano05i] Anonymous. Editorial Board. Journal of Algorithms, 56(1):??, July 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740500057X.

Anonymous:2005:EBf

[Ano05j] Anonymous. Editorial Board. Journal of Algorithms, 56(2): ??, August 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405000696.

Anonymous:2005:EBg

[Ano05k] Anonymous. Editorial Board. Journal of Algorithms, 57(1):??, September 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405000878.

Anonymous:2005:EBh

[Ano05l] Anonymous. Editorial Board. Journal of Algorithms, 57(2):??, November 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405000982.

Anonymous:2005:PAFa

[Ano05m] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 54(1):126, January 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001622.

Anonymous:2005:PAFb

[Ano05n] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 54(2):222, February 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000076.

Anonymous:2005:PAFc

[Ano05o] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 55(1):99, April 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000325.

Anonymous:2005:PAFd

[Ano05p] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 55(2):203, May 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000441.

Anonymous:2005:PAFe

[Ano05q] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 56(1):76, July 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740500060X.

Anonymous:2005:PAFf

[Ano05r] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 56(2):154, August 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000726.

Anonymous:2005:PAFg

[Ano05s] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 57(1):74, September 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000908.

Anonymous:2005:PAFh

[Ano05t] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 57(2):148, November 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740500101X.

Anonymous:2006:AIVa

[Ano06a] Anonymous. Author index for volume 58. Journal of Algorithms, 58(2):166, February 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740500129X.

Anonymous:2006:AIVb

[Ano06b] Anonymous. Author index for volume 59. Journal of Algorithms, 59(2):163, May 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677406000198.

Anonymous:2006:AIVc

[Ano06c] Anonymous. Author index for volume 60. Journal of Algorithms, 60(2):169, August 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677406000460.

Anonymous:2006:AIVd

[Ano06d] Anonymous. Author index for volume 61. Journal of Algorithms, 61(2):90, November 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677406000708.

Anonymous:2006:EBa

[Ano06e] Anonymous. Editorial Board. Journal of Algorithms, 58(1): ??, January 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405001161.

Anonymous:2006:EBb

[Ano06f] Anonymous. Editorial Board. Journal of Algorithms, 58(2):??, February 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677405001252.

Anonymous:2006:EBc

[Ano06g] Anonymous. Editorial Board. Journal of Algorithms, 59(1): ??, April 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677406000034.

Anonymous:2006:EBd

[Ano06h] Anonymous. Editorial Board. Journal of Algorithms, 59(2):??, May 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677406000150.

Anonymous:2006:EBe

[Ano06i] Anonymous. Editorial Board. Journal of Algorithms, 60(1):??, July 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677406000307.

Anonymous:2006:EBf

[Ano06j] Anonymous. Editorial Board. Journal of Algorithms, 60(2): ??, August 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677406000423.

Anonymous:2006:EBg

[Ano06k] Anonymous. Editorial Board. Journal of Algorithms, 61(1):??, September 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677406000575.

Anonymous:2006:EBh

[Ano06l] Anonymous. Editorial Board. Journal of Algorithms, 61(2):??, November 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677406000666.

Anonymous:2006:PAFa

[Ano06m] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 58(1):79, January 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405001197.

Anonymous:2006:PAFb

[Ano06n] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 58(2):165, February 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405001288.

Anonymous:2006:PAFc

[Ano06o] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 59(1):79, April 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740600006X.

Anonymous:2006:PAFd

[Ano06p] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 59(2):162, May 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677406000186.

Anonymous:2006:PAFe

[Ano06q] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 60(1):84, July 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677406000332.

Anonymous:2006:PAFf

[Ano06r] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 60(2):168, August 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677406000459.

Anonymous:2006:PAFg

[Ano06s] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 61(1):45, September 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677406000605.

Anonymous:2006:PAFh

[Ano06t] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 61(2):89, November 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677406000691.

Anonymous:2007:AIV

[Ano07a] Anonymous. Author index for volume 62. Journal of Algorithms, 62(3–4):168, July/October 2007. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677407000375.

Anonymous:2007:EBa

[Ano07b] Anonymous. Editorial Board. Journal of Algorithms, 62(1): ??, January 2007. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667740700003X.

Anonymous:2007:EBb

[Ano07c] Anonymous. Editorial Board. Journal of Algorithms, 62(2): ??, April 2007. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677407000144.

Anonymous:2007:EBc

[Ano07d] Anonymous. Editorial Board. Journal of Algorithms, 62(3–4):??, July/October 2007. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667740700034X.

Anonymous:2007:PAF

[Ano07e] Anonymous. Papers to appear in forthcoming issues. Journal of Algorithms, 62(1):47, January 2007. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677407000065.

Anonymous:2008:EBa

[Ano08a] Anonymous. Editorial Board. Journal of Algorithms, 63(1–3):??, January/July 2008. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677408000126.

Anonymous:2008:EBb

[Ano08b] Anonymous. Editorial Board. Journal of Algorithms, 63(4): ??, October 2008. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677408000357.

Anonymous:2009:EBa

[Ano09a] Anonymous. Editorial Board. Journal of Algorithms, 64(1): ??, January 2009. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677409000121.

Anonymous:2009:EBb

[Ano09b] Anonymous. Editorial Board. Journal of Algorithms, 64(2–3): ??, April/July 2009. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677409000261.

Anonymous:2009:EBc

[Ano09c] Anonymous. Editorial Board. Journal of Algorithms, 64(4): ??, October 2009. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667740900039X.

Azar:1995:CLA

[ANR95] Y. Azar, J. Naor, and R. Rom. The competitiveness of online assignments. Journal of Algorithms, 18(2):221–237, March 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710085.

Anstee:1985:APT

[Ans85] R. P. Anstee. An algorithmic proof of Tutte’s f-factor theorem. Journal of Algorithms, 6(1):112–131, March 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900227.

Aroya:1997:RST

[ANS97] Ishai Ben Aroya, Ilan Newman, and Assaf Schuster. Randomized single-target hot-potato routing. Journal of Algorithms, 23(1): 101–120, April 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908111.

Ahuja:2000:FAI

[AO00] Ravindra K. Ahuja and James B. Orlin. A faster algorithm for the inverse spanning tree problem. Journal of Algorithms, 34(1):177– 193, January 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910520.

Arkin:1986:CC

[AP86] Esther M. Arkin and Christos H. Papadimitriou. On the complexity of circulations. Journal of Algorithms, 7(1):134–145, March 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677486900428.

Adhar:1990:PAC

[AP90] Gur Saran Adhar and Shietung Peng. Parallel algorithms for cographs and parity graphs with applications. Journal of Algorithms, 11(2):252–284, June 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090006Z.

Andersson:1998:AIL

[AP98] Arne Andersson and Ola Petersson. Approximate indexed lists. Journal of Algorithms, 29(2):256–276, November 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909518.

Agarwal:2003:AAP

[AP03] Pankaj K. Agarwal and Cecilia M. Procopiuc. Approximation algorithms for projective clustering. Journal of Algorithms, 46(2):115–139, February 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740200295X.

Auletta:1998:PRG

[APP98] Vincenzo Auletta, Domenico Parente, and Giuseppe Persiano. Placing resources on a growing line. Journal of Algorithms, 26 (1):87–100, 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

Afek:1993:SPR

[AR93] Y. Afek and M. Ricklin. Sparser: a paradigm for running distributed algorithms. Journal of Algorithms, 14(2):316–328, March 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710163.

Allauzen:2000:SOS

[AR00] Cyril Allauzen and Mathieu Raffinot. Simple optimal string matching algorithm. Journal of Algorithms, 36(1):102–116, July 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400910873.

Adler:2005:TBP

[AR05] Micah Adler and Adi Rosén. Tight bounds for the performance of Longest In System on DAGs. Journal of Algorithms, 55(2): 101–112, May 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000720.

Aiello:1998:DPP

[ARV98] William Aiello, S. Raj Rajagopalan, and Ramarathnam Venkatesan. Design of practical and provably good random number generators. Journal of Algorithms, 29(2):358–389, November 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749890952X.

Aspvall:1980:KLP

[AS80] Bengt Aspvall and Richard E. Stone. Khachiyan’s linear programming algorithm. Journal of Algorithms, 1(1):1–13, March 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677480900024.

Attiya:1991:BCA

[AS91] Hagit Attiya and Marc Snir. Better computing on the anonymous ring. Journal of Algorithms, 12(2):204–238, June 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190002G.

Apostolico:1992:SAW

[AS92] Alberto Apostolico and Wojciech Szpankowski. Self-alignments in words and their applications. Journal of Algorithms, 13(3):446– 467, September 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290049I.

Agarwal:1993:CSS

[AS93] P. K. Agarwal and M. Sharir. Circle shooting in a simple polygon. Journal of Algorithms, 14(1):69–87, January 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710047.

Agarwal:1996:SMM

[AS96a] Pankaj K. Agarwal and Sandeep Sen. Selection in monotone matrices and computing kth nearest neighbors. Journal of Algorithms, 20(3):581–601, May 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900280.

Agarwal:1996:RSA

[AS96b] Pankaj K. Agarwal and Micha Sharir. Ray shooting amidst convex polygons in 2D. Journal of Algorithms, 21(3):508–519, November 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900565.

Afek:1997:OTS

[AS97] Yehuda Afek and Gideon Stupp. Optimal time-space tradeoff for shared memory leader election. Journal of Algorithms, 25(1):95– 117, October 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908524.

Alon:1999:TSP

[AS99] Noga Alon and Benny Sudakov. On two segmentation problems. Journal of Algorithms, 33(1):173–184, October 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910246.

Alon:2003:TS

[AS03] Noga Alon and Asaf Shapira. Testing satisfiability. Journal of Algorithms, 47(2):87–103, July 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000191.

Aspvall:1980:RDI

[Asp80] Bengt Aspvall. Recognizing disguised NR(1) instances of the satisfiability problem. Journal of Algorithms, 1(1):97–103, March 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677480900073.

Aspnes:1993:TSE

[Asp93] J. Aspnes. Time- and space-efficient randomized consensus. Journal of Algorithms, 14(3):414–431, May 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710229.

Aspnes:2002:FDC

[Asp02] James Aspnes. Fast deterministic consensus in a noisy environment. Journal of Algorithms, 45(1):16–39, October 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002201.

Agarwal:1994:APS

[AST94] P. K. Agarwal, M. Sharir, and S. Toledo. Applications of parametric searching in geometric optimization. Journal of Algorithms, 17(3):292–318, November 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710388.

Aggarwal:1995:IAT

[AT95] A. Aggarwal and T. Tokuyama. An improved algorithm for the Traveler’s Problem. Journal of Algorithms, 19(2):318–330, September 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710395.

Alstrup:2000:OPA

[AT00] Stephen Alstrup and Mikkel Thorup. Optimal pointer algorithms for finding nearest common ancestors in dynamic trees. Journal of Algorithms, 35(2):169–188, May 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910794.

Atallah:1986:CCH

[Ata86] Mikhail J. Atallah. Computing the convex hull of line intersections. Journal of Algorithms, 7(2):285–288, June 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900106.

Atkinson:1987:OAG

[Atk87] M. D. Atkinson. An optimal algorithm for geometrical congruence. Journal of Algorithms, 8(2):159–172, June 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900368.

Attiya:2000:ERS

[Att00] Hagit Attiya. Efficient and robust sharing of memory in messagepassing systems. Journal of Algorithms, 34(1):109–127, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910258.

Alborzi:2001:CP

[ATUW01] Houman Alborzi, Eric Torng, Patchrawat Uthaisombut, and Stephen Wagner. The k-client problem. Journal of Algorithms, 41(2):115–173, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911824.

Aurenhammer:1988:IAD

[Aur88] F. Aurenhammer. Improved algorithms for discs and balls using power diagrams. Journal of Algorithms, 9(2):151–161, June 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900351.

Avery:1992:APS

[Ave92] Peter Avery. An algorithmic proof that semiorders are representable. Journal of Algorithms, 13(1):144–147, March 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290010A.

Agarwal:1993:IQC

[AVO93] P. K. Agarwal, M. Vankreveld, and M. Overmars. Intersection queries in curved objects. Journal of Algorithms, 15(2):229–266, September 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710400.

Afek:1999:ISC

[AW99] Yehuda Afek and Eytan Weisberger. The instancy of snapshots and commuting objects. Journal of Algorithms, 30(1):68– 105, January 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909695.

Asano:2002:IAA

[AW02] Takao Asano and David P. Williamson. Improved approximation algorithms for MAX SAT. Journal of Algorithms, 42(1):173– 202, January 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401912027.

Aspnes:2005:CCD

[AW05] James Aspnes and Orli Waarts. Compositional competitiveness for distributed algorithms. Journal of Algorithms, 54(2):127– 151, February 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001038.

Awerbuch:1990:EFD

[Awe90] Baruch Awerbuch. On the effects of feedback in dynamic network protocols. Journal of Algorithms, 11(3):342–373, September 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749090018A.

Andrews:2004:MEE

[AZ04] Matthew Andrews and Lisa Zhang. Minimizing end-to-end delay in high-speed networks with a simple coordinated schedule. Journal of Algorithms, 52(1):57–81, July 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000458.

Baker:1985:NPF

[Bak85] Brenda S. Baker. A new proof for the first-fit decreasing binpacking algorithm. Journal of Algorithms, 6(1):49–70, March 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677485900185.

Bergeron:1989:ACU

[BBBD89] F. Bergeron, J. Berstel, S. Brlek, and C. Duboc. Addition chains using continued fractions. Journal of Algorithms, 10(3):403–412, September 1989. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677489900369.

Biedl:2001:EAP

[BBDL01] Therese C. Biedl, Prosenjit Bose, Erik D. Demaine, and Anna Lubiw. Efficient algorithms for Petersen’s matching theorem. Journal of Algorithms, 38(1):110–134, January 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911325.

Bocker:2000:AAT

[BBDS00] Sebastian Böcker, David Bryant, Andreas W. M. Dress, and Mike A. Steel. Algorithmic aspects of tree amalgamation. Journal of Algorithms, 37(2):522–537, November 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911167.

Bareli:1994:ONR

[BBFY94] E. Bareli, P. Berman, A. Fiat, and P. Y. Yan. Online navigation in a Room. Journal of Algorithms, 17(3):319–341, November 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667748471039X.

Berry:2006:WRA

[BBH^{+}06] Anne Berry, Jean-Paul Bordat, Pinar Heggernes, Geneviève Simonet, and Yngve Villanger. A wide-range algorithm for minimal triangulation from an arbitrary ordering. Journal of Algorithms, 58(1):33–66, January 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001142.

Baker:1981:ATD

[BBK81] Brenda S. Baker, Donna J. Brown, and Howard P. Katseff. A 54 algorithm for two-dimensional packing. Journal of Algorithms, 2(4):348–368, December 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900341.

Baker:1982:TDP

[BC82] B. S. Baker and E. G. Coffman, Jr. A two-dimensional binpacking model of preemptive, FIFO storage allocation. Journal of Algorithms, 3(4):303–316, December 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900281.

Berger:1995:SCB

[BC95] B. Berger and L. Cowen. Scheduling with concurrency-based constraints. Journal of Algorithms, 18(1):98–123, January 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710036.

Bender:1999:APM

[BC99] Edward A. Bender and E. Rodney Canfield. An approximate probabilistic model for structured Gaussian elimination. Journal of Algorithms, 31(2):271–290, May 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910088.

Brandstadt:1999:DAT

[BCD99] Andreas Brandstädt, Victor Chepoi, and Feodor Dragan. Distance approximating trees for chordal and dually chordal graphs. Journal of Algorithms, 30(1):166–184, January 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909622.

Burkov:2009:ELP

[BCd09] Andriy Burkov and Brahim Chaib-draa. Effective learning in the presence of adaptive counterparts. Journal of Algorithms, 64(4):127–138, October 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677409000340.

Brinkmann:2002:CEF

[BCH02] Gunnar Brinkmann, Gilles Caporossi, and Pierre Hansen. A constructive enumeration of fusenes and benzenoids. Journal of Algorithms, 45(2):155–166, November 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002158.

Berman:2000:LLB

[BCK00] Piotr Berman, Moses Charikar, and Marek Karpinski. On-line load balancing for related machines. Journal of Algorithms, 35(1): 108–121, April 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910702.

Breslauer:1998:CCS

[BCT98] Dany Breslauer, Livio Colussi, and Laura Toniolo. On the comparison complexity of the string prefix-matching problem. Journal of Algorithms, 29(1):18–67, October 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909488.

Brinkmann:1997:CEF

[BD97] Gunnar Brinkmann and Andreas W. M. Dress. A constructive enumeration of fullerenes. Journal of Algorithms, 23(2):345–358, May 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908068.

Borchers:1998:SRR

[BDGW98] Al Borchers, Ding-Zhu Du, Biao Gao, and Pengjun Wan. The ksteiner ratio in the rectilinear plane. Journal of Algorithms, 29(1): 1–17, October 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909464.

Bender:2004:LPC

[BDIW04] Michael A. Bender, Ziyang Duan, John Iacono, and Jing Wu. A locality-preserving cache-oblivious dynamic dictionary. Journal of Algorithms, 53(2):115–136, November 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000707.

Berman:2003:AAM

[BDM03] Piotr Berman, Bhaskar DasGupta, and S. Muthukrishnan. Approximation algorithms for MAX-MIN tiling. Journal of Algorithms, 47(2):122–134, July 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000154.

Berman:2001:EAA

[BDMR01] Piotr Berman, Bhaskar DasGupta, S. Muthukrishnan, and Suneeta Ramaswami. Efficient approximation algorithms for tiling and packing problems with rectangles. Journal of Algorithms, 41(2):443–470, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911885.

Berman:1997:RBL

[BDP97] Piotr Berman, Krzysztof Diks, and Andrzej Pelc. Reliable broadcasting in logarithmic time with Byzantine link failures.

Journal of Algorithms, 22(2):199–211, February 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749690810X.

Brakerski:2003:DPP

[BDPS03] Zvika Brakerski, Vladimir Dreizin, and Boaz Patt-Shamir. Dispatching in perfectly-periodic schedules. Journal of Algorithms, 49(2):219–239, November 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000890.

Below:2004:CFS

[BDRG04] Alexander Below, Jesús A. De Loera, and Jürgen Richter-Gebert. The complexity of finding small triangulations of convex 3-polytopes. Journal of Algorithms, 50(2):134–167, February 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000920.

Bach:1993:FR

[BDS93] E. Bach, J. Driscoll, and J. Shallit. Factor refinement. Journal of Algorithms, 15(2):199–222, September 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710382.

Bourdon:2002:DAE

[BDV02] Jérémie Bourdon, Benoit Daireaux, and Brigitte Vallée. Dynamical analysis of α-Euclidean algorithms. Journal of Algorithms, 44(1):246–285, July 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002183.

Bodlaender:1997:DT

[BE97] Hans L. Bodlaender and Joost Engelfriet. Domino treewidth. Journal of Algorithms, 24(1):94–123, July 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908548.

Beigel:2005:CTI

[BE05] Richard Beigel and David Eppstein. 3-coloring in time O(n^{1.3289}). Journal of Algorithms, 54(2):168–204, February 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001117.

Bedregal:2007:NFW

[Bed07] Benjamín Callejas Bedregal. A normal form which preserves tautologies and contradictions in a class of fuzzy logics. Journal of Algorithms, 62(3–4):135–147, July/October 2007. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677407000314.

Bampis:1995:PRH

[BEMS95] E. Bampis, M. Elhaddad, Y. Manoussakis, and M. Santha. A parallel reduction of Hamiltonian cycle to Hamiltonian path in tournaments. Journal of Algorithms, 19(3):432–440, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710450.

Bentley:1980:PAC

[Ben80] Jon Louis Bentley. A parallel algorithm for constructing minimum spanning trees. Journal of Algorithms, 1(1):51–59, March 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677480900048.

Bernstein:2005:FCE

[Ber05] Daniel J. Bernstein. Factoring into coprimes in essentially linear time. Journal of Algorithms, 54(1):1–30, January 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000732.

Bespamyatnikh:2003:CHS

[Bes03] Sergei Bespamyatnikh. Computing homotopic shortest paths in the plane. Journal of Algorithms, 49(2):284–303, November 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000907.

Beynon:1983:FAS

[Bey83] W. M. Beynon. A formal account of some elementary continued fraction algorithms. Journal of Algorithms, 4(3):221–240, September 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900238.

Boshernitzan:1984:LAN

[BF84] M. Boshernitzan and A. S. Fraenkel. A linear algorithm for nonhomogeneous spectra of numbers. Journal of Algorithms, 5(2): 187–198, June 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900269.

Bodlaender:2002:APO

[BF02] Hans L. Bodlaender and Fedor V. Fomin. Approximation of pathwidth of outerplanar graphs. Journal of Algorithms, 43(2):190– 200, May 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402000019.

Bender:2005:LCA

[BFCP^{+}05] Michael A. Bender, Martín Farach-Colton, Giridhar Pemmasani, Steven Skiena, and Pavel Sumazin. Lowest common ancestors in trees and directed acyclic graphs. Journal of Algorithms, 57(2):75–94, November 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000854.

Bollobas:1996:BCH

[BFF96] B. Bollobás, T. I. Fenner, and A. M. Frieze. On the best case of Heapsort. Journal of Algorithms, 20(2):205–217, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900115. For other analyses of Heap-

sort, see [LP93, SS93].

Becker:1996:ESO

[BFG^{+}96] Bruno Becker, Paolo Giulio Franciosa, Stephan Gschwind, Stefano Leonardi, Thomas Ohler, and Peter Widmayer. Enclosing a set of objects by two minimum area rectangles. Journal of Algorithms, 21(3):520–541, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900577.

Baker:1986:PCU

[BFM86] B. S. Baker, S. J. Fortune, and S. R. Mahaney. Polygon containment under translation. Journal of Algorithms, 7(4):532–548, December 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900179.

Bradford:1997:MEP

[BFS97] Phillip G. Bradford, Rudolf Fleischer, and Michiel Smid. More efficient parallel totally monotone matrix searching. Journal of Algorithms, 23(2):386–400, May 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749690824X.

Bertsimas:1999:AOA

[BG99] Dimitris Bertsimas and David Gamarnik. Asymptotically optimal algorithms for job shop scheduling and packet routing. Journal of Algorithms, 33(2):296–318, November 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910477.

Baker:2002:SDP

[BG02] Brenda S. Baker and Raffaele Giancarlo. Sparse dynamic programming for longest common subsequence from fragments. Journal of Algorithms, 42(2):231–254, February 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402912149.

Basch:1999:DSM

[BGH99] Julien Basch, Leonidas J. Guibas, and John Hershberger. Data structures for mobile data. Journal of Algorithms, 31(1):1– 28, April 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909889.

Bodlaender:1995:ATP

[BGHK95] H. L. Bodlaender, J. R. Gilbert, H. Hafsteinsson, and T. Kloks. Approximating treewidth, pathwidth, frontsize, and shortest elimination tree. Journal of Algorithms, 18(2):238–255, March 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710097.

Booth:1999:FAI

[BGLR99] Heather D. Booth, Rajeev Govindan, Michael A. Langston, and Siddharthan Ramachandramurthi. Fast algorithms for K _{4} immersion testing. Journal of Algorithms, 30(2):344–378, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909919.

Bradford:2002:OPF

[BGLR02] Phil Bradford, Mordecai J. Golin, Lawrence L. Larmore, and Wojciech Rytter. Optimal prefix-free codes for unequal letter costs: Dynamic programming with the Monge property. Journal of Algorithms, 42(2):277–303, February 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402912137.

Blackburn:2006:RNP

[BGPGS06] Simon R. Blackburn, Domingo Gomez-Perez, Jaime Gutierrez, and Igor E. Shparlinski. Reconstructing noisy polynomial evaluation in residue rings. Journal of Algorithms, 61(2):47–59, November 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404001154.

Bekesi:1997:GAL

[BGPW97] József Békési, Gábor Galambos, Ulrich Pferschy, and Gerhard J. Woeginger. Greedy algorithms for on-line data compression. Journal of Algorithms, 25(2):274–289, November 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908853.

Brent:1980:FST

[BGY80] Richard P. Brent, Fred G. Gustavson, and David Y. Y. Yun. Fast solution of Toeplitz systems of equations and computation of Padé approximants. Journal of Algorithms, 1(3):259–295, September 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677480900139.

Berthome:2000:SSP

[BHNS00] Pascal Berthomé, Torben Hagerup, Ilan Newman, and Assaf Schuster. Self-simulation for the passive optical star. Journal of Algorithms, 34(1):128–147, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910362.

Barequet:2001:EAM

[BHP01] Gill Barequet and Sariel Har-Peled. Efficiently approximating the minimum-volume bounding box of a point set in three dimensions. Journal of Algorithms, 38(1):91–109, January 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911271.

Baswana:2007:IDA

[BHS07] Surender Baswana, Ramesh Hariharan, and Sandeep Sen. Improved decremental algorithms for maintaining transitive closure and all-pairs shortest paths. Journal of Algorithms, 62(2):74– 92, April 2007. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001452.

Bilardi:2003:P

[BI03] Gianfranco Bilardi and Giuseppe F. Italiano. Preface. Journal of Algorithms, 49(1):1, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000841.

Breslauer:1997:RPS

[BJJ97] Dany Breslauer, Tao Jiang, and Zhigen Jiang. Rotations of periodic strings and short superstrings. Journal of Algorithms, 24(2):340–353, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908610.

Bang-Jensen:2000:SEW

[BJJ00] Jørgen Bang-Jensen and Tibor Jordán. Splitting off edges within a specified subset preserving the edge-connectivity of the graph. Journal of Algorithms, 37(2):326–343, November 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911192.

Berman:1990:GPM

[BJL^{+}90] Fran Berman, David Johnson, Tom Leighton, Peter W. Shor, and Larry Snyder. Generalized planar matching. Journal of Algorithms, 11(2):153–184, June 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090001U.

Baumgarten:1994:DPL

[BJM94] N. Baumgarten, H. Jung, and K. Mehlhorn. Dynamic point location in general subdivisions. Journal of Algorithms, 17(3):342– 380, November 1994. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677484710406.

Bang-Jensen:1992:PAH

[BJMT92] Jørgen Bang-Jensen, Yannis Manoussakis, and Carsten Thomassen. A polynomial algorithm for Hamiltonian-connectedness in semicomplete digraphs. Journal of Algorithms, 13(1):114–127, March 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290008Z.

Bang-Jensen:2001:MSS

[BJY01] Jørgen Bang-Jensen and Anders Yeo. The minimum spanning strong subdigraph problem for extended semicomplete digraphs and semicomplete bipartite digraphs. Journal of Algorithms, 41(1):1–19, October 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911745.

Boyar:1987:CPG

[BK87] Joan F. Boyar and Howard J. Karloff. Coloring planar graphs in parallel. Journal of Algorithms, 8(4):470–479, December 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900460.

Bodlaender:1993:SLT

[BK93] H. Bodlaender and T. Kloks. A simple linear time algorithm for triangulating three-colored graphs. Journal of Algorithms, 15(1):

160–172, July 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710357.

Bodlaender:1996:ECA

[BK96] Hans L. Bodlaender and Ton Kloks. Efficient and constructive algorithms for the pathwidth and treewidth of graphs. Journal of Algorithms, 21(2):358–402, September 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900498.

Bhattacharya:1997:AMC

[BK97] Binay K. Bhattacharya and Damon Kaller. An O(m+n log n) algorithm for the maximum-clique problem in circular-arc graphs. Journal of Algorithms, 25(2):336–358, November 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908683.

Benczur:2000:AUE

[BK00] András A. Benczúr and David R. Karger. Augmenting undirected edge connectivity in Õ(n^{2}) time. Journal of Algorithms, 37(1): 2–36, October 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400910939.

Bodlaender:2002:KN

[BK02] Hans L. Bodlaender and Dieter Kratsch. Kayles and nimbers. Journal of Algorithms, 43(1):106–119, April 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402912150.

Barkan:2006:PAT

[BK06] Arye Barkan and Haim Kaplan. Partial alphabetic trees. Journal of Algorithms, 58(2):81–103, February 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000672.

Badger:2004:SBE

[BKL^{+}04] Jonathan Badger, Paul Kearney, Ming Li, John Tsang, and Tao Jiang. Selecting the branches for an evolutionary tree.: a polynomial time approximation scheme. Journal of Algorithms, 51(1): 1–14, April 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000865.

Bhatia:2000:LTS

[BKN00] Randeep Bhatia, Samir Khuller, and Joseph (Seffi) Naor. The loading time scheduling problem. Journal of Algorithms, 36(1): 1–33, July 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400910769.

Boyar:2004:SRA

[BKN04] Joan Boyar, Susan Krarup, and Morten N. Nielsen. Seat reservation allowing seat changes. Journal of Algorithms, 52(2):169– 192, August 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000410.

Barilan:1993:HAN

[BKP93] J. Barilan, G. Kortsarz, and D. Peleg. How to allocate network centers. Journal of Algorithms, 15(3):385–415, November 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710473.

Buss:1996:PAP

[BKRS96] Jonathan F. Buss, Paris C. Kanellakis, Prabhakar L. Ragde, and Alex Allister Shvartsman. Parallel algorithms with processor failures and delays. Journal of Algorithms, 20(1):45–86, January 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900036.

Barbeau:1997:EAC

[BKSD97] M. Barbeau, F. Kabanza, and R. St-Denis. An efficient algorithm for controller synthesis under full observation. Journal of Algorithms, 25(1):144–161, October 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908774.

Barve:2000:CPD

[BKVV00] Rakesh Barve, Mahesh Kallahalla, Peter J. Varman, and Jeffrey Scott Vitter. Competitive parallel disk prefetching and buffer management. Journal of Algorithms, 36(2):152–181, August 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400910897.

Birk:1992:FNI

[BL92] Yitzhak Birk and Jeffrey B. Lotspiech. On finding nonintersecting straightline connections of grid points to the boundary. Journal of Algorithms, 13(4):636–656, December 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290060P.

Blaha:1992:MBP

[Bla92] Kenneth D. Blaha. Minimum bases for permutation groups: the greedy approximation. Journal of Algorithms, 13(2):297– 306, June 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290020D.

Blaser:2004:AAA

[Blä04] Markus Bläser. An 8/13-approximation algorithm for the asymmetric maximum TSP. Journal of Algorithms, 50(1):23–48, January 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403001123.

Blumer:1987:HMD

[Blu87] Janet A. Blumer. How much is that DAWG in the window? A moving window algorithm for the directed acyclic word graph. Journal of Algorithms, 8(4):451–469, December 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900459.

Blum:2000:SDP

[Blu00] Norbert Blum. Speeding up dynamic programming without omitting any optimal solution and some applications in molecular biology. Journal of Algorithms, 35(2):129–168, May 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910782.

Bern:1987:LTC

[BLW87] M. W. Bern, E. L. Lawler, and A. L. Wong. Linear-time computation of optimal subgraphs of decomposable graphs. Journal of Algorithms, 8(2):216–235, June 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900393.

Broder:1997:CMW

[BM97] Andrei Z. Broder and Ernst W. Mayr. Counting minimum weight spanning trees. Journal of Algorithms, 24(1):171–176, July 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908512.

Bartal:2005:RSA

[BM05] Yair Bartal and Manor Mendel. Randomized k-server algorithms for growth-rate bounded graphs. Journal of Algorithms, 55(2): 192–202, May 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000914.

Berkman:1998:TLP

[BMR98] Omer Berkman, Yossi Matias, and Prabhakar Ragde. Triplylogarithmic parallel upper and lower bounds for minimum and range minima over small domains. Journal of Algorithms, 28(2):197–215, August 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497909056.

Blake:2006:RMA

[BMX06] Ian F. Blake, V. Kumar Murty, and Guangwu Xu. Refinements of Miller’s algorithm for computing the Weil/Tate pairing. Journal of Algorithms, 58(2):134–149, February 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000271.

Biran:1990:CCD

[BMZ90] Ofer Biran, Shlomo Moran, and Shmuel Zaks. A combinatorial characterization of the distributed 1-solvable tasks. Journal of Algorithms, 11(3):420–440, September 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090020F.

Bahig:2002:SPN

[BN02] Hatem M. Bahig and Ken Nakamula. Some properties of nonstar steps in addition chains and new cases where the Scholz Conjecture is true. Journal of Algorithms, 42(2):304–316, February 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402912125. See erratum [BN03].

Bahig:2003:ESP

[BN03] Hatem M. Bahig and Ken Nakamula. Erratum to “Some properties of nonstar steps in addition chains and new cases where the Scholz conjecture is true”: [J. Algorithms 42 (2002) 304– 316]. Journal of Algorithms, 47(1):60–61, April 2003. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000166. See [BN02].

Bar-Noy:2000:SMG

[BNHK^{+}00] Amotz Bar-Noy, Magnús M. Halldórsson, Guy Kortsarz, Ravit Salman, and Hadas Shachnai. Sum multicoloring of graphs. Journal of Algorithms, 37(2):422–450, November 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911064.

Bar-Noy:1998:MCS

[BNK98] Amotz Bar-Noy and Guy Kortsarz. Minimum color sum of bipartite graphs. Journal of Algorithms, 28(2):339–365, August 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909385.

Bar-Noy:2003:CLS

[BNL03] Amotz Bar-Noy and Richard E. Ladner. Competitive online stream merging algorithms for media-on-demand. Journal of Algorithms, 48(1):59–90, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000452.

Bar-Noy:2004:EWC

[BNM04] Amotz Bar-Noy and Grzegorz Malewicz. Establishing wireless conference calls under delay constraints. Journal of Algorithms, 51(2):145–169, May 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001664.

Bodlaender:1990:PAG

[Bod90] Hans L. Bodlaender. Polynomial algorithms for graph isomorphism and chromatic index on partial k-trees. Journal of Algorithms, 11(4):631–643, December 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900135.

Bodlaender:1993:LTM

[Bod93] H. L. Bodlaender. On linear time minor tests with depth-first search. Journal of Algorithms, 14(1):1–23, January 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710011.

Borwein:1985:CCF

[Bor85] Peter B. Borwein. On the complexity of calculating factorials. Journal of Algorithms, 6(3):376–380, September 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900069.

Boyer:2005:SCA

[Boy05] John M. Boyer. Simple constant amortized time generation of fixed length numeric partitions. Journal of Algorithms, 54(1):31– 39, January 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000197.

Becker:1983:SAT

[BP83] Ronald I. Becker and Yehoshua Perl. Shifting algorithms for tree partitioning with general weighting functions. Journal of Algorithms, 4(2):101–120, June 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677483900391.

Babai:2004:SBG

[BP04] László Babai and Igor Pak. Strong bias of group generators: an obstacle to the “product replacement algorithm”. Journal of Algorithms, 50(2):215–231, February 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000919.

Boutalis:2000:NMR

[BPK00] Y. Boutalis, C. Papaodysseus, and E. Koukoutsis. A new multichannel recursive least squares algorithm for very robust and efficient adaptive filtering. Journal of Algorithms, 37(2):283–308, November 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911131.

Babel:1996:IPD

[BPT96] L. Babel, I. N. Ponomarenko, and G. Tinhofer. The isomorphism problem for directed path graphs and for rooted directed path graphs. Journal of Algorithms, 21(3):542–564, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900589.

Bruck:1991:HPB

[BR91] Jehoshua Bruck and Vwani P. Roychowdhury. How to play bowling in parallel on the grid. Journal of Algorithms, 12(3):516–529, September 1991. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749190017S.

Berman:1994:IAS

[BR94] P. Berman and V. Ramaiyer. Improved approximations for the Steiner tree problem. Journal of Algorithms, 17(3):381–408, November 1994. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677484710418.

Bartal:1997:DSP

[BR97] Yair Bartal and Adi Rosén. The distributed k-server problem — a competitive distributed translator for k-server algorithms. Journal of Algorithms, 23(2):241–264, May 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908263.

Braga:2007:SIL

[Bra07] Christiano Braga. Special issue: LSFA’06. Journal of Algorithms, 62(3–4):93–94, July/October 2007. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677407000284.

Breslauer:1995:DMU

[Bre95] D. Breslauer. Dictionary-matching on unbounded alphabets: Uniform length dictionaries. Journal of Algorithms, 18(2):278– 295, March 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710115.

Bernstein:1989:AAS

[BRG89] David Bernstein, Michael Rodeh, and Izidor Gertner. Approximation algorithms for scheduling arithmetic expressions on pipelined machines. Journal of Algorithms, 10(1):120–139, March 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677489900278.

Brinck:1986:DTB

[Bri86] Keith Brinck. On deletion in threaded binary trees. Journal of Algorithms, 7(3):395–411, September 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900301.

Bridgland:1987:UTS

[Bri87] Michael F. Bridgland. Universal traversal sequences for paths and cycles. Journal of Algorithms, 8(3):395–404, September 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900198.

Balasubramanian:1997:FST

[BRS97] R. Balasubramanian, Venkatesh Raman, and G. Srinivasaragavan. Finding scores in tournaments. Journal of Algorithms, 24(2):380–394, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677497908658.

Brun:2008:SST

[Bru08] Yuriy Brun. Solving satisfiability in the tile assembly model with a constant-size tileset. Journal of Algorithms, 63(4):151– 166, October 2008. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677408000199.

Bentley:1980:DSP

[BS80] Jon Louis Bentley and James B. Saxe. Decomposable searching problems I. Static-to-dynamic transformation. Journal of Algorithms, 1(4):301–358, December 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677480900152.

Bagchi:1985:WHS

[BS85a] A. Bagchi and P. K. Srimani. Weighted heuristic search in networks. Journal of Algorithms, 6(4):550–576, December 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748590032X.

Bollobas:1985:RRI

[BS85b] Béla Bollobás and Istvan Simon. Repeated random insertion into a priority queue. Journal of Algorithms, 6(4):466–477, December 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677485900288.

Bergen:1991:PAC

[BS91a] James R. Bergen and Haim Shvaytser (Schweitzer). A probabilistic algorithm for computing Hough transforms. Journal of Algorithms, 12(4):639–656, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190037Y.

Bienstock:1991:MGS

[BS91b] D. Bienstock and Paul Seymour. Monotonicity in graph searching. Journal of Algorithms, 12(2):239–245, June 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190003H.

Benasher:1995:CDR

[BS95a] Y. Benasher and A. Schuster. The complexity of data reduction on a reconfigurable linear array. Journal of Algorithms, 18(2):322– 357, March 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710139.

Blum:1995:CRS

[BS95b] A. Blum and J. Spencer. Coloring random and semi-random kcolorable graphs. Journal of Algorithms, 19(2):204–234, September 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710346.

Berger:1997:TBM

[BS97a] Bonnie Berger and Peter W. Shor. Tight bounds for the maximum acyclic subgraph problem. Journal of Algorithms, 25(1): 1–18, October 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908646.

Bhattacharya:1997:SPO

[BS97b] Binay K. Bhattacharya and Sandeep Sen. On a simple, practical, optimal, output-sensitive randomized planar convex hull algorithm. Journal of Algorithms, 25(1):177–193, October 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908695.

Bryant:2001:COT

[BS01] David Bryant and Mike Steel. Constructing optimal trees from quartets. Journal of Algorithms, 38(1):237–259, January 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911337.

Bshouty:1999:LBC

[Bsh99] Nader H. Bshouty. Lower bounds for the complexity of functions in a realistic RAM model. Journal of Algorithms, 32(1): 1–20, July 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909944.

Bazgan:1999:AFN

[BST99] Cristina Bazgan, Miklos Santha, and Zsolt Tuza. On the approximation of finding a(nother) Hamiltonian cycle in cubic Hamiltonian graphs. Journal of Algorithms, 31(1):249–268, April 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909981.

Berkman:1993:ODL

[BSV93] O. Berkman, B. Schieber, and U. Vishkin. Optimal doubly logarithmic parallel algorithms based on finding all nearest smaller values. Journal of Algorithms, 14(3):344–370, May 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710187.

Bhattacharya:1983:EAC

[BT83] Binay K. Bhattacharya and Godfried T. Toussaint. Efficient algorithms for computing the maximum distance between two finite planar sets. Journal of Algorithms, 4(2):121–136, June 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900408.

Booth:1993:FMC

[BT93] H. Booth and R. E. Tarjan. Finding the minimum-cost maximum flow in a series-parallel network. Journal of Algorithms, 15(3):416–446, November 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677483710485.

Bose:1995:GTB

[BT95a] P. Bose and G. Toussaint. Growing a tree from its branches. Journal of Algorithms, 19(1):86–103, July 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710280.

Buchsbaum:1995:CPD

[BT95b] A. L. Buchsbaum and R. E. Tarjan. Confluently persistent deques via data-structural bootstrapping. Journal of Algorithms, 18(3):

513–547, May 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710206.

Bodlaender:1999:GBM

[BT99] Hans L. Bodlaender and Dimitrios M. Thilikos. Graphs with branchwidth at most three. Journal of Algorithms, 32(2):167– 194, August 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910118.

Burge:1981:STL

[Bur81] William H. Burge. Stacks in a two-level store. Journal of Algorithms, 2(2):178–185, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900183.

Burley:1996:TLG

[Bur96] William R. Burley. Traversing layered graphs using the work function algorithm. Journal of Algorithms, 20(3):479–511, May 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900243.

Butler:1983:CNP

[But83] Gregory Butler. Computing normalizers in permutation groups. Journal of Algorithms, 4(2):163–175, June 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900433.

Bonizzoni:1999:AMD

[BV99] Paola Bonizzoni and Gianluca Della Vedova. An algorithm for the modular decomposition of hypergraphs. Journal of Algorithms, 32(2):65–86, August 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910167.

Bender:1985:TAB

[BW85] Edward A. Bender and Herbert S. Wilf. A theoretical analysis of backtracking in the graph coloring problem. Journal of Algorithms, 6(2):275–282, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900446.

Bar-Yehuda:2001:UHW

[BY01] Reuven Bar-Yehuda. Using homogeneous weights for approximating the partial cover problem. Journal of Algorithms, 39(2): 137–144, May 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911507.

Bar-Yehuda:1981:LTA

[BYE81] R. Bar-Yehuda and S. Even. A linear-time approximation algorithm for the weighted vertex cover problem. Journal of Algorithms, 2(2):198–203, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900201.

Bar-Yehuda:2005:AGP

[BYES05] Reuven Bar-Yehuda, Guy Even, and Shimon (Moni) Shahar. On approximating a geometric prize-collecting traveling salesman problem with time windows. Journal of Algorithms, 55(1): 76–92, April 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001688.

Bar-Yehuda:1988:FTD

[BYK88] Reuven Bar-Yehuda and Shay Kutten. Fault tolerant distributed majority commitment. Journal of Algorithms, 9(4):568–582, December 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748890017X.

Bar-Yehuda:2002:AEW

[BYR02] Reuven Bar-Yehuda and Dror Rawitz. Approximating elementweighted vertex deletion problems for the complete k-partite property. Journal of Algorithms, 42(1):20–40, January 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911964.

Bhattacharya:1988:STD

[BZ88] B. K. Bhattacharya and J. Zorbas. Solving the two-dimensional findpath problem using a line-triangle representation of the robot.

Journal of Algorithms, 9(4):449–469, December 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900120.

Bitan:1993:OLB

[BZ93] S. Bitan and S. Zaks. Optimal linear broadcast. Journal of Algorithms, 14(2):288–315, March 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710151.

Cabello:2007:AAS

[Cab07] Sergio Cabello. Approximation algorithms for spreading points. Journal of Algorithms, 62(2):49–73, April 2007. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001129.

Cantor:1981:IPI

[Can81] David G. Cantor. Irreducible polynomials with integral coefficients have succinct certificates. Journal of Algorithms, 2(4):385– 392, December 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900365.

Cary:2001:TOA

[Car01] Matthew Cary. Toward optimal ϵ-approximate nearest neighbor algorithms. Journal of Algorithms, 41(2):417–428, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677401911939.

Chekuri:2001:EAA

[CB01] Chandra Chekuri and Michael Bender. An efficient approximation algorithm for minimizing makespan on uniformly related machines. Journal of Algorithms, 41(2):212–224, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677401911848.

Charron-Bost:2004:UCH

[CBS04] Bernadette Charron-Bost and André Schiper. Uniform consensus is harder than consensus. Journal of Algorithms, 51(1):15– 37, April 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001652.

Chan:2000:EAF

[CC00] Wun-Tat Chan and Francis Y. L. Chin. Efficient algorithms for finding the maximum number of disjoint paths in grids. Journal of Algorithms, 34(2):337–369, February 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910544.

Caruso:2002:DRS

[CCMS02] Antonio Caruso, Stefano Chessa, Piero Maestrini, and Paolo Santi. Diagnosability of regular systems. Journal of Algorithms, 45(2):126–143, November 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740200250X.

Charikar:1999:AAD

[CCyC^{+}99] Moses Charikar, Chandra Chekuri, To yat Cheung, Zuo Dai, Ashish Goel, Sudipto Guha, and Ming Li. Approximation algorithms for directed Steiner problems. Journal of Algorithms, 33(1):73–91, October 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910428.

Caspi:1995:ECS

[CD95] Y. Caspi and E. Dekel. Edge coloring series parallel graphs. Journal of Algorithms, 18(2):296–321, March 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710127.

Cortes:2009:BST

[CDBPL^{+}09] C. Cortés, J. M. Díaz-Báñez, P. Pérez-Lantero, C. Seara, J. Urrutia, and I. Ventura. Bichromatic separability with two boxes: a general approach. Journal of Algorithms, 64(2–3):79–88, April/ July 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677409000029.

Cooper:2001:MCR

[CDF01] Colin Cooper, Martin Dyer, and Alan Frieze. On Markov chains for randomly H-coloring a graph. Journal of Algorithms, 39(1): 117–134, April 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911428.

Chen:2003:FOP

[CDHX03] Danny Z. Chen, Ovidiu Daescu, Xiaobo (Sharon) Hu, and Jinhui Xu. Finding an optimal path without growing the tree. Journal of Algorithms, 49(1):13–41, October 2003. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000804.

Colbourn:1989:URS

[CDN89] Charles J. Colbourn, Robert P. J. Day, and Louis D. Nel. Unranking and ranking spanning trees of a graph. Journal of Algorithms, 10(2):271–286, June 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900163.

Chepoi:2006:DRL

[CDV06] Victor Chepoi, Feodor F. Dragan, and Yann Vaxès. Distance and routing labeling schemes for non-positively curved plane graphs. Journal of Algorithms, 61(2):60–88, November 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001415.

Chan:2001:FCM

[CE01] Timothy M. Chan and Alon Efrat. Fly cheaply: On the minimum fuel consumption problem. Journal of Algorithms, 41(2):330–337, November 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911897.

Cohen:1996:WBT

[CF96] David Cohen and Michael L. Fredman. Weighted binary trees for concurrent searching. Journal of Algorithms, 20(1):87–112, January 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900048.

Costantini:2009:MPC

[CF09] Stefania Costantini and Andrea Formisano. Modeling preferences and conditional preferences on resource consumption and production in ASP. Journal of Algorithms, 64(1):3–15, January 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677409000042.

Calinescu:1998:BAA

[CFFK98] Gruia Calinescu, Cristina G. Fernandes, Ulrich Finkler, and Howard Karloff. A better approximation algorithm for finding planar subgraphs. Journal of Algorithms, 27(2):269–302, May 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497909202.

Csirik:1991:PAA

[CFGK91] J. Csirik, J. B. G. Frenk, G. Galambos, and A. H. G. Rinnooy Kan. Probabilistic analysis of algorithms for dual bin packing problems. Journal of Algorithms, 12(2):189–203, June 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190001F.

Calinescu:2003:MUG

[CFR03] Gruia Calinescu, Cristina G. Fernandes, and Bruce Reed. Multicuts in unweighted graphs and digraphs with bounded degree and bounded tree-width. Journal of Algorithms, 48(2):333–359, September 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000737.

Corneil:1984:NFA

[CG84] Derek Corneil and Mark Goldberg. A non-factorial algorithm for canonical numbering of a graph. Journal of Algorithms, 5(3):345– 362, September 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900154.

Chou:1989:LBR

[CG89] Ching-Tsun Chou and Inder S. Gopal. Linear broadcast routing. Journal of Algorithms, 10(4):490–517, December 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900023.

Chakrabarti:1992:GBF

[CG92] P. P. Chakrabarti and S. Ghose. A general best first search algorithm in AND/OR graphs. Journal of Algorithms, 13(2):177– 187, June 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677492900144.

Camerini:1992:RPP

[CGM92] P. M. Camerini, G. Galbiati, and F. Maffioli. Random pseudo-polynomial algorithms for exact matroid problems. Journal of Algorithms, 13(2):258–273, June 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900188.

Chandru:1995:AML

[CGM^{+}95] V. Chandru, S. K. Ghosh, A. Maheshwari, V. T. Rajan, and S. J. Saluja. NC-algorithms for minimum link path and related problems. Journal of Algorithms, 19(2):173–203, September 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710334.

Czumaj:1997:SPA

[CGPR97] Artur Czumaj, Leszek Gasieniec, Marek Piotrów, and Wojciech Rytter. Sequential and parallel approximation of shortest superstrings. Journal of Algorithms, 23(1):74–100, April 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908238.

Chrobak:2002:FBG

[CGR02] Marek Chrobak, Leszek Gasieniec, and Wojciech Rytter. Fast broadcasting and gossiping in radio networks. Journal of Algorithms, 43(2):177–189, May 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402000044.

Cheng:1988:DWD

[CH88] Ying Cheng and Frank K. Hwang. Diameters of weighted double loop networks. Journal of Algorithms, 9(3):401–410, September 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900302.

Chandra:2001:AAD

[CH01a] Barun Chandra and Magnús M. Halldórsson. Approximation algorithms for dispersion problems. Journal of Algorithms, 38(2):438–465, February 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911453.

Chandra:2001:GLI

[CH01b] Barun Chandra and Magnús M. Halldórsson. Greedy local improvement and weighted set packing approximation. Journal of Algorithms, 39(2):223–240, May 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911556.

Chen:2003:ADT

[CH03] Wei-Mei Chen and Hsien-Kuei Hwang. Analysis in distribution of two randomized algorithms for finding the maximum in a broadcast communication model. Journal of Algorithms, 46(2):140– 177, February 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002936.

Chan:1998:DAD

[Cha98] Timothy M. Chan. Deterministic algorithms for 2-d convex programming and 3-d online linear programming. Journal of Algorithms, 27(1):147–166, April 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909147.

Chan:2003:PTA

[Cha03] Timothy M. Chan. Polynomial-time approximation schemes for packing and piercing fat objects. Journal of Algorithms, 46(2):178–189, February 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002948.

Chen:1999:CDQ

[CHC99] Wei-Mei Chen, Hsien-Kuei Hwang, and Gen-Huey Chen. The cost distribution of queue-Mergesort, optimal Mergesorts, and power-of-2 rules. Journal of Algorithms, 30(2):423–448, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909865.

Chen:1996:OCS

[Che96] Danny Z. Chen. Optimally computing the shortest weakly visible subedge of a simple polygon. Journal of Algorithms, 20(3): 459–478, May 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900231.

Chen:1998:QS

[Che98a] Jing-Chao Chen. Quadripartite sort. Journal of Algorithms, 28(2):258–271, August 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749890935X.

Chen:1998:MEA

[Che98b] Weimin Chen. More efficient algorithm for ordered tree inclusion. Journal of Algorithms, 26(2):370–385, February 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908993.

Chen:1998:EAS

[Che98c] Zhi-Zhong Chen. Efficient approximation schemes for maximization problems on K_{3,3}-free or K_{5}-free graphs. Journal of Algorithms, 26(1):166–187, January 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908944.

Chen:2001:NAO

[Che01a] Weimin Chen. New algorithm for ordered tree-to-tree correction problem. Journal of Algorithms, 40(2):135–158, August 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677401911708.

Chen:2001:AAI

[Che01b] Zhi-Zhong Chen. Approximation algorithms for independent sets in map graphs. Journal of Algorithms, 41(1):20–40, October 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677401911782.

Chang:1989:CCC

[CHS89] Yukon Chang, Susanne Hambrusch, and Janos Simon. On the computational complexity of continuous routing. Journal of Algorithms, 10(1):86–108, March 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900254.

Chern:2002:ATC

[CHT02] Hua-Huai Chern, Hsien-Kuei Hwang, and Tsung-Hsi Tsai. An asymptotic theory for Cauchy–Euler differential equations with applications to the analysis of algorithms. Journal of Algorithms, 44(1):177–225, July 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002080.

Chung:1987:TAS

[Chu87] Moon Jung Chung. O(n^{2.5}) time algorithms for the subgraph homeomorphism problem on trees. Journal of Algorithms, 8(1):106–112, March 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677487900307.

Chu:1992:OAN

[Chu92] Jeff Chu. Optimal algorithm for the nearest common dominator problem. Journal of Algorithms, 13(4):693–697, December 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290063I.

Caprara:2002:WRO

[CIM^{+}02] Alberto Caprara, Giuseppe F. Italiano, G. Mohan, Alessandro Panconesi, and Aravind Srinivasan. Wavelength rerouting in optical networks, or the Venetian Routing Problem. Journal of Algorithms, 45(2):93–125, November 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002146.

Cheng:1991:EMU

[CJ91] Siu Wing Cheng and Ravi Janardan. Efficient maintenance of the union of intervals on a line, with applications. Journal of Algorithms, 12(1):57–74, March 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190023R.

Cheng:1992:ARS

[CJ92] Siu Wing Cheng and Ravi Janardan. Algorithms for ray-shooting and intersection searching. Journal of Algorithms, 13(4):670–692, December 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290062H.

Coffman:2002:DSM

[CJM02] E. G. Coffman, Jr., Predrag Jelenković, and Petar Momcilović. The dyadic stream merging algorithm. Journal of Algorithms, 43(1):120–137, April 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402912204.

Chen:2001:VCF

[CKJ01] Jianer Chen, Iyad A. Kanj, and Weijia Jia. Vertex cover: Further observations and further improvements. Journal of Algorithms, 41(2):280–301, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911861.

Chew:1998:VDL

[CKS^{+}98] L. Paul Chew, Klara Kedem, Micha Sharir, Boaz Tagansky, and Emo Welzl. Voronoi diagrams of lines in 3-space under polyhedral convex distance functions. Journal of Algorithms, 29(2):238–255, November 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909579.

Chrobak:1991:FAT

[CL91] Marek Chrobak and Lawrence L. Larmore. On fast algorithms for two servers. Journal of Algorithms, 12(4):607–614, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190035W.

Chrobak:1994:GHC

[CL94] M. Chrobak and L. L. Larmore. Generosity helps or an 11-competitive algorithm for three servers. Journal of Algorithms, 16(2):234–263, March 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667748471011X.

Chong:1995:FCC

[CL95] K. W. Chong and T. W. Lam. Finding connected components in O(log n log log n) time on the EREW PRAM. Journal of Algorithms, 18(3):378–402, May 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710164.

Cohen:1999:AMM

[CL99] Edith Cohen and David D. Lewis. Approximating matrix multiplication for pattern recognition tasks. Journal of Algorithms, 30(2):211–252, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909890.

Chassaing:2002:RBB

[CL02] Philippe Chassaing and Guy Louchard. Reflected Brownian bridge area conditioned on its local time at the origin. Journal of Algorithms, 44(1):29–51, July 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002092.

Clarkson:1998:SP

[Cla98] Kenneth L. Clarkson. SODA ’95 papers. Journal of Algorithms, 29(2):203, November 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909762.

Chrobak:1997:PMA

[CLRW97] Marek Chrobak, Lawrence L. Larmore, Nick Reingold, and Jeffery Westbrook. Page migration algorithms using work functions. Journal of Algorithms, 24(1):124–157, July 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908536.

Cook:1987:PCD

[CM87] Stephen A. Cook and Pierre McKenzie. Problems complete for deterministic logarithmic space. Journal of Algorithms, 8(3):385– 394, September 1987. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677487900186.

Cheriyan:1988:FNI

[CM88] J. Cheriyan and S. N. Maheshwari. Finding nonseparating induced cycles and independent spanning trees in 3connected graphs. Journal of Algorithms, 9(4):507–537, December 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900156.

Chazelle:1996:LTD

[CM96] Bernard Chazelle and Jirí Matousek. On linear-time deterministic algorithms for optimization problems in fixed dimension. Journal of Algorithms, 21(3):579–597, November 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900607.

Cormode:2005:IDS

[CM05] Graham Cormode and S. Muthukrishnan. An improved data stream summary: the count-min sketch and its applications. Journal of Algorithms, 55(1):58–75, April 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001913.

Colbourn:1996:TAU

[CMN96] Charles J. Colbourn, Wendy J. Myrvold, and Eugene Neufeld. Two algorithms for unranking arborescences. Journal of Algorithms, 20(2):268–281, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900140.

Chiba:1989:HCP

[CN89] Norishige Chiba and Takao Nishizeki. The Hamiltonian cycle problem is linear-time solvable for 4-connected planar graphs. Journal of Algorithms, 10(2):187–211, June 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900126.

Chrobak:1990:IEC

[CN90] Marek Chrobak and Takao Nishizeki. Improved edge-coloring algorithms for planar graphs. Journal of Algorithms, 11(1):102– 116, March 1990. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749090032A.

Chrobak:2000:CAR

[CN00] Marek Chrobak and John Noga. Competitive algorithms for relaxed list update and multilevel caching. Journal of Algorithms, 34(2):282–308, February 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910611.

Chaudhuri:2003:LER

[CND03] Jeet Chaudhuri, Subhas C. Nandy, and Sandip Das. Largest empty rectangle among a point set. Journal of Algorithms, 46(1): 54–78, January 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002857.

Chiba:1981:LCA

[CNS81] Norishige Chiba, Takao Nishizeki, and Nobuji Saito. A linear 5-coloring algorithm of planar graphs. Journal of Algorithms, 2(4):317–327, December 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900316.

Coja-Oghlan:2007:SNH

[CO07] Amin Coja-Oghlan. Solving NP-hard semirandom graph problems in polynomial expected time. Journal of Algorithms, 62(1): 19–46, January 2007. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667740400118X.

Cohen:1996:EPS

[Coh96] Edith Cohen. Efficient parallel shortest-paths in digraphs with a separator decomposition. Journal of Algorithms, 21(2):331–357, September 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900486.

Cohen:1997:USP

[Coh97] Edith Cohen. Using selective path-doubling for parallel shortestpath computations. Journal of Algorithms, 22(1):30–56, January 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908135.

Coja-Oghlan:2006:HSC

[COKN06] Amin Coja-Oghlan, Sven O. Krumke, and Till Nierhoff. A heuristic for the Stacker Crane Problem on trees which is almost surely exact. Journal of Algorithms, 61(1):1–19, September 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001208.

Cole:1986:SSS

[Col86] Richard Cole. Searching and storing similar lists. Journal of Algorithms, 7(2):202–220, June 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900040.

Colussi:1994:FPM

[Col94] L. Colussi. Fastest pattern matching in strings. Journal of Algorithms, 16(2):163–189, March 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667748471008X.

Cowen:2001:CRM

[Cow01] Lenore J. Cowen. Compact routing with minimum stretch. Journal of Algorithms, 38(1):170–183, January 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911349.

Caprara:2003:PCU

[CPR03] Alberto Caprara, Alessandro Panconesi, and Romeo Rizzi. Packing cycles in undirected graphs. Journal of Algorithms, 48(1):239– 256, August 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667740300052X.

Calimeri:2008:EPI

[CPR08] F. Calimeri, S. Perri, and F. Ricca. Experimenting with parallelism for the instantiation of ASP programs. Journal of Algorithms, 63(1–3):34–54, January/July 2008. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677408000059.

Culberson:1994:CPH

[CR94] J. C. Culberson and R. A. Reckhow. Covering polygons is hard. Journal of Algorithms, 17(1):2–44, July 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667748471025X.

Czumaj:2006:BAR

[CR06] Artur Czumaj and Wojciech Rytter. Broadcasting algorithms in radio networks with unknown topology. Journal of Algorithms, 60(2):115–143, August 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001439.

Cattell:2000:FAG

[CRS^{+}00] Kevin Cattell, Frank Ruskey, Joe Sawada, Micaela Serra, and C. Robert Miers. Fast algorithms to generate necklaces, unlabeled necklaces, and irreducible polynomials over GF(2). Journal of Algorithms, 37(2):267–282, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0196677400911088. See addendum [SR03].

Capoyleas:1991:GC

[CRW91] Vasilis Capoyleas, Günter Rote, and Gerhard Woeginger. Geometric clusterings. Journal of Algorithms, 12(2):341–356, June 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190007L.

Cypher:1992:CPA

[CS92] Robert Cypher and Jorge L. C. Sanz. Cubesort: a parallel algorithm for sorting N data items with S-sorters. Journal of Algorithms, 13(2):211–234, June 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900166.

Chudak:1999:AAP

[CS99] Fabián A. Chudak and David B. Shmoys. Approximation algorithms for precedence-constrained scheduling problems on parallel machines that run at different speeds. Journal of Algorithms, 30(2):323–343, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909877.

Cohen:2006:MTD

[CS06] Edith Cohen and Martin J. Strauss. Maintaining time-decaying stream aggregates. Journal of Algorithms, 59(1):19–36, April 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677405000234.

Csirik:1993:PBF

[Csi93] J. Csirik. The parametric behavior of the first-fit decreasing bin packing algorithm. Journal of Algorithms, 15(1):1–28, July 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667748371028X.

Cypher:1989:HSE

[CSS89] R. Cypher, J. L. C. Sanz, and L. Snyder. Hypercube and shuffle-exchange algorithms for image component labeling. Journal of Algorithms, 10(1):140–150, March 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748990028X.

Caro:1996:RGS

[CST96] Yair Caro, András Seboő, and Michael Tarsi. Recognizing greedy structures. Journal of Algorithms, 20(1):137–156, January 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900061.

Cheriyan:1999:FAS

[CT99] Joseph Cheriyan and Ramakrishna Thurimella. Fast algorithms for k-shredders and k-node connectivity augmentation. Journal of Algorithms, 33(1):15–50, October 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910404.

Chen:2006:CBD

[CT06] Zhi-Zhong Chen and Tatsuie Tsukiji. Computing bounded-degree phylogenetic roots of disconnected graphs. Journal of Algorithms, 59(2):125–148, May 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000027.

Chandra:1995:CRC

[CV95] B. Chandra and S. Vishwanathan. Constructing reliable communication networks of small weight online. Journal of Algorithms, 18(1):159–175, January 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667748571005X.

Chaudhary:2001:AAA

[CV01] Amitabh Chaudhary and Sundar Vishwanathan. Approximation algorithms for the achromatic number. Journal of Algorithms, 41(2):404–416, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0196677401911927. See erratum [Ano02i].

Csirik:2002:RAO

[CW02] János Csirik and Gerhard J. Woeginger. Resource augmentation for online bounded space bin packing. Journal of Algorithms, 44(2):308–320, August 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740200202X.

Cowen:2004:CRR

[CW04] Lenore J. Cowen and Christopher G. Wagner. Compact roundtrip routing in directed networks. Journal of Algorithms, 50(1):79– 95, January 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001275.

Cole:1987:SP

[CY87] Richard Cole and Chee K. Yap. Shape from probing. Journal of Algorithms, 8(1):19–38, March 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900253.

Chrobak:1989:FAE

[CY89] Marek Chrobak and Moti Yung. Fast algorithms for edge-coloring planar graphs. Journal of Algorithms, 10(1):35–51, March 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677489900229.

Chen:1991:PRC

[CY91] Lin Chen and Yaacov Yesha. Parallel recognition of the consecutive ones property with applications. Journal of Algorithms, 12(3):375–392, September 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190010V.

Cole:1995:APA

[CZ95] R. Cole and O. Zajicek. An asynchronous parallel algorithm for undirected graph connectivity. Journal of Algorithms, 18(1):50– 97, January 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710024.

Cohen:2001:APS

[CZ01] Edith Cohen and Uri Zwick. All-pairs small-stretch paths. Journal of Algorithms, 38(2):335–353, February 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911179.

Czumaj:1996:VFA

[Czu96] Artur Czumaj. Very fast approximation of the matrix chain product problem. Journal of Algorithms, 21(1):71–79, July 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900371.

Domanski:1985:CDA

[DA85] B. Domanski and M. Anshel. The complexity of Dehn’s algorithm for word problems in groups. Journal of Algorithms, 6(4):543– 549, December 1985. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677485900318.

Dahlhaus:2000:PAH

[Dah00] Elias Dahlhaus. Parallel algorithms for hierarchical clustering and applications to split decomposition and parity graph recognition. Journal of Algorithms, 36(2):205–240, August 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910903.

Dubois:1997:GUB

[DB97] O. Dubois and Y. Boufkhad. A general upper bound for the satisfiability threshold of random r-SAT formulae. Journal of Algorithms, 24(2):395–420, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908671.

Danger:2009:GCO

[DB09] Roxana Danger and Rafael Berlanga. Generating complex ontology instances from documents. Journal of Algorithms, 64(1): 16–30, January 2009. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677409000054.

deBerg:1992:GAD

[dBCO92] Mark T. de Berg, Svante Carlsson, and Mark H. Overmars. A general approach to dominance in the plane. Journal of Algorithms, 13(2):274–296, June 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900199.

deBerg:2005:TNV

[dBGK^{+}05] Mark de Berg, Joachim Gudmundsson, Matthew J. Katz, Christos Levcopoulos, Mark H. Overmars, and A. Frank van der Stappen. TSP with neighborhoods of varying size. Journal of Algorithms, 57(1):22–36, September 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000246.

Doh:1993:ADV

[DC93] J. I. Doh and K. Y. Chwa. An algorithm for determining visibility of a simple polygon from an internal line segment. Journal of Algorithms, 14(1):139–168, January 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710072.

Dasgupta:1996:MHS

[DCD96] Pallab Dasgupta, P. P. Chakrabarti, and S. C. DeSarkar. Multiobjective heuristic search in AND/OR graphs. Journal of Algorithms, 20(2):282–311, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900152.

Diks:1993:ESP

[DDSV93] K. Diks, H. N. Djidjev, O. Sykora, and I. Vrto. Edge separators of planar and outerplanar graphs with applications. Journal of Algorithms, 14(2):258–279, March 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710138.

Dobkin:1987:SSI

[DE87] David P. Dobkin and Herbert Edelsbrunner. Space searching for intersecting objects. Journal of Algorithms, 8(3):348–361, September 1987. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677487900150.

Devroye:1985:ELL

[Dev85] Luc Devroye. The expected length of the longest probe sequence for bucket searching when the distribution is not uniform. Journal of Algorithms, 6(1):1–9, March 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748590015X.

Dyer:1986:PNC

[DF86] M. E. Dyer and A. M. Frieze. Planar 3DM is NP-complete. Journal of Algorithms, 7(2):174–184, June 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900027.

Dyer:1989:SSR

[DF89] M. E. Dyer and A. M. Frieze. The solution of some random NP-hard problems in polynomial expected time. Journal of Algorithms, 10(4):451–489, December 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900011.

Doignon:1994:PTA

[DF94] J. P. Doignon and J. C. Falmagne. A polynomial time algorithm for unidimensional unfolding representations. Journal of Algorithms, 16(2):218–233, March 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710108.

DellAmico:2002:LTA

[DF02] Mauro Dell’Amico and Lucian Finta. A linear time algorithm for scheduling outforests with communication delays on three processors. Journal of Algorithms, 44(2):287–307, August 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002018.

deFigueiredo:2000:FSP

[dFKKR00] Celina M. H. de Figueiredo, Sulamita Klein, Yoshiharu Kohayakawa, and Bruce A. Reed. Finding skew partitions efficiently. Journal of Algorithms, 37(2):505–521, November 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499911227.

Diks:2004:TEL

[DFKP04] Krzysztof Diks, Pierre Fraigniaud, Evangelos Kranakis, and Andrzej Pelc. Tree exploration with little memory. Journal of Algorithms, 51(1):38–63, April 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001615.

Dyer:2000:MCI

[DG00] Martin Dyer and Catherine Greenhill. On Markov chains for independent sets. Journal of Algorithms, 35(1):17–49, April 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910714.

Dor:2001:AGS

[DG01] Avner Dor and Eitan Greenshtein. An almost-greedy search on random binary vectors and random graphs. Journal of Algorithms, 40(1):102–133, July 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911404.

Duncan:2001:BAR

[DGK01] Christian A. Duncan, Michael T. Goodrich, and Stephen Kobourov. Balanced aspect ratio trees: Combining the advantages of k-d trees and octrees. Journal of Algorithms, 38(1):303– 333, January 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911350.

Dahlhaus:2001:EPA

[DGM01] Elias Dahlhaus, Jens Gustedt, and Ross M. McConnell. Efficient and practical algorithms for sequential modular decomposition. Journal of Algorithms, 41(2):360–387, November 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740191185X.

Duran:2006:PTR

[DGM^{+}06] Guillermo Durán, Agustín Gravano, Ross M. McConnell, Jeremy Spinrad, and Alan Tucker. Polynomial time recognition of unit circular-arc graphs. Journal of Algorithms, 58(1):67–78, January 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404001464.

Dahlhaus:1993:PCH

[DHK93] E. Dahlhaus, P. Hajnal, and M. Karpinski. On the parallel complexity of Hamiltonian cycle and matching problem on dense graphs. Journal of Algorithms, 15(3):367–384, November 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710461.

Dietzfelbinger:1997:RRA

[DHKP97] Martin Dietzfelbinger, Torben Hagerup, Jyrki Katajainen, and Martti Penttonen. A reliable randomized algorithm for the closest-pair problem. Journal of Algorithms, 25(1):19–51, October 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497908737.

Diaconis:1980:ART

[Dia80] Persi Diaconis. Average running time of the fast Fourier transform. Journal of Algorithms, 1(2):187–208, June 1980. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748090022X.

Damian-Iordache:2002:ASM

[DIP02] Mirela Damian-Iordache and Sriram V. Pemmaraju. A (2 + ϵ)- approximation scheme for minimum domination on circle graphs. Journal of Algorithms, 42(2):255–276, February 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401912064.

Dobkin:1985:LAD

[DK85] David P. Dobkin and David G. Kirkpatrick. A linear algorithm for determining the separation of convex polyhedra. Journal of Algorithms, 6(3):381–392, September 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900070.

Delatorre:1995:FPA

[DK95] P. Delatorre and C. P. Kruskal. Fast parallel algorithms for all-sources lexicographic search and path-algebra problems. Journal of Algorithms, 19(1):1–24, July 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710255.

Dolev:1999:BPR

[DKK99] Shlomi Dolev, Evangelos Kranakis, and Danny Krizanc. Bakedpotato routing. Journal of Algorithms, 30(2):379–399, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909816.

Dolev:1982:UDA

[DKR82] Danny Dolev, Maria Klawe, and Michael Rodeh. An O(n log n) unidirectional distributed algorithm for extrema finding in a circle. Journal of Algorithms, 3(3):245–260, September 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900232.

Dawande:2001:BMC

[DKST01] Milind Dawande, Pinar Keskinocak, Jayashankar M. Swaminathan, and Sridhar Tayur. On bipartite and multipartite clique problems. Journal of Algorithms, 41(2):388–403, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740191199X.

Du:1993:MMFa

[DL93a] J. Z. Du and J. Y. T. Leung. Minimizing mean flow time in twomachine open shops and flow shops. Journal of Algorithms, 14(1): 24–44, January 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710023.

Du:1993:MMFb

[DL93b] J. Z. Du and J. Y. T. Leung. Minimizing mean flow time with release time and deadline constraints. Journal of Algorithms, 14(1): 45–68, January 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710035.

daLuz:2007:RET

[dLFM07] Robson da Luz, Mírian Halfeld Ferrari, and Martin A. Musicante. Regular expression transformations to extend regular languages (with application to a Datalog XML schema validator). Journal of Algorithms, 62(3–4):148–167, July/October 2007. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677407000326.

DelLungo:2005:GSP

[DLMM05] Alberto Del Lungo, Guy Louchard, Claudio Marini, and Franco Montagna. The Guessing Secrets problem: a probabilistic approach. Journal of Algorithms, 55(2):142–176, May 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000422.

Demaine:2003:LLB

[DLO03] Erik D. Demaine and Alejandro López-Ortiz. A linear lower bound on index size for text retrieval. Journal of Algorithms, 48(1):2–15, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000439.

Datta:1995:SDA

[DLSS95] A. Datta, H. P. Lenhof, C. Schwarz, and M. Smid. Static and dynamic algorithms for k-point Clustering problems. Journal of Algorithms, 19(3):474–503, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710486.

delaTorre:1997:UAA

[dlTK97] Pilar de la Torre and David T. Kao. A uniform approach to the analysis of trie structures that store prefixing-keys. Journal of Algorithms, 22(2):270–295, February 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908226.

Deng:2003:AAP

[DLZZ03] Xiaotie Deng, Guojun Li, Wenan Zang, and Yi Zhou. A 2approximation algorithm for path coloring on a restricted class of trees of rings. Journal of Algorithms, 47(1):1–13, April 2003. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000038.

Dobkin:1985:EUP

[DM85] David P. Dobkin and J. Ian Munro. Efficient uses of the past. Journal of Algorithms, 6(4):455–465, December 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900276.

Duch:2002:APO

[DM02] Amalia Duch and Conrado Martínez. On the average performance of orthogonal range search in multidimensional data structures. Journal of Algorithms, 44(1):226–245, July 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002134.

Dumitrescu:2003:AAT

[DM03] Adrian Dumitrescu and Joseph S. B. Mitchell. Approximation algorithms for TSP with neighborhoods in the plane. Journal of Algorithms, 48(1):135–159, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000476.

Dinitz:1999:AAF

[DN99] Yefim Dinitz and Zeev Nutov. A 3-approximation algorithm for finding optimum 4, 5-vertex-connected spanning subgraphs. Journal of Algorithms, 32(1):31–40, July 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910076.

Dolev:1982:BGS

[Dol82] Danny Dolev. The Byzantine generals strike again. Journal of Algorithms, 3(1):14–30, March 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677482900049.

Dor:1998:GSA

[Dor98] Avner Dor. The greedy search algorithm on binary vectors. Journal of Algorithms, 27(1):42–60, April 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908932.

Diaz:2001:ALP

[DPPS01] Josep Díaz, Mathew D. Penrose, Jordi Petit, and María Serna. Approximating layout problems on random geometric graphs. Journal of Algorithms, 39(1):78–116, April 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911490.

Dietz:1999:SRS

[DR99] Paul F. Dietz and Rajeev Raman. Small-rank selection in parallel, with applications to heap construction. Journal of Algorithms, 30(1):33–51, January 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909713.

DiGaspero:2008:SLS

[DR08] Luca Di Gaspero and Andrea Roli. Stochastic local search for large-scale instances of the haplotype inference problem by pure parsimony. Journal of Algorithms, 63(1–3):55–69, January/July 2008. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677408000060.

Dragan:2005:EAP

[Dra05] Feodor F. Dragan. Estimating all pairs shortest paths in restricted graph families: a unified approach. Journal of Algorithms, 57(1):1–21, September 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001488.

Devroye:1993:CHR

[DT93] L. Devroye and G. Toussaint. Convex hulls for random lines. Journal of Algorithms, 14(3):381–394, May 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710205.

Duval:1983:FWO

[Duv83] Jean Pierre Duval. Factorizing words over an ordered alphabet. Journal of Algorithms, 4(4):363–381, December 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900172.

Deberg:1995:TDT

[DVS95] M. Deberg, M. Vankreveld, and J. Snoeyink. Two-dimensional and three-dimensional point location in rectangular subdivisions. Journal of Algorithms, 18(2):256–277, March 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710103.

Dixon:1983:RSU

[DW83] John D. Dixon and Herbert S. Wilf. The random selection of unlabeled graphs. Journal of Algorithms, 4(3):205–213, September 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677483900214.

Dolev:1984:SPG

[DW84] Danny Dolev and Manfred K. Warmuth. Scheduling precedence graphs of bounded height. Journal of Algorithms, 5(1): 48–59, March 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900397.

Dumitrescu:2004:AOM

[DW04] Sorina Dumitrescu and Xiaolin Wu. Algorithms for optimal multiresolution quantization. Journal of Algorithms, 50(1):1–22, January 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000993.

ElGindy:1981:LAC

[EA81] H. El Gindy and D. Avis. A linear algorithm for computing the visibility polygon from a point. Journal of Algorithms, 2(2):186– 197, June 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900195.

Eberly:1994:LDC

[Ebe94] W. Eberly. Logarithmic depth circuits for Hermite interpolation. Journal of Algorithms, 16(3):335–360, May 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710169.

Everett:1990:RVG

[EC90] H. Everett and D. G. Corneil. Recognizing visibility graphs of spiral polygons. Journal of Algorithms, 11(1):1–26, March 1990. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090026B.

Edelsbrunner:1985:CED

[Ede85] H. Edelsbrunner. Computing the extreme distances between two convex polygons. Journal of Algorithms, 6(2):213–224, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677485900392.

Epstein:2005:ONP

[EF05] Leah Epstein and Lene M. Favrholdt. Optimal non-preemptive semi-online scheduling on two related machines. Journal of Algorithms, 57(1):49–73, September 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001567.

Ellis:2004:DSP

[EFF04] J. Ellis, H. Fan, and M. Fellows. The dominating set problem is fixed parameter tractable for graphs of bounded genus. Journal of Algorithms, 52(2):152–168, August 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000409.

Even:1981:MLG

[EG81] S. Even and O. Goldreich. The minimum-length generator sequence problem is NP-hard. Journal of Algorithms, 2(3):311–313, September 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900298.

Ehrenfeucht:1994:DCA

[EGMS94] A. Ehrenfeucht, H. N. Gabow, R. M. Mcconnell, and S. J. Sullivan. An O(n^{2}) divide-and-conquer algorithm for the prime tree decomposition of two-structures and modular decomposition of graphs. Journal of Algorithms, 16(2):283–294, March 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710133.

Eilam:2003:CRS

[EGP03] Tamar Eilam, Cyril Gavoille, and David Peleg. Compact routing schemes with low stretch factor. Journal of Algorithms, 46(2):97– 114, February 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000026.

Eu:1996:EAL

[EGT96] D. Eu, E. Guévremont, and G. T. Toussaint. On envelopes of arrangements of lines. Journal of Algorithms, 21(1):111–148, July 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900401.

Eppstein:1997:CSM

[EH97] David Eppstein and Daniel S. Hirschberg. Choosing subsets with maximum weighted average. Journal of Algorithms, 24(1):177– 193, July 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908494.

Ehrlich:1981:SSR

[Ehr81] Gideon Ehrlich. Searching and sorting real numbers. Journal of Algorithms, 2(1):1–12, March 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667748190002X.

Eppstein:1992:MMS

[EIT^{+}92] David Eppstein, Giuseppe F. Italiano, Roberto Tamassia, Robert E. Tarjan, Jeffery Westbrook, and Moti Yung. Maintenance of a minimum spanning forest in a dynamic plane graph. Journal of Algorithms, 13(1):33–54, March 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

019667749290004V. See erratum [EIT^{+}93].

Eppstein:1993:EVN

[EIT^{+}93] D. Eppstein, G. F. Italiano, R. Tamassia, R. E. Tarjan, J. Westbrook, and M. Yung. Erratum: Volume 13, number 1 (1992), in the article “Maintenance of a Minimum Spanning Forest in a Dynamic Plane Graph,” by David Eppstein, Giuseppe F. Italiano, Roberto Tamassia, Robert E. Tarjan, Jeffery Westbrook, and Moti Yung, pages 33–54. Journal of Algorithms, 15(1):173, July 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710369. See [EIT^{+}92].

Elkin:2004:LIR

[EK04a] Michael Elkin and Guy Kortsarz. Logarithmic inapproximability of the radio broadcast problem. Journal of Algorithms, 52(1): 8–25, July 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001901.

Evans:2004:ROB

[EK04b] William Evans and David Kirkpatrick. Restructuring ordered binary trees. Journal of Algorithms, 50(2):168–193, February 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000944.

Engels:2003:TSR

[EKK^{+}03] Daniel W. Engels, David R. Karger, Stavros G. Kolliopoulos, Sudipta Sengupta, R. N. Uma, and Joel Wein. Techniques for scheduling with rejection. Journal of Algorithms, 49(1):175– 191, October 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000786.

Elmasry:2004:PSA

[Elm04] Amr Elmasry. Parameterized self-adjusting heaps. Journal of Algorithms, 52(2):103–119, August 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000434.

Even:1997:CSD

[ELW97] Shimon Even, Ami Litman, and Peter Winkler. Computing with snakes in directed networks of automata. Journal of Algorithms, 24(1):158–170, July 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496908408.

Engel:1995:ADV

[Eng95] K. Engel. An algorithm for the determination of the variance of a partially ordered set. Journal of Algorithms, 19(3):441–448, November 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710462.

Edelsbrunner:1985:BDS

[EO85] Herbert Edelsbrunner and Mark H. Overmars. Batched dynamic solutions to decomposable searching problems. Journal of Algorithms, 6(4):515–542, December 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900306.

Eppstein:1990:SCM

[Epp90] David Eppstein. Sequence comparison with mixed convex and concave costs. Journal of Algorithms, 11(1):85–101, March 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677490900319.

Eppstein:1994:OAD

[Epp94] D. Eppstein. Offline algorithms for dynamic minimum spanning tree problems. Journal of Algorithms, 17(2):237–250, September 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710339.

Eppstein:1997:MRB

[Epp97] David Eppstein. Minimum range balanced cuts via dynamic subset sums. Journal of Algorithms, 23(2):375–385, May 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749690841X.

Eppstein:2000:IDM

[Epp00] David Eppstein. Incremental and decremental maintenance of planar width. Journal of Algorithms, 37(2):570–577, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911076.

Er:1985:CGC

[Er85] M. C. Er. The complexity of the generalised cyclic Towers of Hanoi problem. Journal of Algorithms, 6(3):351–358, September 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677485900045.

Eades:1981:OT

[ES81] Peter Eades and John Staples. On optimal trees. Journal of Algorithms, 2(4):369–384, December 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900353.

Eddy:1995:HMC

[ES95] W. F. Eddy and M. J. Schervish. How many comparisons does Quicksort use? Journal of Algorithms, 19(3):402–431, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710449. For other Quicksort bounds,

see [MH96, FJ02, NR02].

Elmallah:1998:PGR

[ES98] E. S. Elmallah and L. K. Stewart. Polygon graph recognition. Journal of Algorithms, 26(1):101–140, January 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749790883X.

Erlebach:2003:ISA

[ES03] Thomas Erlebach and Frits C. R. Spieksma. Interval selection: Applications, algorithms, and lower bounds. Journal of Algorithms, 46(1):27–53, January 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002912.

Epstein:2003:VAP

[ET03] Leah Epstein and Tamir Tassa. Vector assignment problems: a general framework. Journal of Algorithms, 48(2):360–384, September 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000555.

Etzion:1986:ACA

[Etz86] Tuvi Etzion. An algorithm for constructing m-ary de Bruijn sequences. Journal of Algorithms, 7(3):331–340, September 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677486900258.

Fernandez-Baca:1989:SPP

[FBS89] David Fernández-Baca and Giora Slutzki. Solving parametric problems on trees. Journal of Algorithms, 10(3):381–402, September 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677489900357.

Fernandez-Baca:1997:OPS

[FBS97] David Fernández-Baca and Giora Slutzki. Optimal parametric search on graphs of bounded tree-width. Journal of Algorithms, 22(2):212–240, February 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908160.

Farach-Colton:2000:LRA

[FCL00] Martin Farach-Colton and Vincenzo Liberatore. On local register allocation. Journal of Algorithms, 37(1):37–65, October 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400910952.

Feig:1981:SBF

[Fei81] Ephraim Feig. On systems of bilinear forms whose minimal division-free algorithms are all bilinear. Journal of Algorithms, 2(3):261–281, September 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900262.

Feig:1983:CSB

[Fei83a] Ephraim Feig. Certain systems of bilinear forms whose minimal algorithms are all quadratic. Journal of Algorithms, 4(2):137– 149, June 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748390041X.

Feig:1983:MAB

[Fei83b] Ephraim Feig. Minimal algorithms for bilinear forms may have divisions. Journal of Algorithms, 4(1):81–84, March 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900366.

Ferguson:1987:NAC

[Fer87] Helaman R. P. Ferguson. A noninductive GL(n, Z) algorithm that constructs integral linear relations for n Z-linearly dependent real numbers. Journal of Algorithms, 8(1):131–145, March 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900332.

Ferragina:1997:DTI

[Fer97] Paolo Ferragina. Dynamic text indexing under string updates. Journal of Algorithms, 22(2):296–328, February 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908147.

Fernandes:1998:BAR

[Fer98] Cristina G. Fernandes. A better approximation ratio for the minimum size k-edge-connected spanning subgraph problem. Journal of Algorithms, 28(1):105–124, July 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909312.

Flaxman:2004:ECA

[FFU04] Abraham Flaxman, Alan Frieze, and Eli Upfal. Efficient communication in an ad-hoc network. Journal of Algorithms, 52 (1):1–7, July 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667740400029X.

Flajolet:1980:SOA

[FFV80] P. Flajolet, J. Françon, and J. Vuillemin. Sequence of operations analysis for dynamic data structures. Journal of Algorithms, 1(2): 111–141, June 1980. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677480900206.

Frederickson:1993:NEM

[FG93] G. N. Frederickson and D. J. Guan. Nonpreemptive ensemble motion planning on a tree. Journal of Algorithms, 15(1):29–60, July 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710291.

Fredman:1994:TS

[FG94] M. L. Fredman and D. L. Goldsmith. Three stacks. Journal of Algorithms, 17(1):45–70, July 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710261.

Ferragina:1999:IDT

[FG99] Paolo Ferragina and Roberto Grossi. Improved dynamic text indexing. Journal of Algorithms, 31(2):291–319, May 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909993.

Ferraro:2003:OMM

[FG03] Pascal Ferraro and Christophe Godin. Optimal mappings with minimum number of connected components in tree-to-tree comparison problems. Journal of Algorithms, 48(2):385–406, September 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000798.

Flajolet:2001:CAP

[FGP01] P. Flajolet, X. Gourdon, and D. Panario. The complete analysis of a polynomial factorization algorithm over finite fields. Journal of Algorithms, 40(1):37–81, July 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911587.

Fischer:1992:PAM

[FGS92] Anat Fischer, Itzhak Gilboa, and Moshe Shpitalni. A polynomial algorithm for minimal interval representation. Journal of Algorithms, 13(4):546–563, December 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290055H.

Foata:1998:TW

[FH98] Dominique Foata and Guo-Niu Han. Transformations on words. Journal of Algorithms, 28(1):172–191, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909111.

Fellows:2003:ADM

[FHS03] Michael Fellows, Michael Hallett, and Ulrike Stege. Analogs & duals of the MAST problem for sequences & trees. Journal of Algorithms, 49(1):192–216, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000816.

Fillebrown:1992:FCB

[Fil92] Sandra Fillebrown. Faster computation of Bernoulli numbers. Journal of Algorithms, 13(3):431–445, September 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290048H.

Frederickson:1983:FPC

[FJ83] Greg N. Frederickson and Donald B. Johnson. Finding kth paths and p-centers by generating and searching good data structures. Journal of Algorithms, 4(1):61–80, March 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900354.

Fill:2002:QA

[FJ02] James Allen Fill and Svante Janson. Quicksort asymptotics. Journal of Algorithms, 44(1):4–28, July 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S019667740200216X. For other Quicksort bounds, see [ES95, MH96, NR02].

Frieze:1996:GCH

[FJM^{+}96] Alan Frieze, Mark Jerrum, Michael Molloy, Robert Robinson, and Nicholas Wormald. Generating and counting Hamilton cycles in random regular graphs. Journal of Algorithms, 21(1):176–198, July 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900425.

Fredman:1995:DST

[FJMO95] M. L. Fredman, D. S. Johnson, L. A. Mcgeoch, and G. Ostheimer. Data structures for traveling salesmen. Journal of Algorithms, 18(3):432–479, May 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677485710188.

Farber:1985:DPG

[FK85] Martin Farber and J. Mark Keil. Domination in permutation graphs. Journal of Algorithms, 6(3):309–321, September 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667748590001X.

Fredman:1996:CDM

[FK96] Michael L. Fredman and Leonid Khachiyan. On the complexity of dualization of monotone disjunctive normal forms. Journal of Algorithms, 21(3):618–628, November 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900620.

Fiat:2003:MDS

[FK03] Amos Fiat and Haim Kaplan. Making data structures confluently persistent. Journal of Algorithms, 48(1):16–58, August 2003. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000440.

Fekete:1997:NFT

[FKK^{+}97] Sándor P. Fekete, Samir Khuller, Monika Klemmstein, Balaji Raghavachari, and Neal Young. A network-flow technique for finding low-weight bounded-degree spanning trees. Journal of Algorithms, 24(2):310–324, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908622.

Fiat:1991:CPA

[FKL^{+}91] Amos Fiat, Richard M. Karp, Michael Luby, Lyle A. McGeoch, Daniel D. Sleator, and Neal E. Young. Competitive paging algorithms. Journal of Algorithms, 12(4):685–699, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190041V.

Feige:2002:IAM

[FKL02] Uriel Feige, Marek Karpinski, and Michael Langberg. Improved approximation of Max-Cut on graphs of bounded degree. Journal of Algorithms, 43(2):201–219, May 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402000056.

Faigle:1999:GLA

[FKN99] U. Faigle, W. Kern, and W. M. Nawijn. A greedy online algorithm for the k-track assignment problem. Journal of Algorithms, 31(1):196–210, April 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749891001X.

Friesen:1986:EMB

[FL86] D. K. Friesen and M. A. Langston. Evaluation of a MULTIFITbased scheduling algorithm. Journal of Algorithms, 7(1):35– 59, March 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900374.

Feige:2001:AAM

[FL01] Uriel Feige and Michael Langberg. Approximation algorithms for maximization problems arising in graph partitioning. Journal of Algorithms, 41(2):174–211, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911836.

Feige:2006:RRT

[FL06] Uriel Feige and Michael Langberg. The RPR^{2} rounding technique for semidefinite programs. Journal of Algorithms, 60(1): 1–23, July 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001580.

Fleischer:1999:BCC

[Fle99] Lisa Fleischer. Building chain and cactus representations of all minimum cuts from Hao–Orlin in the same asymptotic run time. Journal of Algorithms, 33(1):51–72, October 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910398.

Frederiksen:2003:DTA

[FLNU03] Jens S. Frederiksen, Kim S. Larsen, John Noga, and Patchrawat Uthaisombut. Dynamic TCP acknowledgment in the LogP model. Journal of Algorithms, 48(2):407–428, September 2003. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000580.

Frieze:1999:SEG

[FM99] Alan M. Frieze and Michael Molloy. Splitting an expander graph. Journal of Algorithms, 33(1):166–172, October 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910234.

Feder:2002:WCT

[FM02] Tomás Feder and Rajeev Motwani. Worst-case time bounds for coloring and satisfiability problems. Journal of Algorithms, 45(2):192–201, November 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002249.

Filaseta:2005:ITL

[FM05] Michael Filaseta and Douglas B. Meade. Irreducibility testing of lacunary 0,1-polynomials. Journal of Algorithms, 55(1):21– 28, April 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001543.

Feder:2007:CSP

[FMO^{+}07] Tomás Feder, Rajeev Motwani, Liadan O’Callaghan, Chris Olston, and Rina Panigrahy. Computing shortest paths with uncertainty. Journal of Algorithms, 62(1):1–18, January 2007. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001166.

Frigioni:2000:FDA

[FMSN00] Daniele Frigioni, Alberto Marchetti-Spaccamela, and Umberto Nanni. Fully dynamic algorithms for maintaining shortest paths trees. Journal of Algorithms, 34(2):251–281, February 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910489.

Frigioni:2003:FDS

[FMSN03] Daniele Frigioni, Alberto Marchetti-Spaccamela, and Umberto Nanni. Fully dynamic shortest paths in digraphs with arbitrary arc weights. Journal of Algorithms, 49(1):86–113, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000828.

Fernau:2001:EEA

[FN01] Henning Fernau and Rolf Niedermeier. An efficient exact algorithm for constraint bipartite vertex cover. Journal of Algorithms, 38(2):374–410, February 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911416.

Feinberg:1989:FFP

[FP89] Ellen B. Feinberg and Christos H. Papadimitriou. Finding feasible paths for a two-point body. Journal of Algorithms, 10(1):109– 119, March 1989. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677489900266.

Flener:2009:SNC

[FP09] Pierre Flener and Justin Pearson. Solving necklace constraint problems. Journal of Algorithms, 64(2–3):61–73, April/July 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677408000333.

Furer:1994:AMD

[FR94] M. Furer and B. Raghavachari. Approximating the minimumdegree Steiner tree to within one of optimal. Journal of Algorithms, 17(3):409–423, November 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667748471042X.

Furer:1995:EPA

[FR95] M. Furer and B. Raghavachari. An efficient parallel algorithm for finding Hamiltonian cycles in dense directed graphs. Journal of Algorithms, 18(2):203–220, March 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710073.

Frank:1981:WMI

[Fra81] András Frank. A weighted matroid intersection algorithm. Journal of Algorithms, 2(4):328–336, December 1981. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900328.

Fredman:1981:SBM

[Fre81] M. L. Fredman. The spanning bound as a measure of range query complexity. Journal of Algorithms, 2(1):77–87, March 1981. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900092.

Frederickson:1994:EF

[Fre94] G. N. Frederickson. Editor foreword. Journal of Algorithms, 17(3):291, November 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710376.

Frederickson:1995:UCG

[Fre95] G. N. Frederickson. Using cellular graph embeddings in solving all pairs shortest paths problems. Journal of Algorithms, 19(1): 45–85, July 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710279.

Frederickson:1997:DSD

[Fre97] Greg N. Frederickson. A data structure for dynamically maintaining rooted trees. Journal of Algorithms, 24(1):37–65, July 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908354.

Frieze:1988:AFH

[Fri88] A. M. Frieze. An algorithm for finding Hamilton cycles in random directed graphs. Journal of Algorithms, 9(2):181–204, June 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900375.

Froda:2000:APR

[Fro00] Sorana Froda. On assessing the performance of randomized algorithms. Journal of Algorithms, 37(2):344–362, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911143.

Filaseta:1994:MEN

[FRW94] M. Filaseta, M. L. Robinson, and F. S. Wheeler. The minimal Euclidean norm of an algebraic number is effectively computable. Journal of Algorithms, 16(2):309–333, March 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710157.

Flajolet:1986:CGE

[FS86] Philippe Flajolet and Nasser Saheb. The complexity of generating an exponentially distributed variate. Journal of Algorithms, 7(4):463–488, December 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900143.

Fernandezbaca:1994:PPG

[FS94] D. Fernandezbaca and G. Slutzki. Parametric problems on graphs of bounded tree-width. Journal of Algorithms, 16(3):408–430, May 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710194.

Frieze:1996:ATS

[FS96] Alan Frieze and Stephen Suen. Analysis of two simple heuristics on a random instance of k- sat. Journal of Algorithms, 20(2):312– 355, March 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900164.

Frederickson:1999:IWM

[FSO99] Greg N. Frederickson and Roberto Solis-Oba. Increasing the weight of minimum spanning trees. Journal of Algorithms, 33(2):244–266, November 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749991026X.

Fingerhut:1997:DLC

[FST97] J. Andrew Fingerhut, Subhash Suri, and Jonathan S. Turner. Designing least-cost nonblocking broadband networks. Journal of Algorithms, 24(2):287–309, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749790866X.

Fu:1997:DGP

[Fu97] James Jianghai Fu. Directed graph pattern matching and topological embedding. Journal of Algorithms, 22(2):372–391, February 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908184.

Fujito:1999:AND

[Fuj99] Toshihiro Fujito. Approximating node-deletion problems for matroidal properties. Journal of Algorithms, 31(1):211–227, April 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909956.

Fredman:1982:CPM

[FV82] Michael L. Fredman and Dennis J. Volper. The complexity of partial match retrieval in a dynamic setting. Journal of Algorithms, 3(1):68–78, March 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900098.

Formann:1993:RTD

[FWW93] M. Formann, D. Wagner, and F. Wagner. Routing through a dense channel with minimum total wire length. Journal of Algorithms, 15(2):267–283, September 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710412.

Franzblau:1982:BPH

[FZ82] D. S. Franzblau and Doron Zeilberger. A bijective proof of the hook-length formula. Journal of Algorithms, 3(4):317–343, December 1982. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677482900293.

Gabow:1995:CRS

[Gab95] H. N. Gabow. Centroids, representations, and submodular flows. Journal of Algorithms, 18(3):586–628, May 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667748571022X.

Gabow:1998:AGP

[Gab98] Harold N. Gabow. Algorithms for graphic polymatroids and parametric s-sets. Journal of Algorithms, 26(1):48–86, January 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497909044.

Galli:2000:ACG

[Gal00] Nicola Galli. Average costs of a graph exploration: Upper and lower bounds. Journal of Algorithms, 34(1):148–176, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910532.

Gavril:1987:GMS

[Gav87] Fănică Gavril. Generating the maximum spanning trees of a weighted graph. Journal of Algorithms, 8(4):592–597, December 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900538.

Guttmann-Beck:1997:AAM

[GBH97] Nili Guttmann-Beck and Refael Hassin. Approximation algorithms for min-max tree partition. Journal of Algorithms, 24(2):266–286, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496908482.

Georgakopoulos:2004:STR

[Geo04] George F. Georgakopoulos. Splay trees: a reweighing lemma and a proof of competitiveness vs. dynamic balanced trees. Journal of Algorithms, 51(1):64–76, April 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001573.

Giancarlo:1997:MDP

[GG97] Raffaele Giancarlo and Roberto Grossi. Multi-dimensional pattern matching with dimensional wildcards: Data structures and optimal on-line search algorithms. Journal of Algorithms, 24(2):223–265, August 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496908445.

Gavoille:1998:WCB

[GG98] Cyril Gavoille and Eric Guévremont. Worst case bounds for shortest path interval routing. Journal of Algorithms, 27(1):1–25, April 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497909159.

Garay:1997:ELC

[GGK^{+}97] Juan A. Garay, Inder S. Gopal, Shay Kutten, Yishay Mansour, and Moti Yung. Efficient on-line call control algorithms. Journal of Algorithms, 23(1):180–194, April 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908214.

Goldberg:2001:CGP

[GGP^{+}01] Leslie Ann Goldberg, Paul W. Goldberg, Mike Paterson, Pavel Pevzner, Süleyman Cenk Sahinalp, and Elizabeth Sweedyk. The complexity of gene placement. Journal of Algorithms, 41(2):225– 243, November 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911721.

Goldberg:1998:CCV

[GGPS98] Leslie Ann Goldberg, Paul W. Goldberg, Cynthia A. Phillips, and Gregory B. Sorkin. Constructing computer virus phylogenies. Journal of Algorithms, 26(1):188–208, January 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749790897X.

Golumbic:1988:SCA

[GH88] Martin Charles Golumbic and Peter L. Hammer. Stability in circular arc graphs. Journal of Algorithms, 9(3):314–320, September 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900235.

Guerra-Hernandez:2009:CAS

[GHCMS09] Alejandro Guerra-Hernández, José Martín Castro-Manzano, and Amal El Fallah Seghrouchni. CTL AgentSpeak(L): a specification language for agent programs. Journal of Algorithms, 64(1):31– 40, January 2009. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677409000066.

Guha:2003:CVC

[GHKO03] Sudipto Guha, Refael Hassin, Samir Khuller, and Einat Or. Capacitated vertex covering. Journal of Algorithms, 48(1):257– 270, August 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000531.

Ghosh:1991:CVP

[Gho91] Subir Kumar Ghosh. Computing the visibility polygon from a convex set and related problems. Journal of Algorithms, 12(1): 75–95, March 1991. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749190024S.

Ghosh:1996:NCV

[Gho96] Subir Kumar Ghosh. A note on computing the visibility polygon from a convex chain. Journal of Algorithms, 21(3):657–662, November 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900644.

Goel:2005:OTC

[GHP05] Ashish Goel, Monika R. Henzinger, and Serge Plotkin. An online throughput-competitive algorithm for multicast routing and admission control. Journal of Algorithms, 55(1):1–20, April 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404001555.

Goel:2003:SDT

[GHPT03] Ashish Goel, Monika R. Henzinger, Serge Plotkin, and Eva Tardos. Scheduling data transfers in a network and the set scheduling problem. Journal of Algorithms, 48(2):314–332, September 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000543.

Gilbert:1984:STG

[GHT84] John R. Gilbert, Joan P. Hutchinson, and Robert Endre Tarjan. A separator theorem for graphs of bounded genus. Journal of Algorithms, 5(3):391–407, September 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900191.

Gil:1999:HPT

[GI99] Joseph Gil and Alon Itai. How to pack trees. Journal of Algorithms, 32(2):108–132, August 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910143.

Gaur:2002:CRA

[GIK02] Daya Ram Gaur, Toshihide Ibaraki, and Ramesh Krishnamurti. Constant ratio approximation algorithms for the Rectangle Stabbing Problem and the Rectilinear Partitioning Problem. Journal of Algorithms, 43(1):138–152, April 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402912216.

Gupta:1995:FRG

[GJS95] P. Gupta, R. Janardan, and M. Smid. Further results on generalized intersection searching problems: Counting, reporting, and dynamization. Journal of Algorithms, 19(2):282–317, September 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710383.

Guha:1999:GSB

[GK99] Sudipto Guha and Samir Khuller. Greedy strikes back: Improved facility location algorithms. Journal of Algorithms, 31(1):228– 248, April 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909932.

Gibbons:1991:TCP

[GKR^{+}91] Phillip Gibbons, Richard Karp, Vijaya Ramachandran, Danny Soroker, and Robert Tarjan. Transitive compaction in parallel via branchings. Journal of Algorithms, 12(1):110–125, March 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190026U.

Garg:2000:PAA

[GKR00] Naveen Garg, Goran Konjevod, and R. Ravi. A polylogarithmic approximation algorithm for the group Steiner tree problem. Journal of Algorithms, 37(1):66–84, October 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910964.

Golumbic:1995:GSP

[GKS95] M. C. Golumbic, H. Kaplan, and R. Shamir. Graph sandwich problems. Journal of Algorithms, 19(3):449–473, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710474.

Gandhi:2004:AAP

[GKS04] Rajiv Gandhi, Samir Khuller, and Aravind Srinivasan. Approximation algorithms for partial covering problems. Journal of Algorithms, 53(1):55–84, October 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000689.

Gabow:2001:UMM

[GKT01] Harold N. Gabow, Haim Kaplan, and Robert E. Tarjan. Unique maximum matching algorithms. Journal of Algorithms, 40(2):159–183, August 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911678.

Galluccio:1996:CPM

[GL96] Anna Galluccio and Martin Loebl. Cycles of prescribed modularity in planar digraphs. Journal of Algorithms, 21(1):51–70, July 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667749690036X.

Galluccio:1998:EDC

[GL98] Anna Galluccio and Martin Loebl. Even directed cycles in H-free digraphs. Journal of Algorithms, 27(1):26–41, April 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749790910X.

Gavril:1994:CPO

[GLD94] F. Gavril, V. T. Laredo, and D. Dewerra. Chordless paths, odd holes, and kernels in graphs without m-obstructions. Journal of Algorithms, 17(2):207–221, September 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710315.

Gum:2005:EM

[GLLF05] Ben Gum, Richard J. Lipton, Andrea LaPaugh, and Faith Fich. Estimating the maximum. Journal of Algorithms, 54(1):105– 114, January 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000677.

Gupta:1983:RUT

[GLW83] U. I. Gupta, D. T. Lee, and C. K. Wong. Ranking and unranking of B-trees. Journal of Algorithms, 4(1):51–60, March 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900342.

Gonnet:1984:ALP

[GM84] Gaston H. Gonnet and J. Ian Munro. The analysis of linear probing sort by the use of a new mathematical transform. Journal of Algorithms, 5(4):451–470, December 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900014.

Gaber:2003:CBM

[GM03] Iris Gaber and Yishay Mansour. Centralized broadcast in multihop radio networks. Journal of Algorithms, 46(1):1–20, January 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402002924.

Gavanelli:2008:REE

[GM08] Marco Gavanelli and Toni Mancini. RCRA 2007: Experimental evaluation of algorithms for solving problems with combinatorial explosion. Journal of Algorithms, 63(1–3):1–2, January/July 2008. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677408000023.

Guha:2003:CFA

[GMM03] Sudipto Guha, Adam Meyerson, and Kamesh Munagala. A constant factor approximation algorithm for the fault-tolerant facility location problem. Journal of Algorithms, 48(2):429–440, September 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000567.

Gupta:1995:PCT

[GN95] A. Gupta and N. Nishimura. The parallel complexity of tree embedding problems. Journal of Algorithms, 18(1):176–200, January 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710061.

Guting:1989:FAD

[GNO89] Ralf-Hartmut Güting, Otto Nurmi, and Thomas Ottmann. Fast algorithms for direct enclosures and direct dominances. Journal of Algorithms, 10(2):170–186, June 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900114.

Goldberg:1982:MSD

[Gol82] Robert N. Goldberg. Minimal string difference encodings. Journal of Algorithms, 3(2):147–156, June 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900141.

Goldberg:1992:EAL

[Gol92] Leslie Ann Goldberg. Efficient algorithms for listing unlabeled graphs. Journal of Algorithms, 13(1):128–143, March 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900092.

Goldberg:1997:EIS

[Gol97] Andrew V. Goldberg. An efficient implementation of a scaling minimum-cost flow algorithm. Journal of Algorithms, 22(1):1– 29, January 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667748570805X.

Goodrich:1989:TPP

[Goo89] Michael T. Goodrich. Triangulating a polygon in parallel. Journal of Algorithms, 10(3):327–351, September 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900321.

Goodrich:2011:RSS

[Goo11] Michael T. Goodrich. Randomized Shellsort: a simple dataoblivious sorting algorithm. Journal of the ACM, 58(6):27:1– 27:??, December 2011. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). For other Shellsort bounds, see [Yao80, Sed86, WS90, Poo93, PS97].

Gordon:1998:SFE

[Gor98] Daniel M. Gordon. A survey of fast exponentiation methods. Journal of Algorithms, 27(1):129–146, April 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909135.

Georgakopoulos:1987:STP

[GP87] George Georgakopoulos and Christos H. Papadimitriou. The 1-Steiner tree problem. Journal of Algorithms, 8(1):122–130, March 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900320.

Grable:2000:FDA

[GP00] David A. Grable and Alessandro Panconesi. Fast distributed algorithms for Brooks–Vizing colorings. Journal of Algorithms, 37(1):85–120, October 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400910976.

Garefalakis:2002:PFF

[GP02] Theodoulos Garefalakis and Daniel Panario. Polynomials over finite fields free from large and small degree irreducible factors. Journal of Algorithms, 44(1):98–120, July 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002079.

Gavoille:2004:DLG

[GPPR04] Cyril Gavoille, David Peleg, Stéphane Pérennes, and Ran Raz. Distance labeling in graphs. Journal of Algorithms, 53(1):85– 112, October 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000884.

Goldman:2000:OSH

[GPS00] Sally A. Goldman, Jyoti Parwatikar, and Subhash Suri. Online scheduling with hard deadlines. Journal of Algorithms, 34(2):370– 389, February 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667749991060X.

Goldberg:1993:STP

[GPV93] A. V. Goldberg, S. A. Plotkin, and P. M. Vaidya. Sublineartime parallel algorithms for matching and related problems. Journal of Algorithms, 14(2):180–213, March 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710096.

Gazit:1998:RPA

[GR98] Hillel Gazit and John H. Reif. A randomized parallel algorithm for planar graph isomorphism. Journal of Algorithms, 28(2):290– 314, August 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909439.

Greenberg:1988:SLD

[Gre88] Harold Greenberg. Solution to a linear Diophantine equation for nonnegative integers. Journal of Algorithms, 9(3):343–353, September 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900259.

Grolmusz:2002:CSS

[Gro02] Vince Grolmusz. Constructing set systems with prescribed intersection sizes. Journal of Algorithms, 44(2):321–337, August 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402002043.

Gurari:1984:IDP

[GS84] Eitan M. Gurari and Ivan Hal Sudborough. Improved dynamic programming algorithms for bandwidth minimization and the MinCut Linear Arrangement problem. Journal of Algorithms, 5(4):531–546, December 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900063.

Gavril:1985:CTP

[GS85] Fanica Gavril and Johanan Schönheim. Constructing trees with prescribed cardinalities for the components of their vertex deleted subgraphs. Journal of Algorithms, 6(2):239–252, June 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900410.

Goemans:2004:CFL

[GS04] Michel X. Goemans and Martin Skutella. Cooperative facility location games. Journal of Algorithms, 50(2):194–214, February 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000981.

Gabow:1984:EAF

[GT84] Harold N. Gabow and Robert E. Tarjan. Efficient algorithms for a family of matroid intersection problems. Journal of Algorithms, 5(1):80–131, March 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677484900427.

Gabow:1988:ATB

[GT88] Harold N. Gabow and Robert E. Tarjan. Algorithms for two bottleneck optimization problems. Journal of Algorithms, 9(3):411– 417, September 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900314.

Goodrich:1997:DRS

[GT97] Michael T. Goodrich and Roberto Tamassia. Dynamic ray shooting and shortest paths in planar subdivisions via balanced geodesic triangulations. Journal of Algorithms, 23(1):51–73, April 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485707973.

Goldberg:2001:CTA

[GT01] Andrew V. Goldberg and Kostas Tsioutsiouliklis. Cut tree algorithms: an experimental study. Journal of Algorithms, 38(1): 51–83, January 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911362.

Guenoche:1983:RST

[Gué83] A. Guénoche. Random spanning tree. Journal of Algorithms, 4(3):214–220, September 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677483900226.

Guibas:1980:Pa

[Gui80a] Leo J. Guibas. Problems. Journal of Algorithms, 1(1):107– 110, March 1980. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677480900097.

Guibas:1980:Pb

[Gui80b] Leo J. Guibas. Problems. Journal of Algorithms, 1(2):209–212, June 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677480900231.

Guibas:1980:Pc

[Gui80c] Leo J. Guibas. Problems. Journal of Algorithms, 1(4):394–395, December 1980. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677480900188.

Guibas:1981:Pa

[Gui81a] Leo J. Guibas. Problems. Journal of Algorithms, 2(1):103– 104, March 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900110.

Guibas:1981:Pb

[Gui81b] Leo J. Guibas. Problems. Journal of Algorithms, 2(2):208–210, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677481900225.

Guibas:1981:Pc

[Gui81c] Leo J. Guibas. Problems. Journal of Algorithms, 2(3):314–315, September 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900304.

Guibas:1982:Pa

[Gui82a] Leo J. Guibas. Problems. Journal of Algorithms, 3(2):177–181, June 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677482900177.

Guibas:1982:Pb

[Gui82b] Leo J. Guibas. Problems. Journal of Algorithms, 3(4):362–380, December 1982. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677482900311.

Guibas:1983:Pa

[Gui83a] Leo J. Guibas. Problems. Journal of Algorithms, 4(1):85– 86, March 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900378.

Guibas:1983:Pb

[Gui83b] Leo J. Guibas. Problems. Journal of Algorithms, 4(2):176–188, June 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677483900445.

Guibas:1983:Pc

[Gui83c] Leo J. Guibas. Problems. Journal of Algorithms, 4(4):382–396, December 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900184.

Guibas:1984:Pa

[Gui84a] Leo J. Guibas. Problems. Journal of Algorithms, 5(1):145– 146, March 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900440.

Guibas:1984:Pb

[Gui84b] Leo J. Guibas. Problems. Journal of Algorithms, 5(4):579–594, December 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900099.

Guibas:1985:P

[Gui85] Leo J. Guibas. Problems. Journal of Algorithms, 6(2):283–290, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677485900458.

Gupta:2001:IBA

[Gup01] Anupam Gupta. Improved bandwidth approximation for trees and chordal graphs. Journal of Algorithms, 40(1):24–36, July 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911180.

Gusfield:1984:BNM

[Gus84] Dan Gusfield. Bounds for naive multiple machine scheduling with release times and deadlines. Journal of Algorithms, 5(1): 1–6, March 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748490035X.

Guting:1984:OCA

[Güt84] Ralf Hartmut Güting. An optimal contour algorithm for isooriented rectangles. Journal of Algorithms, 5(3):303–326, September 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677484900130.

Garg:2004:MCN

[GVY04] Naveen Garg, Vijay V. Vazirani, and Mihalis Yannakakis. Multiway cuts in node weighted graphs. Journal of Algorithms, 50(1): 49–61, January 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001111.

Gupta:1997:CPL

[GW97] Himanshu Gupta and Rephael Wenger. Constructing piecewise linear homeomorphisms of simple polygons. Journal of Algorithms, 22(1):142–157, January 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485708085.

Goodrich:2000:FDP

[GW00] Michael T. Goodrich and Christopher G. Wagner. A framework for drawing planar graphs with curves and polylines. Journal of Algorithms, 37(2):399–421, November 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911155.

Graham:1983:FCH

[GY83] Ronald L. Graham and F. Frances Yao. Finding the convex hull of a simple polygon. Journal of Algorithms, 4(4):324–331, December 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677483900135.

Goldberg:1994:FPS

[GZ94] T. Goldberg and U. Zwick. Faster parallel string matching via larger deterministic samples. Journal of Algorithms, 16(2):295– 308, March 1994. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677484710145.

Hagerup:1995:FDP

[Hag95] T. Hagerup. Fast deterministic processor allocation. Journal of Algorithms, 18(3):629–649, May 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677485710231.

Haldar:1997:APS

[Hal97a] S. Haldar. An ‘All Pairs Shortest Paths’ distributed algorithm using 2n^{2} messages. Journal of Algorithms, 24(1):20–36, July 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908421.

Halldorsson:1997:PLG

[Hal97b] Magnus M. Halldórsson. Parallel and on-line graph coloring. Journal of Algorithms, 23(2):265–280, May 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908366.

Han:2004:DST

[Han04] Yijie Han. Deterministic sorting in O(n log log n) time and linear space. Journal of Algorithms, 50(1):96–105, January 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740300155X.

Hartvigsen:1993:MPB

[Har93] D. Hartvigsen. Minimum path bases. Journal of Algorithms, 15(1):125–142, July 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677483710333.

Hare:1997:CPB

[Har97] D. E. G. Hare. Computing the principal branch of log-Gamma. Journal of Algorithms, 25(2):221–236, November 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908816.

Haastad:1990:TRN

[Hås90]  Johan Håstad.  Tensor rank is NP-complete.  Journal

of Algorithms, 11(4):644–654, December 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900146.

Havet:2000:FOH

[Hav00] Frédéric Havet. Finding an oriented Hamiltonian path in a tournament. Journal of Algorithms, 36(2):253–275, August 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740091085X.

He:1991:EPA

[He91a] Xin He. Efficient parallel algorithms for series parallel graphs. Journal of Algorithms, 12(3):409–430, September 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190012N.

He:1991:IAP

[He91b] Xin He. An improved algorithm for the planar 3-cut problem. Journal of Algorithms, 12(1):23–37, March 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190021P.

He:1993:PAC

[He93] X. He. Parallel algorithm for cograph recognition with applications. Journal of Algorithms, 15(2):284–313, September 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710424.

He:2001:SLT

[He01] Xin He. A simple linear time algorithm for proper box rectangular drawings of plane graphs. Journal of Algorithms, 40(1): 82–101, July 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911617.

Hebert:1998:CIQ

[Heb98] D. J. Hebert. Cyclic interlaced quadtree algorithms for quincunx multiresolution. Journal of Algorithms, 27(1):97–128, April 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909257.

Heitsch:2005:IFK

[Hei05] Christine E. Heitsch. Insufficiency of four known necessary conditions on string unavoidability. Journal of Algorithms, 56(2):96– 123, August 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001403.

Henzinger:1997:SAA

[Hen97] Monika Rauch Henzinger. A static 2-approximation algorithm for vertex connectivity and incremental approximation algorithms for edge and vertex connectivity. Journal of Algorithms, 24(1):

194–220, July 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908555.

Hershberger:1988:SPA

[HG88] John Hershberger and Leonidas J. Guibas. An O(n^{2}) shortest path algorithm for a non-rotating convex body. Journal of Algorithms, 9(1):18–46, March 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667748890003X.

Hentzel:1981:ECS

[HH81] Irvin Roy Hentzel and Leslie Hogben. Exhaustive checking of sparse algebras. Journal of Algorithms, 2(1):44–49, March 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677481900067.

Hester:1986:FCO

[HHHW86] J. H. Hester, D. S. Hirschberg, S.-H. S. Huang, and C. K. Wong. Faster construction of optimal binary split trees. Journal of Algorithms, 7(3):412–424, September 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900313.

Hester:1988:COB

[HHL88] J. H. Hester, D. S. Hirschberg, and L. L. Larmore. Construction of optimal binary split trees in the presence of bounded access probabilities. Journal of Algorithms, 9(2):245–253, June 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900405.

Hubalek:2002:MVR

[HHL^{+}02] Friedrich Hubalek, Hsien-Kuei Hwang, William Lew, Hosam Mahmoud, and Helmut Prodinger. A multivariate view of random bucket digital search trees. Journal of Algorithms, 44(1): 121–158, July 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002109.

Haastad:2003:FPR

[HIL03] Johan Håstad, Lars Ivansson, and Jens Lagergren. Fitting points on the real line and its application to RH mapping. Journal of Algorithms, 49(1):42–62, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740300083X.

Hoffman:2005:GHS

[HJW05] Dean Hoffman, Peter Johnson, and Nadine Wilson. Generating Huffman sequences. Journal of Algorithms, 54(1):115–121, January 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404000653.

Hofri:1986:SAN

[HK86] Micha Hofri and Sami Kamhi. A stochastic analysis of the NFD bin-packing algorithm. Journal of Algorithms, 7(4):489– 509, December 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900155.

Hell:1993:ADC

[HK93] P. Hell and D. G. Kirkpatrick. Algorithms for degree constrained graph factors of minimum deficiency. Journal of Algorithms, 14(1):115–138, January 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710060.

Hassin:2001:A

[HK01] Refael Hassin and Samir Khuller. z-approximations. Journal of Algorithms, 41(2):429–442, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911873.

Halldorsson:2002:TMA

[HK02] Magnús M. Halldórsson and Guy Kortsarz. Tools for multicoloring with applications to planar graphs and partial k-trees. Journal of Algorithms, 42(2):334–366, February 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401912106.

Higham:1997:OAP

[HKAA97] Lisa Higham, David Kirkpatrick, Karl Abrahamson, and Andrew Adler. Optimal algorithms for probabilistic solitude detection on anonymous rings. Journal of Algorithms, 23(2):291–328, May 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908287.

Hirschberg:1992:TP

[HL92] D. S. Hirschberg and L. L. Larmore. The Traveler’s Problem. Journal of Algorithms, 13(1):148–160, March 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290011Z.

Hassin:2003:MST

[HL03] Refael Hassin and Asaf Levin. Minimum spanning tree with hop restrictions. Journal of Algorithms, 48(1):220–238, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000518.

Harvey:2006:SMB

[HLLT06] Nicholas J. A. Harvey, Richard E. Ladner, László Lovász, and Tami Tamir. Semi-matchings for bipartite graphs and load balancing. Journal of Algorithms, 59(1):53–78, April 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000209.

Ho:1994:MMW

[HLW94] K. I. J. Ho, J. Y. T. Leung, and W. D. Wei. Minimizing maximum weighted error for imprecise computation tasks. Journal of Algorithms, 16(3):431–452, May 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710200.

Halperin:2004:MCC

[HLZ04] Eran Halperin, Dror Livnat, and Uri Zwick. MAX CUT in cubic graphs. Journal of Algorithms, 53(2):169–185, November 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404000902.

Hassin:1985:OAF

[HM85] Refael Hassin and Nimrod Megiddo. An optimal algorithm for finding all the jumps of a monotone step-function. Journal of Algorithms, 6(2):265–274, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900434.

Hochbaum:1987:FAA

[HM87] Dorit S. Hochbaum and Wolfgang Maass. Fast approximation algorithms for a nonconvex covering problem. Journal of Algorithms, 8(3):305–323, September 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900125.

Hafner:1989:DRT

[HM89] James Lee Hafner and Kevin S. McCurley. On the distribution of running times of certain integer factoring algorithms. Journal of Algorithms, 10(4):531–556, December 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900047.

Hayward:1991:ACA

[HM91] Ryan Hayward and Colin McDiarmid. Average case analysis of heap building by repeated insertion. Journal of Algorithms, 12(1):126–153, March 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190027V.

Hui:1993:USS

[HM93] L. C. K. Hui and C. Martel. Unsuccessful search in self-adjusting data structures. Journal of Algorithms, 15(3):447–481, November 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710497.

Heun:1996:NEA

[HM96] Volker Heun and Ernst W. Mayr. A new efficient algorithm for embedding an arbitrary binary tree into its optimal hypercube. Journal of Algorithms, 20(2):375–399, March 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900188.

Heun:2002:EDE

[HM02a] Volker Heun and Ernst W. Mayr. Efficient dynamic embeddings of binary trees into hypercubes. Journal of Algorithms, 43(1): 51–84, April 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402912186.

Heun:2002:EGB

[HM02b] Volker Heun and Ernst W. Mayr. Embedding graphs with bounded treewidth into their optimal hypercubes. Journal of Algorithms, 43(1):17–50, April 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402912174.

Heidrich-Meisner:2009:NSE

[HMI09] Verena Heidrich-Meisner and Christian Igel. Neuroevolution strategies for episodic reinforcement learning. Journal of Algorithms, 64(4):152–168, October 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677409000364.

Hagerup:2001:DD

[HMP01] Torben Hagerup, Peter Bro Miltersen, and Rasmus Pagh. Deterministic dictionaries. Journal of Algorithms, 41(1):69–85, October 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740191171X.

Hunt:1998:NAS

[HMR^{+}98] Harry B. Hunt III, Madhav V. Marathe, Venkatesh Radhakrishnan, S. S. Ravi, Daniel J. Rosenkrantz, and Richard E. Stearns. NC-approximation schemes for NP- and PSPACE-hard problems for geometric graphs. Journal of Algorithms, 26(2):238– 274, February 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497909032.

Heusinger:1989:SC

[HN89] H. Heusinger and H. Noltemeier. On separable clusterings. Journal of Algorithms, 10(2):212–227, June 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900138.

Hochbaum:1996:AAN

[HN96] Dorit S. Hochbaum and Joseph (Seffi) Naor. Approximation algorithms for network design problems on bounded subsets. Journal of Algorithms, 21(2):403–414, September 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900504.

Hochbaum:1986:BTB

[HNS86] Dorit S. Hochbaum, Takao Nishizeki, and David B. Shmoys. A better than “best possible” algorithm to edge color multigraphs. Journal of Algorithms, 7(1):79–104, March 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900398.

Habib:1997:GCI

[HNS97] Michel Habib, Lhouari Nourine, and George Steiner. Gray codes for the ideals of interval orders. Journal of Algorithms, 25(1):52– 66, October 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908634.

Halperin:2002:CCG

[HNZ02] Eran Halperin, Ram Nathaniel, and Uri Zwick. Coloring k-colorable graphs using relatively small palettes. Journal of Algorithms, 45(1):72–90, October 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002171.

Hao:1994:FAF

[HO94] J. X. Hao and J. B. Orlin. A faster algorithm for finding the minimum cut in a directed graph. Journal of Algorithms, 17(3):424– 446, November 1994. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677484710431.

Hochbaum:1998:ACB

[Hoc98] Dorit S. Hochbaum. Approximating clique and biclique problems. Journal of Algorithms, 29(1):174–200, October 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909646.

Hofri:1984:PAN

[Hof84] Micha Hofri. A probabilistic analysis of the Next-Fit bin packing algorithm. Journal of Algorithms, 5(4):547–556, December 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677484900075.

Hoogeveen:1996:SMS

[Hoo96] J. A. Hoogeveen. Single-machine scheduling to minimize a function of two or three maximum cost criteria. Journal of Algorithms, 21(2):415–433, September 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900516.

Hell:1983:CFG

[HR83] Pavol Hell and Moshe Rosenfeld. The complexity of finding generalized paths in tournaments. Journal of Algorithms, 4(4):303– 309, December 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900111.

Huang:1999:ISM

[HR99] Ming-Deh A. Huang and Ashwin J. Rao. Interpolation of sparse multivariate polynomials over large finite fields with applications. Journal of Algorithms, 33(2):204–228, November 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910453.

Henzinger:2000:CVC

[HRG00] Monika R. Henzinger, Satish Rao, and Harold N. Gabow. Computing vertex connectivity: New bounds from old techniques.

Journal of Algorithms, 34(2):222–250, February 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910556.

Hamidoune:1989:PSL

[HRJ^{+}89] Yahya Ould Hamidoune, David Roeder, Steven Janke, Todd Feil, and Richard Koo. The probability of splitters in a list. Journal of Algorithms, 10(1):151–154, March 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900291.

Helvig:2003:MTT

[HRZ03] C. S. Helvig, Gabriel Robins, and Alex Zelikovsky. The moving-target traveling salesman problem. Journal of Algorithms, 49(1):153–174, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000750.

Hu:1981:AFN

[HS81] T. C. Hu and M. T. Shing. An O(n) algorithm to find a near-optimum partition of a convex polygon. Journal of Algorithms, 2(2):122–138, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900146.

Hu:1983:MFO

[HS83] T. C. Hu and M. T. Shing. Multiterminal flows in outerplanar networks. Journal of Algorithms, 4(3):241–261, September 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748390024X.

Hakimi:1984:AAS

[HS84] S. Louis Hakimi and Edward F. Schmeichel. An adaptive algorithm for system level diagnosis. Journal of Algorithms, 5(4):526– 530, December 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900051.

Hershberger:1991:FTP

[HS91a] John Hershberger and Subhash Suri. Finding tailored partitions. Journal of Algorithms, 12(3):431–463, September 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190013O.

Hofri:1991:SOL

[HS91b] Micha Hofri and Hadas Shachnai. Self-organizing lists and independent references: a statistical synergy. Journal of Algorithms, 12(4):533–555, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190032T.

Hershberger:1995:PAR

[HS95a] J. Hershberger and S. Suri. A pedestrian approach to ray shooting: Shoot a ray, take a walk. Journal of Algorithms, 18(3): 403–431, May 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710176.

Hsu:1995:ISC

[HS95b] W. L. Hsu and J. P. Spinrad. Independent sets in circulararc graphs. Journal of Algorithms, 19(2):145–160, September 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710310.

Hershberger:1996:LMP

[HS96] John Hershberger and Subhash Suri. Off-line maintenance of planar configurations. Journal of Algorithms, 21(3):453–475, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900541.

Heydari:1997:DPN

[HS97] Mohammad H. Heydari and I. Hal Sudborough. On the Diameter of the Pancake Network. Journal of Algorithms, 25(1):67– 94, October 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908749.

Halperin:2002:CPO

[HSG02] Dan Halperin, Micha Sharir, and Ken Goldberg. The 2-center problem with obstacles. Journal of Algorithms, 42(1):109–134, January 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911940.

Ho:1998:RDE

[HSH^{+}98] Tung-Yang Ho, Ting-Yi Sung, Lih-Hsing Hsu, Chang-Hsiung Tsai, and Jeng-Yan Hwang. The recognition of double Euler trails in series-parallel networks. Journal of Algorithms, 28(2):216– 257, August 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909427.

Hsu:2002:STC

[Hsu02] Wen-Lian Hsu. A simple test for the consecutive ones property. Journal of Algorithms, 43(1):1–16, April 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401912052.

Huang:1990:SVU

[HT90] Ming-Deh A. Huang and Shang-Hua Teng. Security, verifiability, and universality in distributed computing. Journal of Algorithms, 11(3):492–521, September 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900238.

Hambrusch:1997:EWR

[HT97] Susanne E. Hambrusch and Hung-Yi Tu. Edge weight reduction problems in directed acyclic graphs. Journal of Algorithms, 24 (1):66–93, July 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908567.

Huang:1987:OMS

[Hua87] Shou-Hsuan Stephen Huang. Optimal multiway split trees. Journal of Algorithms, 8(1):146–156, March 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900344.

Huang:1990:NAG

[Hua90] Yuejiang Huang. A new algorithm for the generation of binary de Bruijn sequences. Journal of Algorithms, 11(1):44–51, March 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749090028D.

Huang:1991:FPF

[Hua91a] Ming-Deh A. Huang. Factorization of polynomials over finite fields and decomposition of primes in algebraic number fields. Journal of Algorithms, 12(3):482–489, September 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190015Q.

Huang:1991:GRH

[Hua91b] Ming-Deh A. Huang. Generalized Riemann Hypothesis and factoring polynomials over finite fields. Journal of Algorithms, 12(3):464–481, September 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190014P.

Hvalica:1996:BFS

[Hva96] Dusan Hvalica. Best first search algorithm in AND/OR graphs with cycles. Journal of Algorithms, 21(1):102–110, July 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900395.

Huang:1984:OBS

[HW84] Shou-Hsuan Stephen Huang and C. K. Wong. Optimal binary split trees. Journal of Algorithms, 5(1):69–79, March 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900415.

Hwang:1992:SNU

[HW92] F. K. Hwang and J. F. Weng. The shortest network under a given topology. Journal of Algorithms, 13(3):468–488, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290050M.

Hofting:2000:PTA

[HW00a] F. Höfting and E. Wanke. Polynomial-time analysis of toroidal periodic graphs. Journal of Algorithms, 34(1):14–39, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910465.

Huang:2000:EHI

[HW00b] Ming-Deh Huang and Yiu-Chung Wong. Extended Hilbert irreducibility and its applications. Journal of Algorithms, 37(1):121– 145, October 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400910988.

He:1988:BTA

[HY88] Xin He and Yaacov Yesha. Binary tree algebraic computation and parallel algorithms for simple graphs. Journal of Algorithms, 9(1): 92–113, March 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900077.

Halperin:2001:AAM

[HZ01a] Eran Halperin and Uri Zwick. Approximation algorithms for MAX 4-SAT and rounding procedures for semidefinite programs. Journal of Algorithms, 40(2):184–211, August 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911629.

Halperin:2001:ORE

[HZ01b] Shay Halperin and Uri Zwick. Optimal randomized EREW PRAM algorithms for finding spanning forests. Journal of Algorithms, 39(1):1–46, April 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911465.

Imai:1983:FCC

[IA83] Hiroshi Imai and Takao Asano. Finding the connected components and a maximum clique of an intersection graph of rectangles in the plane. Journal of Algorithms, 4(4):310–323, December 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677483900123.

Imai:1987:DOS

[IA87] Hiroshi Imai and Takao Asano. Dynamic orthogonal segment intersection search. Journal of Algorithms, 8(1):1–18, March 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900241.

Iwama:1996:CGN

[II96] Kazuo Iwama and Chuzo Iwamoto. α-connectivity: a gradually nonparallel graph problem. Journal of Algorithms, 20(3):526– 544, 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

Iwama:1994:SPA

[IK94] K. Iwama and Y. Kambayashi. A simpler parallel algorithm for graph connectivity. Journal of Algorithms, 16(2):190–217, March 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710091.

Iturriaga:2003:ELA

[IL03] Claudia Iturriaga and Anna Lubiw. Elastic labels around the perimeter of a map. Journal of Algorithms, 47(1):14–39, April 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740300004X.

Iliopoulos:1982:AAC

[Ili82] Costas S. Iliopoulos. Analysis of an algorithm for composition of binary quadratic forms. Journal of Algorithms, 3(2):157–159, June 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677482900153.

Iwama:2001:LBE

[IM01a] Kazuo Iwama and Eiji Miyano. A lower bound for elementary oblivious routing on three-dimensional meshes. Journal of Algorithms, 39(2):145–161, May 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911520.

Iwama:2001:ORA

[IM01b] Kazuo Iwama and Eiji Miyano. An O(N) oblivious routing algorithm for two-dimensional meshes of constant queue-size. Journal of Algorithms, 41(2):262–279, November 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911769.

Irving:2002:SRP

[IM02] Robert W. Irving and David F. Manlove. The Stable Roommates Problem with ties. Journal of Algorithms, 43(1):85–105, April 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402912198.

Ibarra:1982:GFM

[IMH82] Oscar H. Ibarra, Shlomo Moran, and Roger Hui. A generalization of the fast LUP matrix decomposition algorithm and applications. Journal of Algorithms, 3(1):45–56, March 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900074.

Iwama:1996:RPM

[IMK96] Kazuo Iwama, Eiji Miyano, and Yahiko Kambayashi. Routing problems on the mesh of buses. Journal of Algorithms, 20(3): 613–631, May 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900309.

Indyk:2001:SAM

[Ind01] Piotr Indyk. A small approximately min-wise independent family of hash functions. Journal of Algorithms, 38(1):84–90, January 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911313.

Itai:1985:STN

[IR85] Alon Itai and Michael Rodeh. Scheduling transmissions in a network. Journal of Algorithms, 6(3):409–429, September 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900100.

Irving:1985:EAS

[Irv85] Robert W. Irving. An efficient algorithm for the “stable roommates” problem. Journal of Algorithms, 6(4):577–595, December 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677485900331.

Itai:1996:ASR

[IS96] Alon Itai and Hadas Shachnai. Adaptive source routing in highspeed networks. Journal of Algorithms, 20(2):218–243, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900127.

Ibarra:1993:EPA

[IWJ93] O. H. Ibarra, H. Wang, and T. Jiang. On efficient parallel algorithms for solving set recurrence equations. Journal of Algorithms, 14(2):244–257, March 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710126.

Ibarra:1994:SEA

[IZ94] O. H. Ibarra and Q. Zheng. Some efficient algorithms for permutation graphs. Journal of Algorithms, 16(3):453–469, May 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710212.

Jansen:2000:ARO

[Jan00] Klaus Jansen. Approximation results for the optimum cost chromatic partition problem. Journal of Algorithms, 34(1):54–89, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910222.

Jerrum:1986:CRP

[Jer86] Mark Jerrum. A compact representation for permutation groups. Journal of Algorithms, 7(1):60–78, March 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900386.

Jacobs:2000:CRE

[JJ00] David P. Jacobs and Robert E. Jamison. Complexity of recognizing equal unions in families of sets. Journal of Algorithms, 37(2):495–504, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911210.

Jiang:2000:SOP

[JKL00] Tao Jiang, Paul Kearney, and Ming Li. Some open problems in computational molecular biology. Journal of Algorithms, 34(1):194–201, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677499910507.

Janssen:2000:DOF

[JKNS00] Jeannette Janssen, Danny Krizanc, Lata Narayanan, and Sunil Shende. Distributed online frequency assignment in cellular networks. Journal of Algorithms, 36(2):119–151, August 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910684.

Jakoby:2006:SEA

[JLR06] Andreas Jakoby, Maciej Liśkiewicz, and Rüdiger Reischuk. Space efficient algorithms for directed series-parallel graphs. Journal of Algorithms, 60(2):85–114, August 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001087.

Johnson:1995:PAC

[JM95] D. B. Johnson and P. Metaxas. A parallel algorithm for computing minimum spanning trees. Journal of Algorithms, 19(3):383– 401, November 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710437.

Jadhav:1996:OAI

[JMB96] Shreesh Jadhav, Asish Mukhopadhyay, and Binay Bhattacharya. An optimal algorithm for the intersection radius of a set of convex polygons. Journal of Algorithms, 20(2):244–267, March 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900139.

Jonker:2009:FFQ

[JMP09] Hugo Jonker, Sjouke Mauw, and Jun Pang. A formal framework for quantifying voter-controlled privacy. Journal of Algorithms, 64(2–3):89–105, April/July 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677409000091.

Jain:2002:PDS

[JMVW02] Kamal Jain, Ion Mandoiu, Vijay V. Vazirani, and David P. Williamson. A primal-dual schema based approximation algorithm for the element connectivity problem. Journal of Algorithms, 45(1):1–15, October 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002225.

Jarvisalo:2008:ESB

[JN08] Matti Järvisalo and Ilkka Niemelä. The effect of structural branching on the efficiency of clause learning SAT solving: an experimental study. Journal of Algorithms, 63(1–3):90–113, January/July 2008. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677408000084.

Johnson:1980:SLM

[Joh80] Ellis L. Johnson. Subadditive lifting methods for partitioning and knapsack problems. Journal of Algorithms, 1(1):75–96, March 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677480900061.

Johnson:1981:NCC

[Joh81] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 2(4):393–405, December 1981. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900377.

Johnson:1982:NCCa

[Joh82a] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 3(1):89–99, March 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900116.

Johnson:1982:NCCb

[Joh82b] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 3(2):182–195, June 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900189.

Johnson:1982:NCCc

[Joh82c] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 3(3):288–300, September 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900268.

Johnson:1982:NCCd

[Joh82d] David S. Johnson. The NP-completeness column: an ongoing gulde. Journal of Algorithms, 3(4):381–395, December 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900323.

Johnson:1983:NCCa

[Joh83a] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 4(1):87–100, March 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748390038X.

Johnson:1983:NCCb

[Joh83b] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 4(2):189–203, June 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900457.

Johnson:1983:NCCc

[Joh83c] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 4(3):286–300, September 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900287.

Johnson:1983:NCCd

[Joh83d] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 4(4):397–411, December 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900196.

Johnson:1984:NCCa

[Joh84a] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 5(1):147–160, March 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900452.

Johnson:1984:NCCb

[Joh84b] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 5(2):284–299, June 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900324.

Johnson:1984:NCCc

[Joh84c] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 5(3):433–447, September 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900221.

Johnson:1984:NCCd

[Joh84d] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 5(4):595–609, December 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900105.

Johnson:1985:NCCa

[Joh85a] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 6(1):145–159, March 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900252.

Johnson:1985:NCCb

[Joh85b] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 6(2):291–305, June 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748590046X.

Johnson:1985:NCCc

[Joh85c] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 6(3):434–451, September 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900124.

Johnson:1986:NCCa

[Joh86a] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 7(2):289–305, June 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900118.

Johnson:1986:NCCb

[Joh86b] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 7(4):584–601, December 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900209.

Johnson:1987:NCCa

[Joh87a] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 8(2):285–303, June 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900435.

Johnson:1987:NCCb

[Joh87b] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 8(3):438–448, September 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900216.

Johnson:1988:NCC

[Joh88] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 9(3):426–444, September 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900338.

Johnson:1990:NCC

[Joh90] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 11(1):144–151, March 1990. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090035D.

Johnson:1992:EF

[Joh92a] David S. Johnson. Editor’s foreword. Journal of Algorithms, 13 (1):1, March 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290002T.

Johnson:1992:NCC

[Joh92b] David S. Johnson. The NP-completeness column: an ongoing guide. Journal of Algorithms, 13(3):502–524, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290052E.

Jansen:2002:PTA

[JP02] Klaus Jansen and Lorant Porkolab. Polynomial time approximation schemes for general multiprocessor job shop scheduling. Journal of Algorithms, 45(2):167–191, November 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002481.

Jeronimo:2001:CCF

[JPS01] Gabriela Jeronimo, Susana Puddu, and Juan Sabia. Computing Chow forms and some applications. Journal of Algorithms, 41(1): 52–68, October 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911770.

Joichi:1993:AI

[JS93] J. Joichi and D. Stanton. An algorithmic involution for p(n). Journal of Algorithms, 15(2):223–228, September 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710394.

Johnson:1990:RDS

[JZ90] David B. Johnson and Willy Zwaenepoel. Recovery in distributed systems using optimistic message logging and checkpointing. Journal of Algorithms, 11(3):462–491, September 1990. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900227.

Jia:2004:EPA

[JZC04] Weijia Jia, Chuanlin Zhang, and Jianer Chen. An efficient parameterized algorithm for m-set packing. Journal of Algorithms, 50(1):106–117, January 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001548.

Kaminski:1988:APM

[Kam88] Michael Kaminski. An algorithm for polynomial multiplication that does not depend on the ring constants. Journal of Algorithms, 9(1):137–147, March 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900107.

Kantor:1985:PTA

[Kan85] William M. Kantor. Polynomial-time algorithms for finding elements of prime order and Sylow subgroups. Journal of Algorithms, 6(4):478–514, December 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748590029X.

Kantor:1990:FSN

[Kan90] William M. Kantor. Finding Sylow normalizers in polynomial time. Journal of Algorithms, 11(4):523–563, December 1990. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900094.

Kant:1996:AOG

[Kan96] Goos Kant. Augmenting outerplanar graphs. Journal of Algorithms, 21(1):1–25, July 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900346.

Khuller:1999:OPP

[KAO99] Samir Khuller, Pankaj K. Agarwal, and Joseph O’Rourke. Open problems presented at SCG’98. Journal of Algorithms, 30(2):449– 453, February 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909798.

Karloff:2000:F

[Kar00] Howard Karloff. Foreword. Journal of Algorithms, 37(1):1, October 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400910940.

Kearney:1998:TP

[KC98] Paul E. Kearney and Derek G. Corneil. Tree powers. Journal of Algorithms, 29(1):111–131, October 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498999990.

Knott:1989:HTC

[KdlT89] Gary D. Knott and Pilar de la Torre. Hash table collision resolution with direct chaining. Journal of Algorithms, 10(1):20– 34, March 1989. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677489900217.

Khuller:1999:BPI

[KGW99] Samir Khuller, Manfred Göbel, and Jochen Walter. Bases for polynomial invariants of conjugates of permutation groups. Journal of Algorithms, 32(1):58–61, July 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910179.

Khuller:1997:P

[Khu97] Samir Khuller. Problems. Journal of Algorithms, 23(2):401–403, May 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497908476.

Khuller:1998:P

[Khu98] Samir Khuller. Problems. Journal of Algorithms, 28(1):192–195, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909324.

Kim:2005:DMM

[Kim05] Yoo-Ah Kim. Data migration to minimize the total completion time. Journal of Algorithms, 55(1):42–57, April 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001178.

Kingston:1988:NPG

[Kin88] Feffrey H. Kingston. A new proof of the Garsia–Wachs algorithm. Journal of Algorithms, 9(1):129–136, March 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900090.

Kloks:1995:TCB

[KK95] T. Kloks and D. Kratsch. Treewidth of chordal bipartite graphs. Journal of Algorithms, 19(2):266–281, September 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710371.

Kranakis:1997:DCA

[KK97] Evangelos Kranakis and Danny Krizanc. Distributed computing on anonymous hypercube networks. Journal of Algorithms, 23(1): 32–50, April 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908172.

Kleinberg:2001:WCO

[KK01] Jon Kleinberg and Amit Kumar. Wavelength conversion in optical networks. Journal of Algorithms, 38(1):25–50, January 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911374.

Kashyap:2006:ANU

[KK06] Srinivas Kashyap and Samir Khuller. Algorithms for nonuniform size data placement on parallel disks. Journal of Algorithms, 60(2):144–167, August 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001063.

Kaminski:1988:ARM

[KKB88] Michael Kaminski, David G. Kirkpatrick, and Nader H. Bshouty. Addition requirements for matrix and transposed matrix products. Journal of Algorithms, 9(3):354–364, September 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900260.

Kloks:1999:ABA

[KKM99] Ton Kloks, Dieter Kratsch, and Haiko Müller. Approximating the bandwidth for asteroidal triple-free graphs. Journal of Algorithms, 32(1):41–57, July 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749890997X.

Kranakis:2001:FTB

[KKP01] Evangelos Kranakis, Danny Krizanc, and Andrzej Pelc. Faulttolerant broadcasting in radio networks. Journal of Algorithms, 39(1):47–67, April 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911477.

Kloks:1998:MFC

[KKW98] T. Kloks, D. Kratsch, and C. K. Wong. Minimum fill-in on circle and circular-arc graphs. Journal of Algorithms, 28(2):272– 289, August 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909361.

Khuller:2006:GGB

[KKW06] Samir Khuller, Yoo-Ah Kim, and Yung-Chun (Justin) Wan. On generalized gossiping and broadcasting. Journal of Algorithms, 59(2):81–106, May 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000192.

Klee:1985:FST

[KL85] Victor Klee and Michael C. Laskowski. Finding the smallest triangles containing a given convex polygon. Journal of Algorithms, 6(3):359–375, September 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900057.

Kim:1991:CCE

[KL91] Young Man Kim and Ten-Hwang Lai. The complexity of congestion-1 embedding in a hypercube. Journal of Algorithms, 12(2):246–280, June 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190004I.

Karpinski:1993:ANZ

[KL93] M. Karpinski and M. Luby. Approximating the number of zeroes of a GF[2] polynomial. Journal of Algorithms, 14(2):280– 287, March 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667748371014X.

Kofler:1994:AMT

[KL94] E. T. Kofler and C. T. Leondes. Algorithmic modifications to the theory of evidential reasoning. Journal of Algorithms, 17(2):269– 279, September 1994. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677484710352.

Korsh:1997:GMP

[KL97] James Korsh and Seymour Lipschutz. Generating multiset permutations in constant time. Journal of Algorithms, 25(2):321–335, November 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908890.

Kahale:1998:GDR

[KL98] Nabil Kahale and Tom Leighton. Greedy dynamic routing on arrays. Journal of Algorithms, 29(2):390–410, November 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909580.

Korsh:2000:MPL

[KL00] James F. Korsh and Paul LaFollette. Multiset permutations and loopless generation of ordered trees with specified degree sequence. Journal of Algorithms, 34(2):309–336, February 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910593.

Klarner:1981:ADW

[Kla81] David A. Klarner. An algorithm to determine when certain sets have 0-density. Journal of Algorithms, 2(1):31–43, March 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677481900055.

Klawe:1992:SBM

[Kla92] Maria M. Klawe. Superlinear bounds for matrix searching problems. Journal of Algorithms, 13(1):55–78, March 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290005W.

Klarlund:1999:AOB

[Kla99] Nils Klarlund. An n log n algorithm for online BDD refinement. Journal of Algorithms, 32(2):133–154, August 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910131.

Klein:1993:PPR

[Kle93] P. N. Klein. Parallelism, preprocessing, and reachability: a hybrid algorithm for directed graphs. Journal of Algorithms, 14(3): 331–343, May 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710175.

Klingsberg:1980:CFL

[Kli80] Paul Klingsberg. A combinatorial family of labeled trees. Journal of Algorithms, 1(1):104–106, March 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677480900085.

Klingsberg:1982:GCC

[Kli82] Paul Klingsberg. A Gray code for compositions. Journal of Algorithms, 3(1):41–44, March 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677482900062.

Kunde:1988:SCU

[KLL88] Manfred Kunde, Michael A. Langston, and Jin-Ming Liu. On a special case of uniform processor scheduling. Journal of Algorithms, 9(2):287–296, June 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677488900430.

Karp:1989:MCA

[KLM89] Richard M. Karp, Michael Luby, and Neal Madras. Monte– Carlo approximation algorithms for enumeration problems. Journal of Algorithms, 10(3):429–448, September 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900382.

Kantor:1999:SSP

[KLM99] William M. Kantor, Eugene M. Luks, and Peter D. Mark. Sylow subgroups in parallel. Journal of Algorithms, 31(1):132– 195, April 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499909823.

Kindervater:1989:PCT

[KLS89] Gerard A. P. Kindervater, Jan Karel Lenstra, and David B. Shmoys. The parallel complexity of TSP heuristics. Journal of Algorithms, 10(2):249–270, June 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900151.

Kao:2001:EFM

[KLST01] Ming-Yang Kao, Tak-Wah Lam, Wing-Kin Sung, and Hing-Fung Ting. An even faster and more unifying algorithm for comparing trees via unbalanced bipartite matchings. Journal of Algorithms, 40(2):212–233, August 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911630.

Kannan:1996:DET

[KLW96] Sampath K. Kannan, Eugene L. Lawler, and Tandy J. Warnow. Determining the evolutionary tree using experiments. Journal of Algorithms, 21(1):26–50, July 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900358.

Kaufmann:1986:RTG

[KM86] Michael Kaufmann and Kurt Mehlhorn. Routing through a generalized switchbox. Journal of Algorithms, 7(4):510–531, December 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677486900167.

Kesselman:2003:LBA

[KM03] Alexander Kesselman and Yishay Mansour. Loss-bounded analysis for differentiated services. Journal of Algorithms, 46(1):79– 95, January 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002705.

Krumke:1999:IMC

[KMN^{+}99] S. O. Krumke, M. V. Marathe, H. Noltemeier, R. Ravi, S. S. Ravi, R. Sundaram, and H.-C Wirth. Improving minimum cost spanning trees by upgrading nodes. Journal of Algorithms, 33(1):92– 111, October 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910210.

Kashiwabara:1992:GMI

[KMNF92] Toshinobu Kashiwabara, Sumio Masuda, Kazuo Nakajima, and Toshio Fujisawa. Generation of maximum independent sets of a bipartite graph and maximum cliques of a circular-arc graph. Journal of Algorithms, 13(1):161–174, March 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677492900122.

Kao:1998:OCH

[KMSY98] Ming-Yang Kao, Yuan Ma, Michael Sipser, and Yiqun Yin. Optimal constructions of hybrid algorithms. Journal of Algorithms, 29(1):142–164, October 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909592.

Karchmer:1988:FPA

[KN88] Mauricio Karchmer and Joseph Naor. A fast parallel algorithm to color a graph with Δ colors. Journal of Algorithms, 9(1):83– 91, March 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900065.

Knott:1982:FBB

[Kno82] Gary D. Knott. Fixed-bucket binary storage trees. Journal of Algorithms, 3(3):276–287, September 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900256.

Knott:1984:DCC

[Kno84] Gary D. Knott. Direct-chaining with coalescing lists. Journal of Algorithms, 5(1):7–21, March 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677484900361.

Krivelevich:2001:ACM

[KNS01] Michael Krivelevich, Ram Nathaniel, and Benny Sudakov. Approximating coloring and maximum independent sets in 3-uniform hypergraphs. Journal of Algorithms, 41(1):99–113, October 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677401911733.

Knuth:1985:AOC

[Knu85a] Donald E. Knuth. An analysis of optimum caching. Journal of Algorithms, 6(2):181–199, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900379.

Knuth:1985:DHC

[Knu85b] Donald E. Knuth. Dynamic Huffman coding. Journal of Algorithms, 6(2):163–180, June 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900367.

Knuth:1996:EAS

[Knu96] Donald E. Knuth. An exact analysis of stable allocation. Journal of Algorithms, 20(2):431–442, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900206.

Kochol:2005:CCO

[Koc05] Martin Kochol. 3-coloring and 3-clique-ordering of locally connected graphs. Journal of Algorithms, 54(1):122–125, January 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404000896.

Kutten:2000:VPI

[KOPS00] Shay Kutten, Rafail Ostrovsky, and Boaz Patt-Shamir. The Las-Vegas processor identity problem (how and when to be unique). Journal of Algorithms, 37(2):468–494, November 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911106.

Kozen:1994:F

[Koz94] D. Kozen. Foreword. Journal of Algorithms, 17(1):1, July 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710248.

Kalyanasundaram:1993:OWM

[KP93] B. Kalyanasundaram and K. Pruhs. Online weighted matching. Journal of Algorithms, 14(3):478–488, May 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710266.

Kortsarz:1994:GSS

[KP94] G. Kortsarz and D. Peleg. Generating sparse 2-spanners. Journal of Algorithms, 17(2):222–236, September 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710327.

Kannan:1998:RAS

[KP98a] Sampath Kannan and Todd Proebsting. Register allocation in structured programs. Journal of Algorithms, 29(2):223–237, November 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909567.

Kutten:1998:FDC

[KP98b] Shay Kutten and David Peleg. Fast distributed construction of small k-dominating sets and applications. Journal of Algorithms, 28(1):40–66, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909294.

Kutten:1999:FLD

[KP99] Shay Kutten and David Peleg. Fault-local distributed mending. Journal of Algorithms, 30(1):144–165, January 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909725.

Kalyanasundaram:2001:EMM

[KP01] Bala Kalyanasundaram and Kirk R. Pruhs. Eliminating migration in multi-processor scheduling. Journal of Algorithms, 38(1): 2–24, January 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911283.

Kalyanasundaram:2003:MJC

[KP03] Bala Kalyanasundaram and Kirk R. Pruhs. Maximizing job completions online. Journal of Algorithms, 49(1):63–85, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000749.

Kamath:1998:RAC

[KPP98] Anil Kamath, Omri Palmon, and Serge Plotkin. Routing and admission control in general topology networks with Poisson arrivals. Journal of Algorithms, 27(2):236–258, May 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909238.

Korupolu:2000:ALS

[KPR00] Madhukar R. Korupolu, C. Greg Plaxton, and Rajmohan Rajaraman. Analysis of a local search heuristic for facility location problems. Journal of Algorithms, 37(1):146–188, October 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911003.

Korupolu:2001:PAH

[KPR01] Madhukar R. Korupolu, C. Greg Plaxton, and Rajmohan Rajaraman. Placement algorithms for hierarchical cooperative caching. Journal of Algorithms, 38(1):260–302, January 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911295.

Klein:1997:AAS

[KPRT97] Philip N. Klein, Serge A. Plotkin, Satish Rao, and Éva Tardos. Approximation algorithms for Steiner and directed multicuts. Journal of Algorithms, 22(2):241–269, February 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908330.

Karger:1996:BAA

[KPT96] David R. Karger, Steven J. Phillips, and Eric Torng. A better algorithm for an ancient scheduling problem. Journal of Algorithms, 20(2):400–430, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749690019X.

Koda:1993:GCI

[KR93] Y. Koda and F. Ruskey. A Gray code for the ideals of a forest poset. Journal of Algorithms, 15(2):324–340, September 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710448.

Kelsen:1995:FMT

[KR95a] P. Kelsen and V. Ramachandran. On finding minimal twoconnected subgraphs. Journal of Algorithms, 18(1):1–49, January 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710012.

Klein:1995:NBP

[KR95b] P. Klein and R. Ravi. A nearly best-possible approximation algorithm for node-weighted Steiner trees. Journal of Algorithms, 19(1):104–115, July 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677485710292.

Khuller:1996:IAA

[KR96] Samir Khuller and Balaji Raghavachari. Improved approximation algorithms for uniform connectivity problems. Journal of Algorithms, 21(2):434–450, September 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900528.

Kapoor:2000:ECB

[KR00] Ajai Kapoor and Romeo Rizzi. Edge-coloring bipartite graphs. Journal of Algorithms, 34(2):390–396, February 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910581.

Kesselman:2006:SPC

[KR06] Alex Kesselman and Adi Rosén. Scheduling policies for CIOQ switches. Journal of Algorithms, 60(1):60–83, July 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740400149X.

Krawczyk:1992:HPC

[Kra92] Hugo Krawczyk. How to predict congruential generators. Journal of Algorithms, 13(4):527–545, December 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290054G.

Krizanc:1996:TRT

[Kri96] Danny Krizanc. Time-randomness trade-offs in parallel computation. Journal of Algorithms, 20(1):1–19, January 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900012.

Krivelevich:1997:ASC

[Kri97] Michael Krivelevich. Approximate set covering in uniform hypergraphs. Journal of Algorithms, 25(1):118–143, October 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908725.

Kenyon:1998:BRW

[KRS98] Claire Kenyon, Yuval Rabani, and Alistair Sinclair. Biased random walks, Lyapunov functions, and stochastic analysis of best fit bin packing. Journal of Algorithms, 27(2):218–235, May 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497909196.

King:1994:FDM

[KRT94] V. King, S. Rao, and R. Tarjan. A faster deterministic maximum flow algorithm. Journal of Algorithms, 17(3):447–474, November 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710443.

Karloff:1987:EPA

[KS87] Howard J. Karloff and David B. Shmoys. Efficient parallel algorithms for edge coloring problems. Journal of Algorithms, 8(1): 39–52, March 1987. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677487900265.

Kaib:1996:GGR

[KS96] Michael Kaib and Claus P. Schnorr. The generalized Gauss reduction algorithm. Journal of Algorithms, 21(3):565–578, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900590.

Klein:1997:RPA

[KS97] Philip N. Klein and Sairam Subramanian. A randomized parallel algorithm for single-source shortest paths. Journal of Algorithms, 25(2):205–220, November 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908889.

Kolliopoulos:1998:FRV

[KS98] Stavros G. Kolliopoulos and Clifford Stein. Finding realvalued single-source shortest paths in o(n^{3}) expected time. Journal of Algorithms, 28(1):125–141, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909373.

Knessl:2002:LLH

[KS02] Charles Knessl and Wojciech Szpankowski. Limit laws for the height in PATRICIA tries. Journal of Algorithms, 44(1):63–97, July 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402002122.

Krivelevich:2003:ACU

[KS03] Michael Krivelevich and Benny Sudakov. Approximate coloring of uniform hypergraphs. Journal of Algorithms, 49(1):2–12, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000774.

Kolman:2006:IBU

[KS06] Petr Kolman and Christian Scheideler. Improved bounds for the unsplittable flow problem. Journal of Algorithms, 61(1):20–44, September 2006. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001397.

Karp:1980:LET

[KT80] Richard M. Karp and Robert Endre Tarjan. Linear expectedtime algorithms for connectivity problems. Journal of Algorithms, 1(4):374–393, December 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677480900176.

Ko:1986:NQN

[KT86] Ker-I Ko and Shia-Chung Teng. On the number of queries necessary to identify a permutation. Journal of Algorithms, 7(4):449– 462, December 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900131.

Kantor:1988:PTV

[KT88] W. M. Kantor and D. E. Taylor. Polynomial-time versions of Sylow’s theorem. Journal of Algorithms, 9(1):1–17, March 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900028.

Khuller:1993:AAG

[KT93] S. Khuller and R. Thurimella. Approximation algorithms for graph augmentation. Journal of Algorithms, 14(2):214–225, March 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710102.

Kratochvil:2002:CBC

[KT02] Jan Kratochvíl and Zsolt Tuza. On the complexity of bicoloring clique hypergraphs of graphs. Journal of Algorithms, 45(1):40– 54, October 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402002213.

Kucera:1991:GCB

[Kuc91] Ludek Kucera. The greedy coloring is a bad probabilistic algorithm. Journal of Algorithms, 12(4):674–684, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677491900406.

Kulkarni:1990:GRC

[Kul90] V. G. Kulkarni. Generating random combinatorial objects. Journal of Algorithms, 11(2):185–207, June 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090002V.

Kalvin:1983:GAT

[KV83] Alan D. Kalvin and Yaakov L. Varol. On the generation of all topological sortings. Journal of Algorithms, 4(2):150–162, June 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677483900421.

Khuller:1994:PDP

[KVY94] S. Khuller, U. Vishkin, and N. Young. A primal–dual parallel approximation technique applied to weighted set and vertex covers. Journal of Algorithms, 17(2):280–289, September 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710364.

Lagarias:1980:WCC

[Lag80] J. C. Lagarias. Worst-case complexity bounds for algorithms in the theory of integral quadratic forms. Journal of Algorithms, 1(2):142–186, June 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677480900218.

Lagergren:1996:EPA

[Lag96] Jens Lagergren. Efficient parallel algorithms for graphs of bounded tree-width. Journal of Algorithms, 20(1):20–44, January 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900024.

Larson:1984:AHC

[Lar84] Per-Åke Larson. Analysis of hashing with chaining in the prime area. Journal of Algorithms, 5(1):36–47, March 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900385.

Larmore:1987:SAC

[Lar87] Lawrence L. Larmore. A subquadratic algorithm for constructing approximately optimal binary search trees. Journal of Algorithms, 8(4):579–591, December 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900526.

Lindenbaum:1992:PSG

[LB92] Michael Lindenbaum and Alfred Bruckstein. Parallel strategies for geometric probing. Journal of Algorithms, 13(2):320–349, June 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677492900225.

Li:1992:HAL

[LC92] Keoin Li and Kam Hoi Cheng. Heuristic algorithms for on-line packing in three dimensions. Journal of Algorithms, 13(4):589– 605, December 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290058K.

Lengauer:1984:SIS

[Len84] T. Lengauer. On the solution of inequality systems relevant to IC-layout. Journal of Algorithms, 5(3):408–421, September 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677484900208.

Lengauer:1987:EAF

[Len87] T. Lengauer. Efficient algorithms for finding minimum spanning forests of hierarchically defined graphs. Journal of Algorithms, 8(2):260–284, June 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677487900423.

Leung:1984:FAG

[Leu84] Joseph Y.-T Leung. Fast algorithms for generating all maximal independent sets of interval, circular-arc and chordal graphs. Journal of Algorithms, 5(1):22–35, March 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900373.

Leven:1983:NCF

[LG83] Daniel Leven and Zvi Galil. NP completeness of finding the chromatic index of regular graphs. Journal of Algorithms, 4(1):35– 44, March 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900329.

Liberatore:2004:CAC

[Lib04] Vincenzo Liberatore. Circular arrangements and cyclic broadcast scheduling. Journal of Algorithms, 51(2):185–215, May 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001627.

Lieberherr:1982:AEP

[Lie82] Karl J. Lieberherr. Algorithmic extremal problems in combinatorial optimization. Journal of Algorithms, 3(3):225–244, September 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677482900220.

Liu:2002:BSO

[Liu02] Pangfeng Liu. Broadcast scheduling optimization for heterogeneous cluster systems. Journal of Algorithms, 42(1):135–152, January 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677401912040.

Levcopoulos:1998:QGT

[LK98] Christos Levcopoulos and Drago Krznaric. Quasi-greedy triangulations approximating the minimum weight triangulation. Journal of Algorithms, 27(2):303–338, May 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909184.

Lee:2005:EPE

[LKPC05] Mun-Kyu Lee, Yoonjeong Kim, Kunsoo Park, and Yookun Cho. Efficient parallel exponentiation in GF(qn) using normal basis representations. Journal of Algorithms, 54(2):205–221, February 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404001051.

Levy:1988:CAF

[LL88] Hanoch Levy and David W. Low. A contraction algorithm for finding small cycle cutsets. Journal of Algorithms, 9(4):470–493, December 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900132.

LaMarca:1999:ICP

[LL99] Anthony LaMarca and Richard E. Ladner. The influence of caches on the performance of sorting. Journal of Algorithms, 31(1):66– 104, April 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909853.

Lewandowski:1986:APG

[LLL86] J. L. Lewandowski, C. L. Liu, and J. W. S. Liu. An algorithmic proof of a generalization of the Birkhoff–von Neumann theorem. Journal of Algorithms, 7(3):323–330, September 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900246.

Lloyd:1983:AJP

[Llo83] Errol L. Lloyd. An O(n log m) algorithm for the Josephus Problem. Journal of Algorithms, 4(3):262–270, September 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900251.

Liao:2003:CFP

[LLY03] Chien-Chih Liao, Hsueh-I Lu, and Hsu-Chun Yen. Compact floor-planning via orderly spanning trees. Journal of Algorithms, 48(2):441–451, September 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000579.

Larose:2005:PTA

[LLZ05] Benoit Larose, Cynthia Loten, and László Zádori. A polynomialtime algorithm for near-unanimity graphs. Journal of Algorithms, 55(2):177–191, May 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740400077X.

Li:1990:PBG

[LM90] Liwu Li and T. A. Marsland. Probability-based game tree pruning. Journal of Algorithms, 11(1):27–43, March 1990. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090027C.

Leighton:1994:RRS

[LMRR94] F. T. Leighton, B. M. Maggs, A. G. Ranade, and S. B. Rao. Randomized routing and sorting on fixed-connection networks. Journal of Algorithms, 17(1):157–205, July 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710303.

Lupton:1998:DCS

[LMY98] Robert Lupton, F. Miller Maley, and Neal Young. Data collection for the Sloan Digital Sky Survey — a network-flow heuristic. Journal of Algorithms, 27(2):339–356, May 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909226.

Loebl:1997:LUS

[LN97] Martin Loebl and Jaroslav Nesetril. Linearity and unprovability of set union problem strategies. Journal of Algorithms, 23(2): 207–220, May 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908159.

Liben-Nowell:2002:GSI

[LN02] David Liben-Nowell. Gossip is synteny: Incomplete gossip and the syntenic distance between genomes. Journal of Algorithms, 43(2): 264–283, May 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402000068.

Lam:2004:PGE

[LNT04] Tak-Wah Lam, Tsuen-Wan Johnny Ngan, and Kar-Keung To. Performance guarantee for EDF under overload. Journal of Algorithms, 52(2):193–206, August 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001640.

Lagarias:1987:CAM

[LO87] J. C. Lagarias and A. M. Odlyzko. Computing π(x): an analytic method. Journal of Algorithms, 8(2):173–191, June 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748790037X.

Lingas:2001:EMC

[LOÖ01] Andrzej Lingas, Hans Olsson, and Anna Östlin. Efficient merging and construction of evolutionary trees. Journal of Algorithms, 41(1):41–51, October 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740191168X.

Lotti:1992:FSL

[Lot92] Grazia Lotti. Fast solution of linear systems with polynomial coefficients over the ring of integers. Journal of Algorithms, 13(4):564– 576, December 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290056I.

Loveland:1987:FCS

[Lov87] Donald W. Loveland. Finding critical sets. Journal of Algorithms, 8(3):362–371, September 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677487900162.

Lozinskii:1983:APE

[Loz83] Eliezer L. Lozinskii. An algorithm for parallel evaluation of functions. Journal of Algorithms, 4(3):271–281, September 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900263.

Lipski:1980:FCU

[LP80] Witold Lipski, Jr. and Franco P. Preparata. Finding the contour of a union of iso-oriented rectangies. Journal of Algorithms, 1(3):235–246, September 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677480900115.

See erratum [Ano81c].

Lipski:1981:FAT

[LP81a] Witold Lipski, Jr. and Christos H. Papadimitriou. A fast algorithm for testing for safety and detecting deadlocks in locked transaction systems. Journal of Algorithms, 2(3):211–226, September 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900237.

Lipski:1981:SRC

[LP81b] Witold Lipski, Jr. and Franco P. Preparata. Segments, rectangles, contours. Journal of Algorithms, 2(1):63–76, March 1981. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900080.

Lee:1982:IAR

[LP82] D. T. Lee and F. P. Preparata. An improved algorithm for the rectangle enclosure problem. Journal of Algorithms, 3(3):218–224, September 1982. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677482900219.

Levcopoulos:1993:AH

[LP93] C. Levcopoulos and O. Petersson. Adaptive Heapsort. Journal of Algorithms, 14(3):395–413, May 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677483710217.

For other analyses of Heapsort, see [SS93, BFF96].

Larmore:1998:OAT

[LP98] Lawrence L. Larmore and Teresa M. Przytycka. The optimal alphabetic tree problem revisited. Journal of Algorithms, 28(1): 1–20, July 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909348.

Lu:1998:AML

[LR98] Hsueh-I Lu and R. Ravi. Approximating maximum leaf spanning trees in almost linear time. Journal of Algorithms, 29(1):132– 141, October 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909440.

Lopez:1999:APA

[LR99] Mario A. Lopez and Shlomo Reisner. Algorithms for polyhedral approximation of multidimensional ellipsoids. Journal of Algorithms, 33(1):140–165, October 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910313.

Lai:1983:NLS

[LS83] Ten-Hwang Lai and Sartaj Sahni. Nearly on-line scheduling of multiprocessor systems with memories. Journal of Algorithms, 4(4):353–362, December 1983. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677483900160.

Linial:1985:SOS

[LS85] N. Linial and M. Saks. Searching ordered structures. Journal of Algorithms, 6(1):86–103, March 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900203.

Lai:1987:RCG

[LS87a] Ten-Hwang Lai and Alan Sprague. On the routability of a convex grid. Journal of Algorithms, 8(3):372–384, September 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900174.

Leven:1987:ESM

[LS87b] Daniel Leven and Micha Sharir. An efficient and simple motion planning algorithm for a ladder amidst polygonal barriers. Journal of Algorithms, 8(2):192–215, June 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900381.

Leiserson:1988:MIL

[LS88] Charles E. Leiserson and James B. Saxe. A mixed-integer linear programming problem which is efficiently solvable. Journal of Algorithms, 9(1):114–128, March 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900089.

Larmore:1991:LDP

[LS91] Lawrence L. Larmore and Baruch Schieber. On-line dynamic programming with applications to the prediction of RNA secondary structure. Journal of Algorithms, 12(3):490–515, September 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190016R.

Luczak:1997:PRA

[LS97] Tomasz Luczak and Edyta Szymańska. A parallel randomized algorithm for finding a maximal independent set in a linear hypergraph. Journal of Algorithms, 25(2):311–320, November 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497908841.

Lustig:1999:AHP

[LS99] Aviv Lustig and Oded Shmueli. Acyclic hypergraph projections. Journal of Algorithms, 30(2):400–422, February 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909658.

Lucas:1987:RGB

[Luc87] Joan M. Lucas. The rotation graph of binary trees is Hamiltonian. Journal of Algorithms, 8(4):503–535, December 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900484.

Lueker:1998:ACA

[Lue98] George S. Lueker. Average-case analysis of off-line and online knapsack problems. Journal of Algorithms, 29(2):277–305, November 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909543.

Landau:1989:FPS

[LV89] Gad M. Landau and Uzi Vishkin. Fast parallel and serial approximate string matching. Journal of Algorithms, 10(2):157– 169, June 1989. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677489900102.

Lucas:1993:RGB

[LVR93] J. M. Lucas, D. R. Vanbaronaigien, and F. Ruskey. On rotations and the generation of binary trees. Journal of Algorithms, 15(3):343–366, November 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667748371045X.

Lenstra:1996:CST

[LVV96] Jan Karel Lenstra, Marinus Veldhorst, and Bart Veltman. The complexity of scheduling trees with communication delays. Journal of Algorithms, 20(1):157–173, January 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900073.

Lee:1981:FIR

[LW81] D. T. Lee and C. K. Wong. Finding intersection of rectangles by range search. Journal of Algorithms, 2(4):337–347, December 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667748190033X.

Lai:1995:APR

[LW95] T. H. Lai and S. S. Wei. Algorithms for page retrieval and Hamiltonian paths on forward-convex line graphs. Journal of Algorithms, 18(2):358–375, March 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710140.

Li:1999:TRT

[LZ99] Ming Li and Louxin Zhang. Twist-rotation transformations of binary trees and arithmetic expressions. Journal of Algorithms, 32(2):155–166, August 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910106.

Landau:2001:CSA

[LZU01] Gad M. Landau and Michal Ziv-Ukelson. On the common substring alignment problem. Journal of Algorithms, 41(2):338–359, November 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911915.

Maimon:1988:ALM

[Mai88] Oded Maimon. An algorithm for the Lorenz measure in locational decisions on trees. Journal of Algorithms, 9(4):583–596, December 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900181.

Makino:2002:LTA

[Mak02] Kazuhisa Makino. A linear time algorithm for recognizing regular Boolean functions. Journal of Algorithms, 43(2):155–176, May 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402000032.

Malitz:1994:GGP

[Mal94a] S. M. Malitz. Genus g graphs have pagenumber O(\sqrt{g}). Journal of Algorithms, 17(1):85–109, July 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710285.

Malitz:1994:GEP

[Mal94b] S. M. Malitz. Graphs with E edges have pagenumber O(\sqrt{E}). Journal of Algorithms, 17(1):71–84, July 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710273.

Mastrolilli:2006:LTA

[Mas06] Monaldo Mastrolilli. A linear time approximation scheme for the single machine scheduling problem with controllable processing times. Journal of Algorithms, 59(1):37–52, April 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000210.

Matousek:1993:LOQ

[Mat93] J. Matousek. Linear optimization queries. Journal of Algorithms, 14(3):432–448, May 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677483710230.

Matousek:1996:DCG

[Mat96] Jiří Matousek. Derandomization in computational geometry. Journal of Algorithms, 20(3):545–580, May 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900279.

McKay:1998:IFE

[McK98] Brendan D. McKay. Isomorph-free exhaustive generation. Journal of Algorithms, 26(2):306–324, February 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908981.

Megiddo:1985:PTL

[Meg85] Nimrod Megiddo. Partitioning with two lines in the plane. Journal of Algorithms, 6(3):430–433, September 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900112.

Melville:1981:TST

[Mel81] R. Melville. A time/space tradeoff for in-place array permutation. Journal of Algorithms, 2(2):139–143, June 1981. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900158.

Meyer:2003:ACC

[Mey03] Ulrich Meyer. Average-case complexity of single-source shortestpaths algorithms: lower and upper bounds. Journal of Algorithms, 48(1):91–134, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000464.

McDiarmid:1996:LDQ

[MH96] C. J. H. McDiarmid and R. B. Hayward. Large deviations for Quicksort. Journal of Algorithms, 21(3):476–507, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900553. For other Quicksort bounds,

see [ES95, FJ02, NR02].

Mignotte:1982:IAN

[Mig82] Maurice Mignotte. Identification of algebraic numbers. Journal of Algorithms, 3(3):197–204, September 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900190.

Minsker:1989:THR

[Min89] Steven Minsker. The Towers of Hanoi rainbow problem: Coloring the rings. Journal of Algorithms, 10(1):1–19, March 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900205.

Mirolo:1998:CMG

[Mir98] Claudio Mirolo. Convex minimization on a grid and applications. Journal of Algorithms, 26(2):209–237, February 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909081.

Main:1984:AFA

[ML84] Michael G. Main and Richard J. Lorentz. An O(n log n) algorithm for finding all repetitions in a string. Journal of Algorithms, 5(3):422–432, September 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748490021X.

Milidiu:1999:BCL

[MLP99] Ruy Luiz Milidiú, Eduardo Sany Laber, and Artur Alves Pessoa. Bounding the compression loss of the FGK algorithm. Journal of Algorithms, 32(2):195–211, August 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749991012X.

Manasse:1990:CAS

[MMS90] Mark S. Manasse, Lyle A. McGeoch, and Daniel D. Sleator. Competitive algorithms for server problems. Journal of Algorithms, 11(2):208–230, June 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749090003W.

Mitra:1998:ECR

[MN98] Pinaki Mitra and Subhas C. Nandy. Efficient computation of rectilinear geodesic Voronoi neighbor in the presence of obstacles. Journal of Algorithms, 28(2):315–338, August 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909397.

Makinen:2005:TIS

[MNU05] Veli Mäkinen, Gonzalo Navarro, and Esko Ukkonen. Transposition invariant string matching. Journal of Algorithms, 56(2):124– 153, August 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001427.

Miller:1985:NCM

[MO85] Z. Miller and J. B. Orlin. NP-completeness for minimizing maximum edge length in grid embeddings. Journal of Algorithms, 6(1): 10–16, March 1985. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677485900161.

Mohar:1993:PPL

[Moh93] B. Mohar. Projective planarity in linear time. Journal of Algorithms, 15(3):482–502, November 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710503.

Monier:1980:CSM

[Mon80] Louis Monier. Combinatorial solutions of multidimensional divide-and-conquer recurrences. Journal of Algorithms, 1(1):60– 74, March 1980. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748090005X.

Marti-Oliet:2007:SSS

[MOPV07] Narciso Martí-Oliet, Miguel Palomino, and Alberto Verdejo. Strategies and simulations in a semantic framework. Journal of Algorithms, 62(3–4):95–116, July/October 2007. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677407000296.

Moret:2001:F

[Mor01] Bernard M. E. Moret. Foreword. Journal of Algorithms, 38(1): 1, January 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911398.

Motwani:1998:RMD

[Mot98] Rajeev Motwani. Realization of matrices and directed graphs. Journal of Algorithms, 27(1):61–74, April 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909093.

Moran:1981:CIR

[MP81] Shlomo Moran and Yehoshua Perl. The complexity of identifying redundant and essential elements. Journal of Algorithms, 2(1): 22–30, March 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900043.

Minieka:1983:FCT

[MP83] Edward Minieka and Niranjani H. Patel. On finding the core of a tree with a specified length. Journal of Algorithms, 4(4):345– 352, December 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900159.

Mahmoud:1989:ASS

[MP89] Hosam M. Mahmoud and Boris Pittel. Analysis of the space of search trees under the random insertion algorithm. Journal of Algorithms, 10(1):52–75, March 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900230.

Mansour:1993:GPS

[MP93] Y. Mansour and B. Pattshamir. Greedy packet scheduling on shortest paths. Journal of Algorithms, 14(3):449–465, May 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710242.

McDiarmid:1989:BHF

[MR89] C. J. H. McDiarmid and B. A. Reed. Building heaps fast. Journal of Algorithms, 10(3):352–365, September 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900333.

Munro:1992:SMD

[MR92] J. Ian Munro and Venkatesh Raman. Sorting with minimum data movement. Journal of Algorithms, 13(3):374–393, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290045E.

Mahajan:1999:PAG

[MR99] Meena Mahajan and Venkatesh Raman. Parameterizing above guaranteed values: MaxSat and MaxCut. Journal of Algorithms, 31(2):335–354, May 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677498909968.

Maratea:2008:LBT

[MRFL08] Marco Maratea, Francesco Ricca, Wolfgang Faber, and Nicola Leone. Look-back techniques and heuristics in DLV: Implementation, evaluation, and comparison to QBF solvers. Journal of Algorithms, 63(1–3):70–89, January/July 2008. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677408000072.

Munro:2001:SES

[MRR01] J. Ian Munro, Venkatesh Raman, and S. Srinivasa Rao. Space efficient suffix trees. Journal of Algorithms, 39(2):205–222, May 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400911519.

Marathe:1998:BND

[MRS^{+}98] Madhav V. Marathe, R. Ravi, Ravi Sundaram, S. S. Ravi, Daniel J. Rosenkrantz, and Harry B. Hunt III. Bicriteria network design problems. Journal of Algorithms, 28(1):142–171, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498909300.

Morgan:1980:LAC

[MS80] Christine A. Morgan and Peter J. Slater. A linear algorithm for a core of a tree. Journal of Algorithms, 1(3):247–258, September 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677480900127.

Mohammed:1987:IBI

[MS87] John L. Mohammed and Carlos S. Subi. An improved blockinterchange algorithm. Journal of Algorithms, 8(1):113–121, March 1987. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677487900319.

Mansour:1989:FEC

[MS89] Yishay Mansour and Baruch Schieber. Finding the edge connectivity of directed graphs. Journal of Algorithms, 10(1):76– 85, March 1989. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677489900242.

Mansour:1990:SRP

[MS90a] Yishay Mansour and Leonard Schulman. Sorting on a ring of processors. Journal of Algorithms, 11(4):622–630, December 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677490900124.

Morton:1990:CVP

[MS90b] H. R. Morton and H. B. Short. Calculating the 2-variable polynomial for knots presented as closed braids. Journal of Algorithms, 11(1):117–131, March 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090033B.

Mount:1990:PCP

[MS90c] David M. Mount and Ruth Silverman. Packing and covering the plane with translates of a convex polygon. Journal of Algorithms, 11(4):564–580, December 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090010C.

Ma:1994:AUS

[MS94a] T. H. Ma and J. Spinrad. An O(n^{2}) algorithm for undirected split decomposition. Journal of Algorithms, 16(1):145–160, January 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710078.

Ma:1994:CSC

[MS94b] T. H. Ma and J. P. Spinrad. On the 2-chain subgraph cover and related problems. Journal of Algorithms, 17(2):251–268, September 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710340.

Martel:1994:CCW

[MS94c] C. Martel and R. Subramonian. On the complexity of certified write-all algorithms. Journal of Algorithms, 16(3):361–387, May 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677484710170.

Manne:1995:OPS

[MS95] F. Manne and T. Sorevik. Optimal partitioning of sequences. Journal of Algorithms, 19(2):235–249, September 1995. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710358.

MacKenzie:1998:UET

[MS98] Philip D. MacKenzie and Quentin F. Stout. Ultrafast expected time parallel algorithms. Journal of Algorithms, 26(1):1–33, January 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497908920.

Merlini:2002:FH

[MS02a] Donatella Merlini and Renzo Sprugnoli. Fountains and histograms. Journal of Algorithms, 44(1):159–176, July 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002110.

Meyer:2002:OGT

[MS02b] Ulrich Meyer and Jop F. Sibeyn. Oblivious gossiping on tori. Journal of Algorithms, 42(1):1–19, January 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401912003.

Meyer:2003:SPS

[MS03] U. Meyer and P. Sanders. Δ-stepping: a parallelizable shortest path algorithm. Journal of Algorithms, 49(1):114–152, October 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000762.

Muthukrishnan:2005:AAA

[MS05] S. Muthukrishnan and Torsten Suel. Approximation algorithms for array partitioning problems. Journal of Algorithms, 54(1):85– 104, January 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001676.

Marques-Silva:2008:MCB

[MS08] Joao Marques-Silva. Model checking with Boolean Satisfiability. Journal of Algorithms, 63(1–3):3–16, January/July 2008. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677408000035.

Mehta:2006:PPP

[MSV06] Aranyak Mehta, Scott Shenker, and Vijay V. Vazirani. Posted price profit maximization for multicast by approximating fixed points. Journal of Algorithms, 58(2):150–164, February 2006. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001440.

Matousek:1991:AFT

[MT91] Jirí Matousek and Robin Thomas. Algorithms finding treedecompositions of graphs. Journal of Algorithms, 12(1):1–22, March 1991. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749190020Y.

Moran:1997:LBW

[MT97] Shlomo Moran and Gadi Taubenfeld. A lower bound on wait-free counting. Journal of Algorithms, 24(1):1–19, July 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908378.

Miller:1999:OCU

[MTT99] Gary L. Miller, Dafna Talmor, and Shang-Hua Teng. Optimal coarsening of unstructured meshes. Journal of Algorithms, 31(1): 29–65, April 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909907.

Makino:2001:MER

[MUI01] Kazuhisa Makino, Yushi Uno, and Toshihide Ibaraki. On minimum edge ranking spanning trees. Journal of Algorithms, 38(2):411–437, February 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740091143X.

Muthukrishnan:2000:SOP

[Mut00] S. Muthukrishnan. Simple optimal parallel multiple pattern matching. Journal of Algorithms, 34(1):1–13, January 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910155.

Matias:1991:PHI

[MV91] Yossi Matias and Uzi Vishkin. On parallel hashing and integer sorting. Journal of Algorithms, 12(4):573–606, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190034V.

McKay:1990:UGR

[MW90] Brendan D. McKay and Nicholas C. Wormald. Uniform generation of random regular graphs of moderate degree. Journal of Algorithms, 11(1):52–67, March 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090029E.

Matousek:1992:GSC

[MW92] Jirí Matousek and Emo Welzl. Good splitters for counting points in triangles. Journal of Algorithms, 13(2):307–319, June 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677492900214.

Moses:1994:CTR

[MW94] Y. Moses and O. Waarts. Coordinated traversal: (t + 1)- round Byzantine agreement in polynomial time. Journal of Algorithms, 17(1):110–156, July 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710297.

Megiddo:1986:RAW

[MZ86] Nimrod Megiddo and Eitan Zemel. An O(n log n) randomizing algorithm for the weighted Euclidean 1-center problem. Journal of Algorithms, 7(3):358–368, September 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900271.

Nagamochi:2005:AML

[Nag05] Hiroshi Nagamochi. A 4/3-approximation for the minimum 2-local-vertex-connectivity augmentation in a connected graph. Journal of Algorithms, 56(2):77–95, August 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740500026X.

Naor:2005:FCP

[Nao05] Moni Naor. On fairness in the carpool problem. Journal of Algorithms, 55(1):93–98, April 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000781.

Nandy:2000:SZP

[NBHB00] Subhas C. Nandy, Bhargab B. Bhattacharya, and Antonio Hernández-Barrera. Safety zone problem. Journal of Algorithms, 37(2):538–569, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910927.

Nalon:2007:CRN

[ND07] Cláudia Nalon and Clare Dixon. Clausal resolution for normal modal logics. Journal of Algorithms, 62(3–4):117–134, July/ October 2007. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677407000302.

Nakajima:1982:CRS

[NH82] Kazuo Nakajima and S. Louis Hakimi. Complexity results for scheduling tasks with discrete starting times. Journal of Algorithms, 3(4):344–361, December 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748290030X.

Nagamochi:2001:EAS

[NH01] Hiroshi Nagamochi and Toru Hasunuma. An efficient NC algorithm for a sparse k-edge-connectivity certificate. Journal of Algorithms, 38(2):354–373, February 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911441.

Nagamochi:1999:AEC

[NI99] Hiroshi Nagamochi and Toshihide Ibaraki. Augmenting edgeconnectivity over the entire range in Õ(nm) time. Journal of Algorithms, 30(2):253–301, February 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667749890983X.

Nichitiu:2001:ALE

[NMR01] Codrin Nichitiu, Jacques Mazoyer, and Eric Rémila. Algorithms for leader election by cellular automata. Journal of Algorithms, 41(2):302–329, November 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911757.

Naor:1990:ERG

[NN90] Joseph Naor and Mark B. Novick. An efficient reconstruction of a graph from its line graph in parallel. Journal of Algorithms, 11(1):132–143, March 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749090034C.

Nakayama:1985:LBC

[NNS85] Hiroyuki Nakayama, Takao Nishizeki, and Nobuji Saito. Lower bounds for combinatorial problems on graphs. Journal of Algorithms, 6(3):393–399, September 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677485900082.

Noshita:1985:TEC

[Nos85] Kohei Noshita. A theorem on the expected complexity of Dijkstra’s shortest path algorithm. Journal of Algorithms, 6(3):400– 408, September 1985. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677485900094.

Nikolopoulos:2004:PAC

[NP04] Stavros D. Nikolopoulos and Leonidas Palios. Parallel algorithms for P _{4}-comparability graphs. Journal of Algorithms, 51(1):77– 104, April 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403001895.

Norton:1992:USA

[NPT92] Carolyn Haibt Norton, Serge A. Plotkin, and Éva Tardos. Using separation algorithms in fixed dimension. Journal of Algorithms, 13(1):79–98, March 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749290006X.

Niedermeier:2000:NUB

[NR00] Rolf Niedermeier and Peter Rossmanith. New upper bounds for maximum satisfiability. Journal of Algorithms, 36(1):63–88, July 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677400910757.

Neininger:2002:RCQ

[NR02] Ralph Neininger and Ludger Rüschendorf. Rates of convergence for Quicksort. Journal of Algorithms, 44(1):52–62, July 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402002067. For other Quicksort bounds,

see [ES95, MH96, FJ02].

Niedermeier:2003:EFP

[NR03] Rolf Niedermeier and Peter Rossmanith. On efficient fixedparameter algorithms for weighted vertex cover. Journal of Algorithms, 47(2):63–77, July 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000051.

Narayanan:1996:AAM

[NRP96] H. Narayanan, Subir Roy, and Sachin Patkar. Approximation algorithms for min-k-overlap problems using the principal lattice of partitions approach. Journal of Algorithms, 21(2):306–330, September 1996. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496900474.

Nishimura:2002:GPL

[NRT02] Naomi Nishimura, Prabhakar Ragde, and Dimitrios M. Thilikos. On graph powers for leaf-labeled trees. Journal of Algorithms, 42(1):69–108, January 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677401911952.

Neiger:1990:AIF

[NT90] Gil Neiger and Sam Toueg. Automatically increasing the faulttolerance of distributed algorithms. Journal of Algorithms, 11(3):374–419, September 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749090019B.

Nykanen:2002:EPL

[NU02] Matti Nykänen and Esko Ukkonen. The exact path length problem. Journal of Algorithms, 42(1):41–53, January 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401912015.

Nedev:2000:PTA

[NW00] Zhivko P. Nedev and Peter T. Wood. A polynomial-time algorithm for finding regular simple paths in outerplanar graphs. Journal of Algorithms, 35(2):235–259, May 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910726.

ORourke:1986:OAF

[OAMB86] Joseph O’Rourke, Alok Aggarwal, Sanjeev Maddila, and Michael Baldwin. An optimal algorithm for finding minimal enclosing triangles. Journal of Algorithms, 7(2):258–269, June 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900076.

Oldham:2001:CAA

[Old01] Jeffrey D. Oldham. Combinatorial approximation algorithms for generalized flow problems. Journal of Algorithms, 38(1):135– 169, January 2001. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911301.

Olivos:1981:VAC

[Oli81] Jorge Olivos. On vectorial addition chains. Journal of Algorithms, 2(1):13–21, March 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900031.

Osorio:2009:P

[OO09] Mauricio Osorio and Ivan Olmos. Preface. Journal of Algorithms, 64(1):1–2, January 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677409000030.

Olariu:1988:CPH

[OTZ88] S. Olariu, S. Toida, and M. Zubair. On a conjecture by plaisted and hong. Journal of Algorithms, 9(4):597–598, December 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900193.

Overmars:1996:RSP

[OvdS96] Mark H. Overmars and Frank A. van der Stappen. Range searching and point location among fat objects. Journal of Algorithms, 21(3):629–656, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900632.

Overmars:1981:DOD

[Ove81] Mark H. Overmars. Dynamization of order decomposable set problems. Journal of Algorithms, 2(3):245–260, September 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677481900250.

Overmars:1988:EDS

[Ove88] Mark H. Overmars. Efficient data structures for range searching on a grid. Journal of Algorithms, 9(2):254–275, June 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900417.

Overmars:1988:RV

[OW88] Mark H. Overmars and Derick Wood. On rectangular visibility. Journal of Algorithms, 9(3):372–390, September 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900284.

ODunlaing:1985:RMP

[ÓY85] Colm Ó’Dúnlaing and Chee K. Yap. A “retraction” method for planning the motion of a disc. Journal of Algorithms, 6(1):104– 111, March 1985. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677485900215.

Pachl:1987:LBP

[Pac87] Jan K. Pachl. A lower bound for probabilistic distributed algorithms. Journal of Algorithms, 8(1):53–65, March 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900277.

Palios:1997:CMN

[Pal97] Leonidas Palios. Connecting the maximum number of nodes in the grid to the boundary with nonintersecting line segments. Journal of Algorithms, 22(1):57–92, January 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485707985.

Pan:1980:CVR

[Pan80] V. Ya Pan. Convolution of vectors over the real field of constants by evaluation-interpolation algorithms. Journal of Algorithms, 1(4):297–300, December 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677480900140.

Pan:1981:BCA

[Pan81a] V. Pan. The bit-complexity of arithmetic algorithms. Journal of Algorithms, 2(2):144–163, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667748190016X.

Pan:1981:UAA

[Pan81b] V. Pan. A unified approach to the analysis of bilinear algorithms. Journal of Algorithms, 2(3):301–310, September 1981. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900286.

Pan:1983:ALC

[Pan83] V. Ya Pan. The additive and logical complexities of linear and bilinear arithmetic algorithms. Journal of Algorithms, 4(1):1– 34, March 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900317.

Provan:1999:TCA

[PB99] J. Scott Provan and Roger C. Burk. Two-connected augmentation problems in planar graphs. Journal of Algorithms, 32(2): 87–107, August 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910052.

Przytycka:1991:PAP

[PC91] T. Przytycka and D. G. Corneil. Parallel algorithms for parity graphs. Journal of Algorithms, 12(1):96–109, March 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190025T.

Pinto:2009:SAC

[PCBC^{+}09] David Pinto, Jorge Civera, Alberto Barrón-Cedeño, Alfons Juan, and Paolo Rosso. A statistical approach to crosslingual natural language tasks. Journal of Algorithms, 64(1):51–60, January 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740900008X.

Pearl:1981:SEL

[Pea81] Judea Pearl. A space-efficient on-line method of computing quantile estimates. Journal of Algorithms, 2(2):164–177, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677481900171.

Pellegrini:1996:RHS

[Pel96] Marco Pellegrini. Repetitive hidden surface removal for polyhedra. Journal of Algorithms, 21(1):80–101, July 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900383.

Pemantle:2005:CRA

[Pem05a] Robin Pemantle. Cycles in random k-ary maps and the poor performance of random number generation. Journal of Algorithms, 54(1):72–84, January 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001561.

Pemantle:2005:PMD

[Pem05b] Robin Pemantle. A probabilistic model for the degree of the cancellation polynomial in Gosper’s algorithm. Journal of Algorithms, 54(1):58–71, January 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000744.

Perl:1984:OST

[Per84] Yehoshua Perl. Optimum split trees. Journal of Algorithms, 5(3):367–374, September 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677484900178.

Plaisted:1987:HTA

[PH87] David A. Plaisted and Jiarong Hong. A heuristic triangulation algorithm. Journal of Algorithms, 8(3):405–437, September 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900204.

Pippenger:2002:ACP

[Pip02] Nicholas Pippenger. Analysis of carry propagation in addition: an elementary approach. Journal of Algorithms, 42(2):317–333, February 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677402912162.

Pisinger:1999:LTA

[Pis99] David Pisinger. Linear time algorithms for knapsack problems with bounded weights. Journal of Algorithms, 33(1):1–14, October 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910349.

Pittel:1987:LPP

[Pit87] B. Pittel. Linear probing: the probable largest search time grows logarithmically with the number of records. Journal of Algorithms, 8(2):236–249, June 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667748790040X.

Plaisted:1984:HMG

[Pla84] David A. Plaisted. Heuristic matching for graphs satisfying the triangle inequality. Journal of Algorithms, 5(2):163–179, June 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677484900245.

Poblete:1985:AFH

[PM85] Patricio V. Poblete and J. Ian Munro. The analysis of a fringe heuristic for binary search trees. Journal of Algorithms, 6(3):336– 350, September 1985. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677485900033.

Poblete:1989:LCF

[PM89] Patricio V. Poblete and J. Ian Munro. Last-come-first-served hashing. Journal of Algorithms, 10(2):228–248, June 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748990014X.

Poonen:1993:WCS

[Poo93] B. Poonen. The worst case in Shellsort and related algorithms. Journal of Algorithms, 15(1):101–124, July 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0196677483710321. For other Shellsort bounds, see [Yao80, Sed86, WS90, PS97, Goo11].

Panaite:1999:EUU

[PP99] Petrisor Panaite and Andrzej Pelc. Exploring unknown undirected graphs. Journal of Algorithms, 33(2):281–295, November 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667749991043X.

Panholzer:2004:PPA

[PPR04] Alois Panholzer, Helmut Prodinger, and Marko Riedel. Permuting in place: analysis of two stopping rules. Journal of Algorithms, 51(2):170–184, May 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404000288.

Proskurowski:1985:BTG

[PR85] Andrzej Proskurowski and Frank Ruskey. Binary tree gray codes. Journal of Algorithms, 6(2):225–238, June 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900409.

Paterson:1991:SMB

[PR91] M. S. Paterson and A. A. Razborov. The set of minimal braids is co-NP-complete. Journal of Algorithms, 12(3):393–408, September 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749190011M.

Pagh:2004:CH

[PR04] Rasmus Pagh and Flemming Friche Rodler. Cuckoo hashing. Journal of Algorithms, 51(2):122–144, May 2004. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403001925.

Pritchard:1983:FCP

[Pri83] Paul Pritchard. Fast compact prime number sieves (among others). Journal of Algorithms, 4(4):332–344, December 1983. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677483900147.

Pritchard:1999:CSG

[Pri99] Paul Pritchard. On computing the subset graph of a collection of sets. Journal of Algorithms, 33(2):187–203, November 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910325.

Pruhs:2003:F

[Pru03] Kirk Pruhs. Foreword. Journal of Algorithms, 48(1):1, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000427.

Policastro:2009:LSA

[PRZP09] Claudio A. Policastro, Roseli A. F. Romero, Giovana Zuliani, and Ednaldo Pizzolato. Learning of shared attention in sociable robotics. Journal of Algorithms, 64(4):139–151, October 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677409000352.

Pomerance:1995:CIF

[PS95] C. Pomerance and J. Sorenson. Counting the integers factorable via cyclotomic methods. Journal of Algorithms, 19(2):250–265, September 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S019667748571036X.

Panconesi:1996:CDN

[PS96] Alessandro Panconesi and Aravind Srinivasan. On the complexity of distributed network decomposition. Journal of Algorithms, 20(2):356–374, March 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900176.

Plaxton:1997:LBS

[PS97] C. Greg Plaxton and Torsten Suel. Lower bounds for Shellsort. Journal of Algorithms, 23(2):221–240, May 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0196677496908251. For other Shellsort bounds, see [Yao80, Sed86, WS90, Poo93, Goo11].

Promel:2000:NAA

[PS00] Hans Jürgen Prömel and Angelika Steger. A new approximation algorithm for the Steiner tree problem with performance ratio 5/3. Journal of Algorithms, 36(1):89–101, July 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910861.

Penn:1997:NIA

[PSK97] Michal Penn and Haya Shasha-Krupnik. NOTE improved approximation algorithms for weighted 2- and 3-vertex connectivity augmentation problems. Journal of Algorithms, 22(1):187– 196, January 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485708000.

Peng:1993:ACT

[PSY93] S. T. Peng, A. B. Stephens, and Y. Yesha. Algorithms for a core and k-tree core of a tree. Journal of Algorithms, 15(1): 143–159, July 1993. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677483710345.

Peng:1996:EAF

[PtL96] Shietung Peng and Win tsung Lo. Efficient algorithms for finding a core of a tree with a specified length. Journal of Algorithms, 20(3):445–458, May 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749690022X.

Papadimitriou:1984:TGP

[PV84] Christos H. Papadimitriou and Umesh V. Vazirani. On two geometric problems related to the travelling salesman problem. Journal of Algorithms, 5(2):231–246, June 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900294.

Pearson:1993:ESP

[PV93] D. Pearson and V. V. Vazirani. Efficient sequential and parallel algorithms for maximal bipartite sets. Journal of Algorithms, 14(2):171–179, March 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677483710084.

Panigrahy:1998:AAA

[PV98] Rina Panigrahy and Sundar Vishwanathan. An O(log ∗n) approximation algorithm for the asymmetric p-center problem. Journal of Algorithms, 27(2):259–268, May 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909214.

Prodinger:2002:P

[PV02] Helmut Prodinger and Brigitte Vallée. Preface. Journal of Algorithms, 44(1):1–3, July 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002195.

Pittel:1997:LCS

[PW97] Boris Pittel and Robert S. Weishaar. On-line coloring of sparse random graphs and random trees. Journal of Algorithms, 23(1): 195–205, April 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908275.

Propp:1998:HGP

[PW98] James Gary Propp and David Bruce Wilson. How to get a perfectly random sample from a generic Markov chain and generate a random spanning tree of a directed graph. Journal of Algorithms, 27(2):170–217, May 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497909172.

Paterson:1986:PRP

[PY86] Michael S. Paterson and F. Frances Yao. Point retrieval for polygons. Journal of Algorithms, 7(3):441–447, September 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900337.

Paterson:1992:OBS

[PY92] Michael S. Paterson and F. Frances Yao. Optimal binary space partitions for orthogonal objects. Journal of Algorithms, 13(1): 99–113, March 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749290007Y.

Queyranne:2002:AAG

[QS02] Maurice Queyranne and Maxim Sviridenko. A (2 + ϵ)- approximation algorithm for the generalized preemptive open shop problem with minsum objective. Journal of Algorithms, 45(2):202–212, November 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677402002511.

Ragde:1993:PSC

[Rag93] P. Ragde. The parallel simplicity of compaction and chaining. Journal of Algorithms, 14(3):371–380, May 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710199.

Rajasekaran:1995:RSC

[Raj95] S. Rajasekaran. k–k routing, k–k sorting, and cut-through routing on the mesh. Journal of Algorithms, 19(3):361–382, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710425.

Rajasekaran:1998:SMC

[Raj98] Sanguthevar Rajasekaran. Selection on mesh connected computers with fixed and reconfigurable buses. Journal of Algorithms, 29(1):68–81, October 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667749890966X.

Ralston:1981:NMA

[Ral81] A. Ralston. A new memoryless algorithm for de Bruijn sequences. Journal of Algorithms, 2(1):50–62, March 1981. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677481900079.

Ramachandran:1988:FMF

[Ram88] Vijaya Ramachandran. Finding a minimum feedback arc set in reducible flow graphs. Journal of Algorithms, 9(3):299–313, September 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900223.

Ramachandran:1995:EF

[Ram95a] V. Ramachandran. Editor’s foreword. Journal of Algorithms, 18 (3):377, May 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710152.

Ramesh:1995:TLG

[Ram95b] H. Ramesh. On traversing layered graphs on-line. Journal of Algorithms, 18(3):480–512, May 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667748571019X.

Ramachandran:1997:PAR

[Ram97] Vijaya Ramachandran. Parallel algorithms for reducible flow graphs. Journal of Algorithms, 23(1):1–31, April 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908202.

Rangan:1983:MNA

[Ran83] C. Pandu Rangan. On the minimum number of additions required to compute a quadratic form. Journal of Algorithms, 4(3):282– 285, September 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900275.

Ramanan:1989:LPL

[RBLL89] Prakash Ramanan, Donna J. Brown, C. C. Lee, and D. T. Lee. On-line bin packing in linear time. Journal of Algorithms, 10(3):305–326, September 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667748990031X.

Raghavan:1986:SBW

[RCS86] Raghunath Raghavan, James Cohoon, and Sartaj Sahni. Single bend wiring. Journal of Algorithms, 7(2):232–257, June 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677486900064.

Ramanan:1984:PAP

[RDL84] P. Ramanan, J. S. Deogun, and C. L. Liu. A personnel assignment problem. Journal of Algorithms, 5(1):132–144, March 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900439.

Ramanan:1984:NAS

[RH84] Prakash V. Ramanan and Laurent Hyafil. New algorithms for selection. Journal of Algorithms, 5(4):557–578, December 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677484900087.

Richards:1986:FSC

[Ric86] Dana Richards. Finding short cycles in planar graphs using separators. Journal of Algorithms, 7(3):382–394, September 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677486900295.

Rahman:2000:BRD

[RiNN00] Md. Saidur Rahman, Shin ichi Nakano, and Takao Nishizeki. Box-rectangular drawings of plane graphs. Journal of Algorithms, 37(2):363–398, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911052.

Rahman:2004:RDP

[RNG04] Md. Saidur Rahman, Takao Nishizeki, and Shubhashis Ghosh. Rectangular drawings of planar graphs. Journal of Algorithms, 50(1):62–78, January 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403001263.

Robson:1986:AMI

[Rob86] J. M. Robson. Algorithms for maximum independent sets. Journal of Algorithms, 7(3):425–440, September 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900325.

Ronyai:1988:FPF

[Rón88a] Lajos Rónyai. Factoring polynomials over finite fields. Journal of Algorithms, 9(3):391–400, September 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900296.

Ronyai:1988:ZDQ

[Rón88b] Lajos Rónyai. Zero divisors in quaternion algebras. Journal of Algorithms, 9(4):494–506, December 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900144.

Ronn:1990:NCS

[Ron90] Eytan Ronn. NP-complete stable matching problems. Journal of Algorithms, 11(2):285–304, June 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677490900072.

Ronyai:1994:DMC

[Ron94] L. Ronyai. A deterministic method for computing splitting elements in simple algebras over Q. Journal of Algorithms, 16(1): 24–32, January 1994. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677484710029.

Rosen:1982:RLA

[Ros82] Barry K. Rosen. Robust linear algorithms for cutsets. Journal of Algorithms, 3(3):205–217, September 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900207.

Roura:2001:DAC

[Rou01] Salvador Roura. Digital access to comparison-based tree data structures and algorithms. Journal of Algorithms, 40(1):1–23, July 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677401911605.

Ruskey:1990:GBT

[RP90] Frank Ruskey and Andrzej Proskurowski. Generating binary trees by transpositions. Journal of Algorithms, 11(1):68–84, March 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749090030I.

Ramalingam:1996:IAG

[RR96] G. Ramalingam and Thomas Reps. An incremental algorithm for a generalization of the shortest-path problem. Journal of Algorithms, 21(2):267–305, September 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900462.

Robertson:1986:GMI

[RS86] Neil Robertson and P. D. Seymour. Graph minors. II. Algorithmic aspects of tree-width. Journal of Algorithms, 7(3):309–322, September 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677486900234.

Raghavan:2003:RAR

[RS03] Vijay Raghavan and Jeremy Spinrad. Robust algorithms for restricted domains. Journal of Algorithms, 48(1):160–172, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677403000488.

Ruskey:1992:GN

[RSW92] Frank Ruskey, Carla Savage, and Terry Min Yih Wang. Generating necklaces. Journal of Algorithms, 13(3):414–430, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290047G.

Rosenstiehl:1984:GCP

[RT84] Pierre Rosenstiehl and Robert E. Tarjan. Gauss codes, planar Hamiltonian graphs, and stack-sortable permutations. Journal of Algorithms, 5(3):375–390, September 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748490018X.

Rodl:1987:MOL

[RT87] V. Rödl and C. Tovey. Multiple optima in local search. Journal of Algorithms, 8(2):250–259, June 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900411.

Reif:1997:DAA

[RT97] John H. Reif and Stephen R. Tate. On dynamic algorithms for algebraic problems. Journal of Algorithms, 22(2):347–371, February 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485708073.

Rub:1997:ART

[Rüb97] Christine Rüb. On the average running time of odd-even merge sort. Journal of Algorithms, 22(2):329–346, February 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908123.

Ruppert:1995:DRA

[Rup95] J. Ruppert. A Delaunay refinement algorithm for quality 2dimensional mesh generation. Journal of Algorithms, 18(3):548– 585, May 1995. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677485710218.

Ruskey:1988:AIG

[Rus88] Frank Ruskey. Adjacent interchange generation of combinations. Journal of Algorithms, 9(2):162–180, June 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900363.

Rabin:1989:MMG

[RV89] Michael O. Rabin and Vijay V. Vazirani. Maximum matchings in general graphs through randomization. Journal of Algorithms, 10(4):557–567, December 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900059.

Ramarao:1992:FUS

[RV92] K. V. S. Ramarao and S. Venkatesan. On finding and updating shortest paths distributively. Journal of Algorithms, 13(2):235– 257, June 1992. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677492900177.

Remmel:1984:MSF

[RW84] J. B. Remmel and R. Whitney. Multiplying Schur functions. Journal of Algorithms, 5(4):471–487, December 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900026.

Seidel:2000:EWC

[SA00] Raimund Seidel and Udo Adamy. On the exact worst case query complexity of planar point location. Journal of Algorithms, 37(1):189–217, October 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400911015.

Sadakane:2003:NTI

[Sad03] Kunihiko Sadakane. New text indexing functionalities of the compressed suffix arrays. Journal of Algorithms, 48(2):294–313, September 2003. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000877.

Sagan:1980:SRS

[Sag80] Bruce Sagan. On selecting a random shifted young tableau. Journal of Algorithms, 1(3):213–234, September 1980. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677480900103.

Savage:1989:GCS

[Sav89] Carla D. Savage. Gray code sequences of partitions. Journal of Algorithms, 10(4):577–595, December 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900072.

Schmid:1995:NLS

[SB95] U. Schmid and J. Blieberger. On nonpreemptive LCFS scheduling with deadlines. Journal of Algorithms, 18(1):124–158, January 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710048.

Sarkar:1994:IGA

[SCGD94] U. K. Sarkar, P. P. Chakrabarti, S. Ghose, and S. C. Desarkar. Improving greedy algorithms by lookahead-search. Journal of Algorithms, 16(1):1–23, January 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710017.

Schnorr:1982:RAI

[Sch82] C. P. Schnorr. Refined analysis and improvements on some factoring algorithms. Journal of Algorithms, 3(2):101–127, June 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677482900128.

Schaffer:1988:TUB

[Sch88a] Alejandro A. Schäffer. A tighter upper bound on the worst case behavior of Conway’s parallel sorting algorithm. Journal of Algorithms, 9(3):321–342, September 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900247.

Schnorr:1988:MEA

[Sch88b] C. P. Schnorr. A more efficient algorithm for lattice basis reduction. Journal of Algorithms, 9(1):47–62, March 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900041.

Schonhage:1988:PCI

[Sch88c] A. Schönhage. Probabilistic computation of integer polynomial GCDs. Journal of Algorithms, 9(3):365–371, September 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900272.

Schieber:1998:CMW

[Sch98]  Baruch Schieber.  Computing a minimum weight k-link

path in graphs with the concave Monge property. Journal of Algorithms, 29(2):204–222, November 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909555.

Schirokauer:2005:VL

[Sch05a] Oliver Schirokauer. Virtual logarithms. Journal of Algorithms, 57(2):140–147, November 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001579.

Schuler:2005:ASP

[Sch05b] Rainer Schuler. An algorithm for the satisfiability problem of formulas in conjunctive normal form. Journal of Algorithms, 54(1): 40–44, January 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000768.

Sibeyn:1997:DPR

[SCK97] Jop F. Sibeyn, Bogdan S. Chlebus, and Michael Kaufmann. Deterministic permutation routing on meshes. Journal of Algorithms, 22(1):111–141, January 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485708048.

Stathis:2009:P

[SdGG09] Kostas Stathis, Artur d’Avila Garcez, and Robert Givan. Preface. Journal of Algorithms, 64(4):125–126, October 2009. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677409000339.

Sedgewick:1986:NUB

[Sed86] Robert Sedgewick. A new upper bound for Shellsort. Journal of Algorithms, 7(2):159–173, June 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900015.

For other Shellsort bounds, see [Yao80, WS90, Poo93, PS97, Goo11].

Sellers:1980:TCE

[Sel80] Peter H. Sellers. The theory and computation of evolutionary distances: Pattern recognition. Journal of Algorithms, 1(4):359– 373, December 1980. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677480900164.

Semba:1984:EAG

[Sem84] Ichiro Semba. An efficient algorithm for generating all k-subsets (1 ≤ k ≤ m ≤ n) of the set [1, 2, …, n] in lexicographical order. Journal of Algorithms, 5(2):281–283, June 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900312.

Sen:1995:FCR

[Sen95] S. D. Sen. Fractional cascading revisited. Journal of Algorithms, 19(2):161–172, September 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677485710322.

Sun:2004:IPS

[SFBY04] Fangting Sun, David Fernández-Baca, and Wei Yu. Inverse parametric sequence alignment. Journal of Algorithms, 53(1):36– 54, October 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000719.

Sgall:1996:RLS

[Sga96] Jiří Sgall. Randomized on-line scheduling of parallel jobs. Journal of Algorithms, 21(1):149–175, July 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900413.

Sunder:1998:SIO

[SH98] Sivaprakasam Sunder and Xin He. Scheduling interval ordered tasks in parallel. Journal of Algorithms, 26(1):34–47, January 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497908956.

Hsu:2000:FCT

[sH00] Tsan sheng Hsu. On four-connecting a triconnected graph. Journal of Algorithms, 35(2):202–234, May 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910770.

Hsu:2002:SFB

[sH02a] Tsan sheng Hsu. Simpler and faster biconnectivity augmentation. Journal of Algorithms, 45(1):55–71, October 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002237.

Stearns:2002:ESQ

[SH02b] Richard E. Stearns and Harry B. Hunt III. Exploiting structure in quantified formulas. Journal of Algorithms, 43(2):220–263, May 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677402000020.

Sherk:1995:SAA

[She95] M. Sherk. Self-adjusting k-ary search trees. Journal of Algorithms, 19(1):25–44, July 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677485710267.

Shiloach:1981:FCC

[Shi81] Yossi Shiloach. Fast canonization of circular strings. Journal of Algorithms, 2(2):107–121, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900134.

Shmoys:2004:F

[Shm04] David Shmoys. Foreword. Journal of Algorithms, 50(2):133, February 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677403000956.

Schmeichel:1990:PFI

[SHOS90] E. Schmeichel, S. L. Hakimi, M. Otsuka, and G. Sullivan. A parallel fault identification algorithm. Journal of Algorithms, 11(2): 231–241, June 1990. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749090004X.

Shparlinski:2000:CJS

[Shp00] Igor E. Shparlinski. Computing Jacobi symbols modulo sparse integers and polynomials and some applications. Journal of Algorithms, 36(2):241–252, August 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910915.

Sibeyn:2006:ES

[Sib06] Jop F. Sibeyn. External selection. Journal of Algorithms, 58(2):104–117, February 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677405000544.

Siegel:2001:MBT

[Sie01] Alan Siegel. Median bounds and their application. Journal of Algorithms, 38(1):184–236, January 2001. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911386.

Suzuki:1992:VPQ

[SIN92] Hitoshi Suzuki, Akira Ishiguro, and Takao Nishizeki. Variablepriority queue and doughnut routing. Journal of Algorithms, 13(4):606–635, December 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749290059L.

Sitters:2005:CPM

[Sit05] René Sitters. Complexity of preemptive minsum scheduling on unrelated parallel machines. Journal of Algorithms, 57(1):37–48, September 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404001130.

Skiena:1991:PCP

[Ski91] Steven S. Skiena. Probing convex polygons with half-planes. Journal of Algorithms, 12(3):359–374, September 1991. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190009N.

Skodinis:2003:CLT

[Sko03] Konstantin Skodinis. Construction of linear tree-layouts which are optimal with respect to vertex separation in linear time. Journal of Algorithms, 47(1):40–59, April 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002250.

Seroussi:1984:SAB

[SL84] G. Seroussi and A. Lempel. On symmetric algorithms for bilinear forms over finite fields. Journal of Algorithms, 5(3):327–344, September 1984. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677484900142.

Sagan:2006:ASR

[SL06] Bruce E. Sagan and Jaejin Lee. An algorithmic sign-reversing involution for special rim-hook tableaux. Journal of Algorithms, 59(2):149–161, May 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740400104X.

Slavik:1997:TAG

[Sla97] Petr Slavík. A tight analysis of the greedy algorithm for set cover. Journal of Algorithms, 25(2):237–254, November 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908877.

Subramanian:2002:GPR

[SM02] C. R. Subramanian and C. E. Veni Madhavan. General partitioning on random graphs. Journal of Algorithms, 42(1):153– 172, January 2002. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677401911976.

Snir:1986:DST

[Sni86] Marc Snir. Depth-size trade-offs for parallel prefix computation. Journal of Algorithms, 7(2):185–201, June 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900039.

Soroker:1988:FPA

[Sor88a] Danny Soroker. Fast parallel algorithms for finding Hamiltonian paths and cycles in a tournament. Journal of Algorithms, 9(2): 276–286, June 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900429.

Soroker:1988:FPS

[Sor88b] Danny Soroker. Fast parallel strong orientation of mixed graphs and related augmentation problems. Journal of Algorithms, 9(2): 205–223, June 1988. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677488900387.

Sorenson:1994:TFG

[Sor94] J. Sorenson. Two fast GCD algorithms. Journal of Algorithms, 16(1):110–144, January 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710066.

Spinrad:1994:RCG

[Spi94] J. Spinrad. Recognition of circle graphs. Journal of Algorithms, 16(2):264–282, March 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710121.

Sprugnoli:1989:ASP

[Spr89] Renzo Sprugnoli. The analysis of a simple in-place merging algorithm. Journal of Algorithms, 10(3):366–380, September 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677489900345.

Squire:1998:GAO

[Squ98] Matthew B. Squire. Generating the acyclic orientations of a graph. Journal of Algorithms, 26(2):275–290, February 1998. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908919.

Sawada:2003:GLB

[SR03] J. Sawada and F. Ruskey. Generating Lyndon brackets: an addendum to: “Fast algorithms to generate necklaces, unlabeled necklaces and irreducible polynomials over GF(2). Journal of Algorithms, 46(1):21–26, January 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002869. See [CRS^{+}00].

Sun:2006:FAO

[SR06] Zheng Sun and John H. Reif. On finding approximate optimal paths in weighted regions. Journal of Algorithms, 58(1):1–32, January 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404001191.

Safro:2006:GML

[SRB06] Ilya Safro, Dorit Ron, and Achi Brandt. Graph minimum linear arrangement by multilevel weighted edge contractions. Journal of Algorithms, 60(1):24–41, July 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001531.

Salowe:1987:SSM

[SS87] Jeffrey Salowe and William Steiger. Simplified stable merging tasks. Journal of Algorithms, 8(4):557–571, December 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900502.

Schaffer:1993:AH

[SS93] R. Schaffer and R. Sedgewick. The analysis of Heapsort. Journal of Algorithms, 15(1):76–100, July 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S019667748371031X. For other analyses of Heapsort, see [LP93, BFF96].

Shi:1999:TWT

[SS99] Hanmao Shi and Thomas H. Spencer. Time-work tradeoffs of the single-source shortest paths problem. Journal of Algorithms, 30(1):19–32, January 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909683.

Shahrokhi:2000:CSD

[SS00a] Farhad Shahrokhi and Weiping Shi. On crossing sets, disjoint sets, and pagenumber. Journal of Algorithms, 34(1):40–53, January 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910490.

Smith:2000:RTM

[SS00b] Adam Smith and Subhash Suri. Rectangular tiling in multidimensional arrays. Journal of Algorithms, 37(2):451–467, November 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740091109X.

Sanders:2001:HHH

[SSO01] Peter Sanders and Roberto Solis-Oba. How helpers hasten hrelations. Journal of Algorithms, 41(1):86–98, October 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911691.

Soisalon-Soininen:1984:OAC

[SSW84] Eljas Soisalon-Soininen and Derick Wood. Optimal algorithms to compute the closure of a set of iso-rectangles. Journal of Algorithms, 5(2):199–214, June 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900270.

Skowronska:1986:ACT

[SSZ86] Miroslawa Skowrońska, Maciej M. Syslo, and Cristina Zamfirescu. An algorithmic characterization of total digraphs. Journal of Algorithms, 7(1):120–133, March 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900416.

Shamir:1999:FSI

[ST99] Ron Shamir and Dekel Tsur. Faster subtree isomorphism. Journal of Algorithms, 33(2):267–280, November 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910441.

Steiner:1987:SDP

[Ste87] George Steiner. Searching in 2-dimensional partial orders. Journal of Algorithms, 8(1):95–105, March 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900290.

Stout:1985:PCS

[Sto85] Quentin F. Stout. Pyramid computer solutions of the closest pair problem. Journal of Algorithms, 6(2):200–212, June 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900380.

Shamir:1984:SDG

[SU84] Eli Shamir and Eli Upfal. Sequential and distributed graph coloring algorithms with performance analysis in random graph spaces. Journal of Algorithms, 5(4):488–501, December 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900038.

Shioura:1997:LTA

[SU97] Akiyoshi Shioura and Takeaki Uno. A linear time algorithm for finding a k-tree core. Journal of Algorithms, 23(2):281–290, May 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667749690838X.

Subramanian:1996:ES

[Sub96] Ashok Subramanian. An explanation of splaying. Journal of Algorithms, 20(3):512–525, May 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677496900255.

Subramanian:1999:MCC

[Sub99] C. R. Subramanian. Minimum coloring k-colorable graphs in polynomial average time. Journal of Algorithms, 33(1):112–123, October 1999. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910350.

Shiloach:1981:FMM

[SV81] Yossi Shiloach and Uzi Vishkin. Finding the maximum, merging, and sorting in a parallel computation model. Journal of Algorithms, 2(1):88–102, March 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900109.

Shiloach:1982:PCA

[SV82a] Yossi Shiloach and Uzi Vishkin. An O(log n) parallel connectivity algorithm. Journal of Algorithms, 3(1):57–67, March 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900086.

Shiloach:1982:PMF

[SV82b] Yossi Shiloach and Uzi Vishkin. An O(n^{2} log n) parallel max-flow algorithm. Journal of Algorithms, 3(2):128–146, June 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748290013X.

Schaffer:1987:CHP

[SV87] Alejandro A. Schäffer and Christopher J. Van Wyk. Convex hulls of piecewise-smooth Jordan curves. Journal of Algorithms, 8(1): 66–94, March 1987. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677487900289.

Swart:1985:FCH

[Swa85] Garret Swart. Finding the convex hull facet by facet. Journal of Algorithms, 6(1):17–48, March 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900173.

John:2003:PSP

[SWMV03] Katherine St. John, Tandy Warnow, Bernard M. E. Moret, and Lisa Vawter. Performance study of phylogenetic methods: (unweighted) quartet methods and neighbor-joining. Journal of Algorithms, 48(1):173–193, August 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S019667740300049X.

Steele:1982:LBA

[SY82] J. Michael Steele and Andrew C. Yao. Lower bounds for algebraic decision trees. Journal of Algorithms, 3(1):1–8, March 1982. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677482900025.

Shasha:1990:FAU

[SZ90] Dennis Shasha and Kaizhong Zhang. Fast algorithms for the unit cost editing distance between trees. Journal of Algorithms, 11(4):581–621, December 1990. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677490900113.

Saks:1999:OSD

[SZ99] Michael Saks and Fotios Zaharoglou. Optimal space distributed order-preserving lists. Journal of Algorithms, 31(2):320–334, May 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677498910008.

Schoenmeyr:2005:FBA

[SZ05a] Tor Schoenmeyr and David Yu Zhang. FFT-based algorithms for the string matching with mismatches problem. Journal of Algorithms, 57(2):130–139, November 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000180.

Sparl:2005:LCA

[ŠŽ05b] Petra Šparl and Janez Žerovnik. 2-local 4/3-competitive algorithm for multicoloring hexagonal graphs. Journal of Algorithms, 55(1):29–41, April 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001476.

Szpankowski:1988:SRA

[Szp88] Wojciech Szpankowski. Some results on V -ary asymmetric tries. Journal of Algorithms, 9(2):224–244, June 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900399.

Szymanski:1985:HTR

[Szy85] Thomas G. Szymanski. Hash table reorganization. Journal of Algorithms, 6(3):322–335, September 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900021.

Thanh:1986:OET

[TAB86] Mai Thanh, V. S. Alagar, and T. D. Bui. Optimal expectedtime algorithms for merging. Journal of Algorithms, 7(3):341– 357, September 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748690026X.

Tamminen:1985:TLG

[Tam85] Markku Tamminen. Two levels are as good as any. Journal of Algorithms, 6(1):138–144, March 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900240.

Tamassia:1996:LPG

[Tam96] Roberto Tamassia. On-line planar graph embedding. Journal of Algorithms, 21(2):201–239, September 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900449.

Tan:1991:NSC

[Tan91] Jimmy J. M. Tan. A necessary and sufficient condition for the existence of a complete stable matching. Journal of Algorithms, 12(1):154–178, March 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190028W.

Tardos:1998:EF

[Tar98] Éva Tardos. Editor’s foreword. Journal of Algorithms, 27(2): 169, May 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497909160.

Teng:1999:LEM

[Ten99] Shang-Hua Teng. Low energy and mutually distant sampling. Journal of Algorithms, 30(1):52–67, January 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909634.

Thomassen:1989:GGP

[Tho89] Carsten Thomassen. The graph genus problem is NP-complete. Journal of Algorithms, 10(4):568–576, December 1989. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677489900060.

Thomassen:1993:ECP

[Tho93] C. Thomassen. The even cycle problem for planar digraphs. Journal of Algorithms, 15(1):61–75, July 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710308.

Thorup:1996:EPS

[Tho96] Mikkel Thorup. Efficient preprocessing of simple binary pattern forests. Journal of Algorithms, 20(3):602–612, May 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900292.

Thorup:1997:PSR

[Tho97] Mikkel Thorup. Parallel shortcutting of rooted trees. Journal of Algorithms, 23(1):139–159, April 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908299.

Thorup:1999:DDC

[Tho99] Mikkel Thorup. Decremental dynamic connectivity. Journal of Algorithms, 33(2):229–243, November 1999. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910337.

Thorup:2000:FIS

[Tho00] Mikkel Thorup. Floats, integers, and single source shortest paths. Journal of Algorithms, 35(2):189–201, May 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400910800.

Thorup:2002:RST

[Tho02] Mikkel Thorup. Randomized sorting in O(n log log n) time and linear space using addition, shift, and bit-wise Boolean operations. Journal of Algorithms, 42(2):205–230, February 2002. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402912113.

Thurimella:1997:SLD

[Thu97] Ramakrishna Thurimella. Sub-linear distributed algorithms for sparse certificates and biconnected components. Journal of Algorithms, 23(1):160–179, April 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908329.

Tiwari:1986:EPA

[Tiw86] Prasoon Tiwari. An efficient parallel algorithm for shifting the root of a depth first spanning tree. Journal of Algorithms, 7(1):105–119, March 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900404.

Tseng:1996:CNS

[TL96] Paul Tseng and Zhi-Quan Luo. On computing the nested sums and infimal convolutions of convex piecewise-linear functions. Journal of Algorithms, 21(2):240–266, September 1996. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900450.

Tolimieri:1982:NMA

[Tol82] R. Tolimieri. A nonquadratic minimal algorithm for a system of quadratic forms. Journal of Algorithms, 3(1):31–40, March 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677482900050.

Tamir:2005:CLP

[TPMRC05] A. Tamir, J. Puerto, J. A. Mesa, and A. M. Rodríguez-Chía. Conditional location of path and tree shaped facilities on trees. Journal of Algorithms, 56(1):50–75, July 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000222.

Treur:2009:PFS

[Tre09] Jan Treur. Past-future separation and normal forms in temporal predicate logic specifications. Journal of Algorithms, 64(2–3):106– 124, April/July 2009. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677409000108.

Tsaban:2003:PGF

[Tsa03] Boaz Tsaban. Permutation graphs, fast forward permutations, and sampling the cycle structure of a permutation. Journal of Algorithms, 47(2):104–121, July 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000178.

Thilikos:2005:CLT

[TSB05a] Dimitrios M. Thilikos, Maria Serna, and Hans L. Bodlaender. Cutwidth I: a linear time fixed parameter algorithm. Journal of Algorithms, 56(1):1–24, July 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000167.

Thilikos:2005:CIA

[TSB05b] Dimitrios M. Thilikos, Maria Serna, and Hans L. Bodlaender. Cutwidth II: Algorithms for partial w-trees of bounded degree. Journal of Algorithms, 56(1):25–49, July 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000179.

Tung:1987:CCD

[Tun87] S. P. Tung. Computational complexities of Diophantine equations with parameters. Journal of Algorithms, 8(3):324–336, September 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900137.

Turner:1988:AAC

[Tur88] Jonathan S. Turner. Almost all k-colorable graphs are easy to color. Journal of Algorithms, 9(1):63–82, March 1988. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677488900053.

Tucker:1984:ACP

[TW84] Alan Tucker and Donna Wilson. An O(N^{2}) algorithm for coloring perfect planar graphs. Journal of Algorithms, 5(1):60–68, March 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677484900403.

Ukkonen:1985:FAP

[Ukk85] Esko Ukkonen. Finding approximate patterns in strings. Journal of Algorithms, 6(1):132–137, March 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900239.

Vakhania:2003:BAS

[Vak03] Nodari Vakhania. A better algorithm for sequencing with release and delivery times on identical machines. Journal of Algorithms, 48(2):273–293, September 2003. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677403000725.

Valiant:1984:SMF

[Val84] L. G. Valiant. Short monotone formulae for the majority function. Journal of Algorithms, 5(3):363–366, September 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900166.

Vallee:1991:GAR

[Val91] Brigitte Vallée. Gauss’ algorithm revisited. Journal of Algorithms, 12(4):556–572, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190033U.

vanBaronaigien:2000:LGC

[vB00] Dominique Roelants van Baronaigien. A loopless Gray-code algorithm for listing k-ary trees. Journal of Algorithms, 35(1):100– 107, April 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677499910738.

Boas:1991:Pa

[vEB91a] Peter van Emde Boas. Problems. Journal of Algorithms, 12(1):183–185, March 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677491900303.

Boas:1991:Pb

[vEB91b] Peter van Emde Boas. Problems. Journal of Algorithms, 12(4):700–703, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/019667749190042W.

Veldhorst:1986:ORD

[Vel86] Marinus Veldhorst. The optimal representation of disjoint isooriented rectangles in two-dimensional trees. Journal of Algorithms, 7(1):1–34, March 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677486900362.

Venkatesan:1987:ICS

[Ven87] Shankar M. Venkatesan. Improved constants for some separator theorems. Journal of Algorithms, 8(4):572–578, December 1987. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677487900514.

Verma:1994:GMM

[Ver94] R. M. Verma. A general method and a master theorem for divide-and-conquer recurrences with applications. Journal of Algorithms, 16(1):67–79, January 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710042.

Vinyals:2008:TSE

[VGC^{+}08] Meritxell Vinyals, Andrea Giovannucci, Jesús Cerquides, Pedro Meseguer, and Juan A. Rodriguez-Aguilar. A test suite for the evaluation of mixed multi-unit combinatorial auctions. Journal of Algorithms, 63(1–3):130–150, January/July 2008. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677408000102.

Villard:2000:PEP

[Vil00] Gilles Villard. Processor efficient parallel solution of linear systems of equations. Journal of Algorithms, 35(1):122–126, April 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677499910696.

Vishkin:1983:ISM

[Vis83] Uzi Vishkin. Implementation of simultaneous memory address access in models that forbid it. Journal of Algorithms, 4(1):45– 50, March 1983. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677483900330.

Vishkin:1992:PBF

[Vis92a] Uzi Vishkin. A parallel blocking flow algorithm for acyclic networks. Journal of Algorithms, 13(3):489–501, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290051D.

Vishwanathan:1992:ROG

[Vis92b] Sundar Vishwanathan. Randomized online graph coloring. Journal of Algorithms, 13(4):657–669, December 1992. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749290061G.

Vishwanathan:2004:AAF

[Vis04] Sundar Vishwanathan. An approximation algorithm for finding long paths in Hamiltonian graphs. Journal of Algorithms, 50(2):246–256, February 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677403000932.

Vitter:1982:DAH

[Vit82] Jeffrey Scott Vitter. Deletion algorithms for hashing that preserve randomness. Journal of Algorithms, 3(3):261–275, September 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677482900244.

vanLeeuwen:1981:MPR

[vLW81] Jan van Leeuwen and Derick Wood. The measure problem for rectangular ranges in d-space. Journal of Algorithms, 2(3):282– 300, September 1981. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0196677481900274.

Valiveti:1993:SOD

[VO93] R. S. Valiveti and B. J. Oommen. Self-organizing doubly-linked lists. Journal of Algorithms, 14(1):88–114, January 1993. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677483710059.

vanOtterlo:2009:IDP

[vO09] Martijn van Otterlo. Intensional dynamic programming. A Rosetta stone for structured dynamic programming. Journal of Algorithms, 64(4):169–191, October 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677409000376.

vanStee:2005:MTC

[vSL05] Rob van Stee and Han La Poutré. Minimizing the total completion time on-line on a single machine, using restarts. Journal of Algorithms, 57(2):95–129, November 2005. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404001506.

vanVliet:1996:AWC

[vV96] André van Vliet. On the asymptotic worst case behavior of Harmonic fit. Journal of Algorithms, 20(1):113–136, January 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667749690005X.

Vaishnavi:1982:RLS

[VW82] V. K. Vaishnavi and D. Wood. Rectilinear line segment intersection, layered segment trees, and dynamization. Journal of Algorithms, 3(2):160–176, June 1982. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677482900165.

Wachs:1989:EDP

[Wac89] Michelle L. Wachs. On an efficient dynamic programming technique of F. F. Yao. Journal of Algorithms, 10(4):518–530, December 1989. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677489900035.

Wahlstrom:2004:EAF

[Wah04] Magnus Wahlström. Exact algorithms for finding minimum transversals in rank-3 hypergraphs. Journal of Algorithms, 51(2):

107–121, May 2004. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000136.

Walsh:1998:GWF

[Wal98] Timothy R. Walsh. Generation of well-formed parenthesis strings in constant worst-case time. Journal of Algorithms, 29(1):165– 173, October 1998. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677498909609.

Wanke:1994:BTW

[Wan94] E. Wanke. Bounded tree-width and LOGCFL. Journal of Algorithms, 16(3):470–491, May 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710224.

Wang:1997:TBS

[Wan97] Biing-Feng Wang. Tighter bounds on the solution of a divideand-conquer maximin recurrence. Journal of Algorithms, 23(2): 329–344, May 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677496908391.

Wang:2000:EPA

[Wan00] Biing-Feng Wang. Efficient parallel algorithms for optimally locating a path and a tree of a specified length in a weighted tree network. Journal of Algorithms, 34(1):90–108, January 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910209.

Wang:2005:LTA

[Wan05] Biing-Feng Wang. Linear time algorithms for the ring loading problem with demand splitting. Journal of Algorithms, 54(1):45– 57, January 2005. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677404000446.

Warshauer:1986:CPS

[War86] Max L. Warshauer. Conway’s parallel sorting algorithm. Journal of Algorithms, 7(2):270–276, June 1986. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900088.

Warnow:1994:TCI

[War94] T. J. Warnow. Tree compatibility and inferring evolutionary history. Journal of Algorithms, 16(3):388–407, May 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710182.

Wang:1993:SLS

[WCP93] B. F. Wang, G. H. Chen, and K. Park. On the set LCS and setset LCS problems. Journal of Algorithms, 14(3):466–477, May 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710254.

Wu:2000:PTA

[WCT00] Bang Ye Wu, Kun-Mao Chao, and Chuan Yi Tang. A polynomial time approximation scheme for optimal product-requirement communication spanning trees. Journal of Algorithms, 36(2):182– 204, August 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400910885.

Weinberger:1984:FNF

[Wei84] P. J. Weinberger. Finding the number of factors of a polynomial. Journal of Algorithms, 5(2):180–186, June 1984. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900257.

Weihe:1997:EDP

[Wei97] Karsten Weihe. Edge-disjoint (s, t)-paths in undirected planar graphs in linear time. Journal of Algorithms, 23(1):121–138, April 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496908317.

Westbrook:2000:LBR

[Wes00] Jeffery Westbrook. Load balancing for response time. Journal of Algorithms, 35(1):1–16, April 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677400910745.

Wang:1997:IAA

[WG97] Lusheng Wang and Dan Gusfield. Improved approximation algorithms for tree alignment. Journal of Algorithms, 25(2):255–273, November 1997. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677497908828.

Wallace:2008:ESV

[WG08] Richard J. Wallace and Diarmuid Grimes. Experimental studies of variable selection strategies based on constraint weights. Journal of Algorithms, 63(1–3):114–129, January/July 2008. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677408000096.

Wilf:1981:USF

[Wil81] Herbert S. Wilf. The uniform selection of free trees. Journal of Algorithms, 2(2):204–207, June 1981. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0196677481900213.

Wilber:1988:CLW

[Wil88] Robert Wilber. The concave least-weight subsequence problem revisited. Journal of Algorithms, 9(3):418–425, September 1988. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0196677488900326.

Wilf:1991:TAS

[Wil91] Herbert S. Wilf. Two algorithms for the sieve method. Journal of Algorithms, 12(1):179–182, March 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190029X.

Williamson:2008:OBP

[Wil08] Jon Williamson. Objective Bayesian probabilistic logic. Journal of Algorithms, 63(4):167–183, October 2008. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677408000205.

Winter:1986:GSP

[Win86] Pawel Winter. Generalized Steiner problem in series-parallel networks. Journal of Algorithms, 7(4):549–566, December 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677486900180.

Winkler:2001:OGD

[Win01] Peter Winkler. Optimality and greed in dynamic allocation. Journal of Algorithms, 41(2):244–261, November 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677401911988.

Wang:2004:EAS

[WLK04] Biing-Feng Wang, Jyh-Jye Lin, and Shan-Chyun Ku. Efficient algorithms for the scaled indexing problem. Journal of Algorithms, 52(1):82–100, July 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S019667740400046X.

Wu:1995:SAA

[WMM95] S. Wu, U. Manber, and E. Myers. A subquadratic algorithm for approximate regular expression matching. Journal of Algorithms, 19(3):346–360, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677485710413.

Wolfson:1986:AEU

[Wol86] Ouri Wolfson. An algorithm for early unlocking of entities in database transactions. Journal of Algorithms, 7(1):146–156, March 1986. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667748690043X.

Wolfson:1987:VLS

[Wol87] Ouri Wolfson. The virtues of locking by symbolic names. Journal of Algorithms, 8(4):536–556, December 1987. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677487900496.

Wood:2002:DCB

[Woo02] David R. Wood. Degree constrained book embeddings. Journal of Algorithms, 45(2):144–154, November 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002493.

Wormald:1984:GRR

[Wor84] Nicholas C. Wormald. Generating random regular graphs. Journal of Algorithms, 5(2):247–280, June 1984. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677484900300.

Werman:1986:BGM

[WPMK86] Michael Werman, Shmuel Peleg, Robert Melter, and T. Y. Kong. Bipartite graph matching for points on a line or a circle. Journal of Algorithms, 7(2):277–284, June 1986. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667748690009X.

Wang:2006:EAC

[WPYK06] Biing-Feng Wang, Shietung Peng, Hong-Yi Yu, and Shan-Chyun Ku. Efficient algorithms for a constrained k-tree core problem in a tree network. Journal of Algorithms, 59(2):107–124, May 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677405000155.

Weiss:1990:TLB

[WS90] Mark Allen Weiss and Robert Sedgewick. Tight lower bounds for Shellsort. Journal of Algorithms, 11(2):242–251, June 1990. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

019667749090005Y. For other Shellsort bounds, see [Yao80, Sed86, Poo93, PS97, Goo11].

Weber:2005:MIG

[WTM05] Kenneth Weber, Vilmar Trevisan, and Luiz Felipe Martins. A modular integer GCD algorithm. Journal of Algorithms, 54(2):152–167, February 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001075.

Wu:1991:OQM

[Wu91] Xiaolin Wu. Optimal quantization by matrix searching. Journal of Algorithms, 12(4):663–673, December 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677491900392.

Wu:2002:PTA

[Wu02] Bang Ye Wu. A polynomial time approximation scheme for the two-source minimum routing cost spanning trees. Journal of Algorithms, 44(2):359–378, August 2002. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677402002055.

Xu:2000:AWC

[Xu00] Kaihong Xu. The asymptotic worst-case behavior of the FFD heuristic for small items. Journal of Algorithms, 37(2):237–246, November 2000. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0196677400911039.

Yannakakis:1994:AMS

[Yan94] M. Yannakakis. On the approximation of maximum satisfiability. Journal of Algorithms, 17(3):475–502, November 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710455.

Yao:1980:AS

[Yao80] Andrew Chi-Chih Yao. An analysis of (h, k, 1)-Shellsort. Journal of Algorithms, 1(1):14–50, March 1980. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0196677480900036. For other Shellsort bounds, see [Sed86, WS90, Poo93, PS97, Goo11].

Yao:1985:OAK

[Yao85] Andrew C. Yao. On optimal arrangements of keys with double hashing. Journal of Algorithms, 6(2):253–264, June 1985. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0196677485900422.

Yeh:2001:SAP

[Yeh01] Wei-Chang Yeh. A simple algorithm for the planar multiway cut problem. Journal of Algorithms, 39(1):68–77, April 2001. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677400911489.

Yellin:1992:RSC

[Yel92] Daniel M. Yellin. Representing sets with constant time equality testing. Journal of Algorithms, 13(3):353–373, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/019667749290044D.

Yeo:1999:PTA

[Yeo99] Anders Yeo. A polynomial time algorithm for finding a cycle covering a given set of vertices in a semicomplete multipartite digraph. Journal of Algorithms, 33(1):124–139, October 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910416.

Hsieh:2000:FIP

[yHHsH^{+}00] Sun yuan Hsieh, Chin-Wen Ho, Tsan sheng Hsu, Ming-Tat Ko, and Gen-Huey Chen. A faster implementation of a parallel tree contraction scheme and its application on distance-hereditary graphs. Journal of Algorithms, 35(1):50–81, April 2000. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677499910647.

Yanez:1999:PDA

[YM99] Javier Yáñez and Javier Montero. A poset dimension algorithm. Journal of Algorithms, 30(1):185–208, January 1999. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909749.

Young:2000:LPA

[You00] Neal E. Young. On-line paging against adversarially biased random inputs. Journal of Algorithms, 37(1):218–235, October 2000. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S019667740091099X.

Yang:2004:DTS

[YW04] Boting Yang and Cao An Wang. Detecting tetrahedralizations of a set of line segments. Journal of Algorithms, 53(1):1–35, October 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677404000690.

Yao:1997:DLO

[YY97] Andrew C. Yao and Frances F. Yao. Dictionary look-up with one error. Journal of Algorithms, 25(1):194–202, October 1997. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677497908750.

Zantema:1991:MSA

[Zan91] H. Zantema. Minimizing sums of addition chains. Journal of Algorithms, 12(2):281–307, June 1991. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
019667749190005J.

Zepeda:2009:PSM

[ZC09] Claudia Zepeda and José Luis Carballido. P-stable models of strong kernel programs. Journal of Algorithms, 64(1):41–50, January 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677409000078.

Zhu:1991:AOG

[ZG91] Binghuan Zhu and Wayne Goddard. An algorithm for outerplanar graphs with parameter. Journal of Algorithms, 12(4):657–662, December 1991. CODEN JOALDV. ISSN 0196-6774 (print),

1090-2678 (electronic). URL http://www.sciencedirect.com/
science/article/pii/019667749190038Z.

Zhang:1998:VTG

[Zha98] Yanjun Zhang. The variance of two game tree algorithms. Journal of Algorithms, 28(1):21–39, July 1998. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677498909282.

Zhu:1995:PAH

[Zhu95] X. D. Zhu. A polynomial algorithm for homomorphisms to oriented cycles. Journal of Algorithms, 19(3):333–345, November 1995. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677485710401.

Zhu:2004:AQP

[Zhu04] An Zhu. Analysis of queueing policies in QoS switches. Journal of Algorithms, 53(2):137–168, November 2004. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677404000756.

Zhou:1996:ECP

[ZiNN96] Xiao Zhou, Shin ichi Nakano, and Takao Nishizeki. Edge-coloring partial k-trees. Journal of Algorithms, 21(3):598–617, November 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677496900619.

Zhang:2009:IAA

[ZLL09] Tongquan Zhang, Weidong Li, and Jianping Li. An improved approximation algorithm for the ATSP with parameterized triangle inequality. Journal of Algorithms, 64(2–3):74–78, April/July 2009. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677408000412.

Zerovnik:1993:CDM

[ZP93] J. Zerovnik and T. Pisanski. Computing the diameter in multipleloop networks. Journal of Algorithms, 14(2):226–243, March 1993. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677483710114.

Zhou:1996:LAE

[ZSN96] Xiao Zhou, Hitoshi Suzuki, and Takao Nishizeki. A linear algorithm for edge-coloring series-parallel multigraphs. Journal of Algorithms, 20(1):174–201, January 1996. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496900085.

Zhou:1997:PAE

[ZSN97] Xiao Zhou, Hitoshi Suzuki, and Takao Nishizeki. An NC parallel algorithm for edge-coloring series-parallel multigraphs. Journal of Algorithms, 23(2):359–374, May 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677496908305.

Zhang:1994:ATM

[ZSW94] K. Z. Zhang, D. Shasha, and J. T. L. Wang. Approximate tree matching in the presence of variable length don’t cares. Journal of Algorithms, 16(1):33–66, January 1994. CO-DEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677484710030.

Zverovich:2006:NKG

[Zve06] Igor E. Zverovich. A new kind of graph coloring. Journal of Algorithms, 58(2):118–133, February 2006. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0196677405000532.